### v2

# Vistoria: A Multimodal System to Support Fictional Story Writing through Instrumental Text–Image Co-Editing

Abstract.
Humans think visually—we remember in images, dream in pictures, and use visual metaphors to communicate. Yet, most creative writing tools remain text-centric, limiting how authors planing and translating ideas. We present Vistoria, a system for synchronized text–image co-editing in fictional story writing that treats visuals and text as co-equal narrative materials. A formative Wizard-of-Oz co-design study with 10 story writers revealed how sketches, images, and annotations serve as essential instruments for ideation and organization. Drawing on theories of Instrumental Interaction and Structural Mapping, Vistoria introduces multimodal operations-lasso, collage, filters, and perspective shifts that enable seamless narrative exploration across modalities. A controlled study with 12 participants shows that co-editing enhances expressiveness, immersion, and collaboration, enabling writers to explore divergent directions, embrace serendipitous randomness, and trace evolving storylines. While multimodality increased cognitive demand, participants reported stronger senses of authorship and agency. These findings demonstrate how multimodal co-editing expands creative potential by balancing abstraction and concreteness in narrative development.
Multimodality, Creativity Support, Storytelling, Creative Writing, Instrumental Interation, Structure Mapping, Direct Manipulation
journalyear: 2026copyright: acmlicensedconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Creant el demà junts, Barcelonabooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Creant el demà junts, Barcelona
Figure 1. Vistoria supports a cyclic workflow in which multimodal artifacts and ideas co-evolve. (A) Instrument Interaction: a unified set of tools—Collage, Lasso, Filter, and Perspective Shift—enables coupled editing of image and text to produce image–text “cards.” (B) Artifacts → Idea: the resulting visual and textual artifacts stimulate new concepts and story directions. (C) Cognitive Support: a structural mapping view aligns image and text variations, making alternatives comparable and keeping narrative structure consistent. (D) Ideas → Artifacts: emerging ideas are materialized back into new cards, closing the loop and driving iterative ideation, exploration, and integration.
1. Introduction
Human cognition is deeply rooted in visual processing (Rezaee, 2011; Huang et al., 2024; Kanellopoulou et al., 2019). We recall experiences as spatial scenes, build mental models through imagery, and rely on techniques such as memory palaces to organize abstract information. Everyday reasoning often involves manipulating internal visual representations—planning a route, rearranging objects in mind, or picturing alternative outcomes. Language use is similarly tied to imagery—comprehension frequently evokes mental pictures, and abstract ideas are often conveyed through spatial metaphors such as path, framework, or perspective (Meng et al., 2023). This aligns with Dual Coding Theory, which argues that cognition integrates both verbal and nonverbal channels, with distinct representational capacities, to support imagination, reasoning, and communication (Clark and Paivio, 1991).
Given this inherent multimodality of thought, a critical question arises: why do our most sophisticated creative practices, such as narrative writing, remain confined to unimodal tools? Fictional story writers often describe richly visualized scenes and manipulate complex spatial relationships in their imagination, yet must translate these processes through the narrow channel of sequential text (Mukramah et al., 2023; Listyani, 2019; Rao et al., 2024; Wang et al., 2025c). This mismatch between the multimodal nature of creative cognition and the unimodal design of creative tools constrains expressive potential (Zhao et al., 2025; Lee et al., 2024).
In fictional story writing, the most demanding phases are planning and translating (Gero et al., 2022; Flower and Hayes, 1981). Authors must organize imagined scenes, characters, and events into coherent structures, then render these abstractions into linear prose (Doyle, 1998; Bal and Veltkamp, 2013; Chung et al., 2025; Dang et al., 2023). Visuals can scaffold these processes by stimulating ideas, supporting organization, and enriching narrative detail (Mukramah et al., 2023; Listyani, 2019; Ali and Parikh, 2021). However, traditional visual aids have been limited in scope and scalability, offering only passive inspiration rather than dynamic, interactive support for the iterative writing process (Kreminski et al., 2024; Choi et al., 2024; Yan et al., 2023; Mukramah et al., 2023; Listyani, 2019). Large language models (LLMs) have unlocked unprecedented generative capabilities, enabling rapid creation of diverse narrative elements through combined text and visual imagery at previously unattainable scales (Fu et al., 2024). LLM-powered systems have explored visual integration through character construction tools (Qin et al., 2024), visual analytics for branching storylines (Mishra et al., 2025; Wang et al., 2025a), script visualization (Rao et al., 2024), diagram-prose synchronization (Masson et al., 2024b), and text-editing metaphors from image manipulation (Masson et al., 2024a).
However, the use of visuals remains peripheral across these approaches. They are often treated as prompts, overlays, or loosely coupled representations with unidirectional interaction. Our formative Wizard-of-Oz (WOz) co-design study with 10 writers revealed that sketches, annotations, and image manipulations are essential instruments for externalizing mental images and organizing narrative ideas. What is missing from existing LLM-powered creativity tools for fictional story writing is a unified multimodal substrate that directly supports iterative oscillation between planning and translating, where text and visuals serve as co-equal, manipulable materials in a bidirectional co-editing loop.
To bridge this gap, we developed Vistoria, a system that transforms fictional story writing from a unimodal text process into a multimodal co-editing experience where synchronized text-image manipulations through polymorphic operations—lasso, collage, perspective shifts, and filters. The design enables writers to fluidly navigate between abstract conceptualization and concrete visualization while creating a unified multimodal substrate for narrative development. These operations enable writers to reorganize narrative elements, explore alternative viewpoints, and refine story details through a bidirectional loop where text and visuals continuously inform one another. The design of Vistoria was grounded in two key theoretical frameworks. Instrumental Interaction guides consistent operations across modalities to reduce switching costs and sustain flow (Beaudouin-Lafon, 2000; Shi et al., 2025; Riche et al., 2025), while Structural Mapping (Gentner, 1983; Norman, 2013) explains how verbal and visual representations can be aligned so edits in one modality inform the other.
We conducted a controlled study with 12 participants to evaluate Vistoria’s effectiveness in supporting multimodal narrative creation. The results indicate that participants successfully integrated Vistoria’s synchronized text-image co-editing into their creative workflows, employing four distinct multimodal strategies to enhance storytelling. Participants demonstrated enhanced creative expression by using images as catalysts for richer descriptions, leveraging visual editing operations to drive narrative pivots, and employing multimodal elements to externalize and interconnect story branches. The cross-modal dialogue facilitated creativity through divergent exploration of multiple narrative directions, serendipitous discoveries from AI-generated content, and culture-driven associations that activated personal memories and references. While multimodality increased cognitive load, study participants reported heightened authorship, agency, and immersion, characterizing the system as a collaborative partner that expanded their imagination. Ultimately, creativity emerged from the dynamic interplay between modalities, where text and imagery continuously shaped one another.
In summary, this work contributes:
  • A WOz co-design study with 10 writers examined the practices and needs of using multimodal elements (text, sketches, and visual imagery) to externalize ideas and develop narratives in the planning and translating phase of fictional story writing;
  • Vistoria, a multimodal co-editing system that unifies text and visual images through polymorphic operations and a bidirectional editing loop to support planning and translating in fictional story writing;
  • A controlled usability study with 12 participants demonstrates the effectiveness of Vistoria, showing that multimodal co-editing enhances expressiveness, coherence, and authorial agency in fictional story writing.
2. Related Work
2.1. Using Visuals to Support the Cognitive Process of Fictional Story Writing
Fictional story writing is distinct from argumentative or expository genres in its emphasis on imagination, world-building, and character development (Doyle, 1998). Writers must invent narrative worlds and characters while ensuring coherence, which poses unique cognitive challenges: abstract, non-verbal mental images must be transformed into structured narrative elements and then into text (Bal and Veltkamp, 2013).
The Cognitive Process Model of Writing (Gero et al., 2022; Flower and Hayes, 1981) frames writing as recursive processes of planning, translating, and reviewing. In fictional story writing, the transition from planning to translating is especially demanding, as writers move from imaginative constructs to linear verbal representation, imposing a high cognitive load. However, visual representations can scaffold this process by externalizing abstract ideas. Research shows that picture prompts improve writing coherence (Mukramah et al., 2023), and visual images stimulate creativity in narrative writing (Listyani, 2019). In practice, sketches, maps, and diagrams help structure plot, setting, and character relationships during planning, and serve as cognitive anchors during translating to support coherence and consistency (Zhao et al., 2025). Dual Coding Theory (Clark and Paivio, 1991) explains these benefits: verbal and imagery systems function separately but also interact, creating richer memory traces when information is encoded in both modalities. In fictional writing, visual representations of narrative elements complement verbal planning, making abstract concepts more concrete and retrievable. When writers encounter difficulties in translation, visual anchors provide alternative access to imaginative content, reducing cognitive load and enabling more fluid expression (Barsalou, 2008).
Building on this foundation, a variety of creativity support tools have incorporated visuals into different phases of writing. Planning-focused systems such as CCI, Sketchar, and CharacterMeet assist authors in character and world development, often through image-guided backgrounds or conversational refinement of characters (Park et al., 2024; Long et al., 2024; Qin et al., 2024). Translation-focused tools like ScriptViz and Script2Screen aim to align textual composition with visual referents, either by retrieving reference visuals from movie databases (Rao et al., 2024) or by synchronizing scriptwriting with audiovisual scene creation (Wang et al., 2025c). Complexity management systems, for example, WhatIF (Mishra et al., 2025), ClueCart (Wang et al., 2025a), and PlotMap (Wang et al., 2024), help writers maintain structural coherence by visualizing branched narratives, organizing narrative clues hierarchically, or integrating spatial layouts with textual plot structures.
Despite these advances, visuals in prior systems remain peripheral: they serve primarily as references, scaffolds, or analytic overlays, rather than as co-equal, manipulable materials. This treatment constrains multimodal creative expression by limiting writers to either text-first workflows or loosely coupled visual aids (Zhao et al., 2025). To bridge this gap, our work introduces a system that treats visuals and text as fully integrated, co-editable artifacts, directly supporting the planning and translating phases of fictional story writing so that imagination can be externalized and refined seamlessly across both modalities.
2.2. LLM-powered Multimodality Tools for Creativity in Content Creation
Recent multimodal creativity tools move beyond linear prompting by enabling direct manipulation of creative elements, helping creators express intentions that language alone cannot capture (Riche et al., 2025; Masson et al., 2024a). Powered by advanced LLMs, these systems address fundamental barriers through three complementary mechanisms. First, externalizing creative structures. Tools like AI-Instruments (Riche et al., 2025) and Brickify (Shi et al., 2025) turn abstract intentions into manipulable interface objects or reusable design tokens, making spatial, stylistic, and compositional relations tangible. Second, reducing cognitive load with multimodal input: DrawTalking (Rosenberg et al., 2024) and Code Shaping (Yen et al., 2025) distribute expression across sketch, voice, or annotations, lowering the reliance on precise verbal specification. Third, supporting iterative refinement. Inkspire (Lin et al., 2025) and AIdeation (Wang et al., 2025b) accelerate variation and exploration, enabling rapid cycles of sketch-to-output or recombination of references. However, most current systems remain fragmented: they either support isolated stages (e.g., ideation, editing) or couple modalities only loosely, often within a one-shot prompting paradigm that underutilizes LLMs’ potential for sustained dialogue and refinement (Dang et al., 2023).
Seeking tighter coupling for fictional story writing, recent systems push integration of multimodal interaction in different ways (Chung et al., 2025; Chung and Kreminski, 2024). WorldSmith grounds worldbuilding in text–sketch edits and region-based filling to support users iteratively visualize and modify elements of their fictional world (Dang et al., 2023). XCreation (Yan et al., 2023) supports cross-modal storybook creation by integrating an interpretable entity-relation graph, improving the usability of the underlying generative structures. Toyteller (Chung et al., 2025) maps symbolic motions to character actions. Visual Writing allows writers co-edit stories by directly manipulating visual representations of entities, actions, and events alongside text, enabling easier navigation, editing, and exploration of narrative variations (Masson et al., 2024b). These systems highlight the transformative potential of LLMs to couple narrative and visual modalities, yet they still lack a unified, synchronized substrate that supports fictional story writing as a continuous, multimodal workflow. Our work advances this direction by introducing such a substrate, where LLM-generated texts and visuals act as co-editable inspirational materials for iterative, coherent story development.
2.3. Instrumental Interaction and Structural Mapping Theory
Instrumental Interaction is central to understanding how users control and refine digital systems. Beaudouin-Lafon (Beaudouin-Lafon, 2000) proposed it as a shift from designing static interface elements to designing instruments that mediate between users and domain objects. A key principle is reification, which transforms abstract commands into persistent, manipulable objects (Kent, 2000). In computing, this elevates implicit system descriptions into explicit first-class entities. In LLM-assisted workflows, this principle appears in modular prompt blocks for structured edits (Yen et al., 2023) and in Textoshop’s reification of abstract text editing commands (e.g., tone adjustment, boolean operations, layers) into direct manipulation tools for text (Masson et al., 2024a). A second principle is polymorphism, where the same instrument applies across contexts (Masson, 2023). This reduces cognitive load by enabling predictable, transferable patterns, e.g., copy–paste works consistently across text, images, and files (Angelini et al., 2015), and scrollbars operate similarly across documents, spreadsheets, and browsers (Masson, 2023). Finally, reuse allows users to replay or adapt prior operations, from macros to redo commands (Riche et al., 2025). Systems like Spacetime exemplify this by objectifying space, time, and actions into persistent containers, enabling edits to be carried forward as manipulable entities (Xia et al., 2018). Together, these principles reduce cognitive burden by externalizing interaction histories, making them manipulable, transferable, and extensible.
In our system, we extend this perspective to AI support for fictional story writing, where imagination must be continuously externalized and revised. We design interaction instruments that reify narrative elements into manipulable objects and, through polymorphism, let the same tools operate across LLM-generated text and images. This integration of reification and polymorphism enables writers to fluidly adjust multimodal outputs as persistent, tangible entities.
Structural Mapping Theory (SMT) is central to understanding how users interpret and compare complex representations. Gentner (Gentner, 1983) proposed SMT to describe how people form abstractions through analogical alignment of relational structures. Rather than relying on surface similarities, SMT highlights the role of alignable differences, which reveal systematic variations and help learners construct new knowledge. Building on this foundation, SMT-informed design emphasizes preserving object structure rather than abstracting it away. At the collection level, showing complete objects allows users to perceive relationships across the dataset (Yan et al., 2022; Wu et al., 2020). At the part level, systems can highlight structural correspondences by surfacing identity matches (literal overlaps) (Glassman et al., 2015), property matches (shared attributes) (Jasim et al., 2022), and conceptual matches (semantic commonalities) (Glassman et al., 2018). These correspondences support the iterative process of aligning, re-aligning, and refining mental models. Prior visualization and learning systems demonstrate the benefits of making such correspondences explicit, between code solutions (Glassman et al., 2015), textual variants (Glassman et al., 2018), or large language model outputs (Arawjo et al., 2024; Gero et al., 2024), helping users notice critical differences, generalize across examples, and avoid disengagement when facing “walls of text.” SMT thus serves as both a cognitive theory and a design constant for interactive systems that deepen users’ reasoning about collections of related objects.
In our system, SMT informs the design of synchronized text–image co-editing: textual fragments are explicitly mapped to visual elements based on their shared properties, matched concepts, or references to the same object, so edits in one modality lead to structurally aligned updates in the other, allowing users to compare, reorganize, and refine story elements across modalities.
3. Formative Study
Previous research shows that multimodal tools enhance fictional story writing by making abstract concepts tangible, reducing cognitive load, and improving creativity and coherence (Chung et al., 2022; Chung and Kreminski, 2024; Dang et al., 2023; Zhao et al., 2025). However, current tools treat images and texts as supplementary rather than integral to the creative process, leaving unclear how writers actually integrate multiple content types.
To address this gap, we conducted a Wizard-of-Oz co-design study (Dahlbäck et al., 1993) examining how creators use multimodal content (images, text, sketches) when planning and drafting fictional stories (Zhao et al., 2025; Gero et al., 2022; Flower and Hayes, 1981). Our investigation focused on three questions: (1) Multimodal information use: what types of multimodal content users employ and how they leverage these materials for idea generation; (2) Iteration and integration: how creators refine and combine multimodal artifacts in world-building and narrative development; and (3) Organization of inspirations: how creators organize, connect, and refine dispersed inspirations through multimodal manipulation. The WoZ setup simulated AI-assisted visual and textual support while sustaining the impression of an intelligent, interactive system.
3.1. Process
We designed a Wizard-of-Oz (WoZ) co-design study, positioning participants as active co-designers and treating text, sketches, and images as shared design materials (Dahlbäck et al., 1993; Sanders and Stappers, 2008; Vaajakallio and Mattelmäki, 2014).
3.1.1. Participants
We recruited 10 participants, each with an two or more years of creative writing experience. The group included three fictional story writers, three animation scriptwriters, two visual film creators, one new media creator, and one online fiction writer. Eight participants held a master’s degree or higher, and two held a bachelor’s degree.
Experimental Setup Three days prior to the session, participants were instructed to prepare a brief fictional story outline consisting of several sentences that followed one of the narrative structures from The Seven Basic Plots (Booker, 2004), which served as the foundation for subsequent ideation and content development. The 90-minute main session took place in Figma(Figma, 2025), which provided a collaborative canvas for free sketching, arranging text–image materials, and treating multimodal inputs as manipulable artifacts.
Session Process During the session, participants engaged in fictional story co-design through activities including generative prompts, collage, and storyboard-like arrangement while interacting with a simulated AI assistant operated by the researcher (wizard). Rather than working toward a fixed output, participants iteratively developed stories of about 300 words while envisioning how the tool itself should behave. They freely requested content through voice, text, or hand-drawn sketches, and called upon the assistant for image and text editing as new needs emerged. The wizard responded in situ to simulate intelligent system behaviors, encouraging naturalistic exploration. To support this process, the experimenter used Claude (Anthropic, 2025) for text generation (drawing on outlines and randomness to stimulate ideation) and Midjourney (Midjourney, 2025) and GPT-4o (OpenAI, 2024) for image generation from prompts, sketches, and editing instructions. Each session concluded with a 30-minute semi-structured interview probing how multimodal materials mediated co-creation, and what interaction patterns and workflows participants desired.
3.2. Formative Study Findings
3.2.1. Using multimodal input to reify vague ideas
As shown in the Table LABEL:tab:formative_grouped, participants utilized multimodal expressions, including text, sketches, and images, as co-design materials to articulate and negotiate intentions with the wizarded system.
Sketches externalized vague intentions and spatial imagination. For example, P4 envisioned a scene where clown JoJo appeared on stage and created a sketch showing “JoJo emerging from a giant spinning wheel” with textual annotations describing the intended atmosphere, hoping AI could elaborate narrative details. Text functioned as the primary medium for conveying intent, allowing participants to express connections between desired content and existing stories while clearly communicating their intentions. Participants frequently used textual annotations to specify story parts or image types for AI generation and guide generation direction. Images communicated style and mood expectations. P3 altered an AI-generated picture’s style by supplying a reference image, while P1 noted, “if possible, I want to use a ’supporting image’–a vague reference picture—as a basis, expecting the AI to generate more detailed images derived from it.”
Table 1. Results of user strategies for manipulating multimodal elements in the WOz co-design study.
Stage Observed behavior User need / Insight N Interaction Example
Multi-modal input Text Generated prose should match the established world while allowing the injection of new elements 83 [Uncaptioned image] [Uncaptioned image]
Multi-modal input Sketches Make spatial relations/composition concrete 8 [Uncaptioned image] [Uncaptioned image]
Planning LLM-Generated images Precisely locate inspiration from images 90 [Uncaptioned image] [Uncaptioned image]
Planning LLM-Generated text Filter usable bits from many generations and reuse them 90 [Uncaptioned image] [Uncaptioned image]
Planning/Translating Text notes / annotations Externalize temporary ideas for re-generation or final writing 30 [Uncaptioned image] [Uncaptioned image]
Planning/Translating Collage elements Recompose fragments from cross-image to form new scenes 12 [Uncaptioned image] [Uncaptioned image]
Planning/Translating Link elements Structure relationships between character / scene / object to connect the plot 40 [Uncaptioned image] [Uncaptioned image]
Translating Reconfigure elements in canvas Reorder scattered ideas by role / time / space into place for global understanding 4 [Uncaptioned image] [Uncaptioned image]
Translating Text (integration) Consolidate AI-generated text and image-inspired content into the draft 10 [Uncaptioned image] [Uncaptioned image]
3.2.2. Text–Image Interplay as Complementary Design Moves
Participants most often sought AI elaboration on characters, objects, and scenes, expressing the need for assistance with character design refinement, setting depictions, or object visualization. Combining image outputs with textual descriptions helped participants enrich their limited knowledge. For example, P7 requested AI-generated designs of an ancient Chinese poison bottle as a narrative element, noting her vague understanding of the concept.
These findings highlight the value of systems that accept heterogeneous multimodal input and help co-designers transform nascent ideas into concrete narrative materials.
Participants perceived text and images as distinct yet complementary in their creative writing. We observed an oscillation between abstraction (text) and concreteness (images) as a recurring co-design pattern, which reveals design opportunities for multimodal authoring tools.
Text as open imagination. Participants described text as a “blank canvas” for associative and structural thinking. P4 noted, “text allows me to imagine many things in my mind,” and P1 emphasized its “infinite imagination on a blank page.” P8 highlighted that text helped set up narrative structure before layering in visuals. This suggests systems should treat text as a flexible space for ideation and intent communication, where ambiguity can be preserved rather than prematurely resolved.
Images as concreteness, inspiration, and feedback. Images grounded abstract ideas, while their randomness often sparked unexpected inspiration. P1 explained: “The randomness in AI-generated images goes beyond what I want or can express; it helps me imagine the next step of the story.” Similarly, P4 refined Lily’s behavior based on an unexpected visual detail, and P3 used images as feedback for progressive refinement. P10 likened it to a “look-and-write exercise” that scaffolds scene construction. These accounts highlight the value of image outputs not just as illustrations but as provocations. Designers can leverage it through image-based iteration, selection, and reinterpretation.
Complementary interplay. Participants emphasized that neither modality sufficed alone: images “set the vibe,” while text reframed meaning. P8 noted, “Images can serve as references for appearance when I don’t have many ideas, while text quickly triggers associations.” We observed participants iteratively moving between text for open-ended imagination and images for concrete grounding, forming a cycle of divergence and convergence. This interplay suggests design opportunities for systems that orchestrate smooth transitions between modalities, enabling users to fluidly oscillate between abstract exploration and concrete elaboration.
3.2.3. Direct Manipulation of Multimodal Artifacts
Participants expressed a strong interest in treating text and images as manipulable, recombinable design materials. Two recurring practices pointed to design needs for more fluid multimodal manipulation.
Collaging and Recombination. Participants frequently merged elements across outputs to spark new ideas. P9 envisioned combining “the house from the first AI-generated images with the street from the second picture” to construct scenes, while P7 highlighted that “randomly combining characters and scenes” could inspire unexpected connections when accompanied by textual descriptions. As she explained, “if I can see an AI-generated image of my protagonist in one of the scenes, it helps me better imagine potential connections between elements that might otherwise seem unrelated.” Such practices illustrate the potential of collage and recombination as creative strategies.
Granular Editing and Annotation. Beyond recombination, participants desired fine-grained control over outputs. Sticky notes captured details for iteration and served as prompts for later development. Participants also wanted more localized operations, such as regenerating specific regions (P1, P3), extracting and reusing circled image elements (P5), or annotating character personas for refinement (P2). They left narrative prompts for later translation, e.g., P2’s note “Ending could be related to why this postman job even exists.” These behaviors emphasize annotation as both a vehicle for iteration and a bridge to subsequent writing.
Together, these findings suggest systems should enable flexible recombination, localized editing, and traceable annotations to help creators turn multimodal inspiration into actionable writing material.
3.2.4. From Fragmented Inspirations to Coherent Storylines
While participants often highlighted text or circled inspiring image details, organizing these dispersed fragments into coherent narratives was a persistent challenge. As P2 noted, “everything quickly became too messy on the canvas”, and P4 likened fragments on the canvas to “many cards that required connections,” where narrative coherence depends on linking passages, characters, and settings from scattered parts. Furthermore, P6 wished for mind map–like tools to scaffold this process. Participants also requested clustering notes, surfacing latent relations (e.g., by character/object/setting), and consolidating materials into reusable “setting cards” for ensuring cross-chapter consistency and avoiding logic conflicts (P4, P7).
These challenges point to opportunities for systems that transform fragmented inspirations into structured storylines by supporting clustering, relation mapping, and the creation of reusable narrative units that preserve coherence across iterations.
3.3. Design Goals
Drawing on insights from our WoZ co-design study, prior work on multimodal LLM tools, and theories of Structural Mapping and Instrumental Interaction (Section 2.3), we identify four design goals for a multimodal content creation interface that supports the planning and translating phase in fictional story writing (Gero et al., 2022; Flower and Hayes, 1981; Zhao et al., 2025).
  • DG1: Reifying Multimodal Intentions into Manipulable Artifacts. Informed by findings in Section 3.2.1, our system should reify multimodal inputs (text, sketches, images) into first-class artifacts that can be annotated, rearranged, and recombined. Drawing on Instrumental Interaction, such reification should preserve fleeting intentions as reusable materials for later iteration; following SMT, these artifacts should function as alignable units that support comparison, linkage, and cross-modal reorganization.
  • DG2: Aligning Text and Images for Iterative Creative Exploration. Informed by findings in Section 3.2.2, our system should enable fluid cross-modal iteration: textual edits can be re-visualized, image refinements can inform descriptions, and alternative versions should make differences explicit. Grounded in Dual Coding Theory and SMT, this alignment should support cycles of divergence and convergence by allowing users to oscillate between abstraction (text) and concreteness (images).
  • DG3: Enabling Polymorphic Cross-Modal Manipulation. Informed by findings in Section 3.2.3, our system should support direct manipulation interactions that apply uniformly across modalities, guided by Instrumental Interaction’s principle of polymorphism. This uniformity should reduce switching costs and empower writers to fluidly manipulate and edit textual and visual fragments while maintaining narrative coherence.
  • DG4: Organizing and Reusing Fragments into Coherent Narratives. Informed by findings in Section 3.2.4, our system should cluster and organize fragments and fleeting ideations, surface hidden connections, and consolidate dispersed inspirations into coherent, evolving narratives. It should further support the reuse of narrative units (e.g., characters, settings, objects) to ensure consistency and foster pathways for story progression.
4. Vistoria System
In this section, we present the key features of Vistoria. As shown in Figure 2, the interface comprises three primary components: a left text editor, a central collapsible cluster panel, and a right canvas interface. The text editor houses the current story draft, serving as contextual information for content generation. The right canvas supports freeform sketching, text input, and comprehensive editing tools. The central cluster panel aggregates highlights and annotations from canvas content, organizing them by characters, objects, and scenes for reference and summary.
Together, these components create a workflow where narrative text, visual artifacts, and organizational structures remain continuously aligned, enabling writers to externalize, manipulate, and recombine ideas and materials across modalities.
4.1. Key Features
Figure 2. Vistoria’s interface: a left text editor, a central Cluster panel(can be collapsed when not used), and a right free-form Canvas. Cards (image-text pairs with main object keywords) are generated from user intentions and serve as the basic units; two modes—Exact Craft and Creative Spark—support precise realization vs. exploratory variation.
4.1.1. Reifying Intention through Multimodal Generation
Writers often externalize early intentions through sketches, short notes, or reference images, but these expressions are typically fleeting and difficult to refine into coherent narrative materials. Supporting the preservation and refinement of such multimodal intentions is therefore essential for sustaining ideation (DG1). To address this, Vistoria reifies multimodal inputs into persistent Cards that bundle an image, a story fragment, and extracted narrative objects (characters, objects, scenes). Drawing on the principle of Reuse, Vistoria reifies sketches and text from ephemeral marks into manipulable and reusable artifacts that remain on the canvas and can be invoked again as elements for regeneration, enabling writers to carry forward and iteratively develop their creative intentions. At the same time, we treat cards as alignable units: text and their corresponding images are designed to carry the same underlying meaning, with images serving as visualizations of the narrative described in text.
Vistoria further balances precision and exploration by offering two complementary generation modes (Figure 3). In Exact Craft mode, single cards closely adhere to the author’s expressed intention to concretize specific ideas. In Creative Spark mode, three cards were generate represent diverse options based on writer’s intention, the system deliberately introduces variation around characters, settings, or objects, providing alternative prompts that can inspire new directions.
Figure 3. Vistoria balances precision and exploration through two complementary generation modes: in Exact Craft mode, a single card closely follows the author’s intention to concretize ideas; in Creative Spark mode, three cards are generated with deliberate variations in characters, settings, or objects, offering diverse options to inspire new directions.
Figure 4. A set of visual instruments for image-text co-editing to enhance planning and translating of fictional story writing: (a) Lasso selects regions for coupled image–text edits. (b) Writers extract elements and compose across cards to discover new narrative directions, using collage alongside other instruments to recombine sketches, images, and text on the canvas. (c) Perspective Shift changes an image’s viewpoint and automatically regenerates the story’s point of view (first/third/second person). (d) Filters align visual style and textual tone (e.g., melancholic/dreamy) by jointly altering image effects and rewriting prose.
4.1.2. Synchronized Image–Text Co-Editing through Visual Instruments
Fictional story writing benefits from fluid movement between abstract textual reasoning and concrete visual imagination, yet existing tools separate these modalities and disrupt creative flow (DG2, DG3). To address this, we introduce a set of visual instruments (Figure 4) designed around three principles: (1) Reification, which draws on familiar operations from image editing (e.g., lassoing, collaging, filtering) and transforms them into persistent, manipulable operations that apply equally to text and image elements (Swerzenski, 2021; Masson et al., 2024a); (2) Polymorphism, which ensures the same operations apply uniformly across text and images; and (3) Structural Mapping, which synchronizes changes so that visual and textual representations carry the same narrative meaning. Together, these principles reduce modality switching costs while maintaining coherence between story segments and imagery.
Lasso. The lasso instrument exemplifies reification by turning the abstract action of “focusing on part of a story” into a manipulable unit: selecting a region in either an image or a fragment of text generates a new card focusing on the selected part with enriched narrative and visual details. Through polymorphism, the same selection logic applies across modalities—whether circling a visual detail or isolating a text segment—providing a consistent interaction pattern. The text within the selected area in the original content will be emphasized to form a new card. The extracted portion of text is used to regenerate the corresponding image. Structural mapping ensures that the lassoed image region and the expanded story fragment correspond to one another, aligning visual and textual perspectives within the same narrative unit (Figure 4 (a) ).
Collage. The collage instrument reifies the abstract act of “recombining inspirations” into a tangible manipulation: fragments of images, sketches, or text can be directly composed within a collage frame to form a new card. Through polymorphism, the same cut–paste–combine logic applies uniformly across modalities—an image region, a text excerpt, or a sketch element can all be treated as compositional materials for intention-based generation. Guided by structural mapping, the system interprets the spatial arrangement of these multimodal pieces as narrative intent, generating a card where textual descriptions and visual depictions are aligned. For instance, merging two character fragments not only produces a combined image but also generates a new story segment situating them together, ensuring that narrative and imagery evolve in sync (Figure 4 (b)).
Filter. Stylistic coherence is critical in fictional story writing, as consistent affective and aesthetic cues sustain narrative transportation (Green and Brock, 2000), activate readers’ interpretive schemas (Arbib, 1992), and enhance the emotional resonance of literariness (Miall and Kuiken, 1999). In Vistoria, the filter instrument reifies this abstraction into a concrete tool: applying a “melancholic” or “dreamy” filter adjusts both the style of the image and the tone of the accompanying prose (Appendix Table 4). Through polymorphism and structural mapping, the same filter operation works seamlessly across modalities, leveraging the correspondence between visual style in images and emotional tone in text to act on both simultaneously. By making intangible stylistic intentions manipulable and synchronized, filters expand expressive possibilities while maintaining narrative immersion (Figure 4 (c)).
Perspective Shift. Fictional story writing often utilizes perspective shifts, as narratology highlights that changes in voice and focalization fundamentally reshape how events and characters are perceived (Genette, 1980). Cognitive poetics further shows that such shifts alter readers’ empathy and immersion. First-person narrations foster intimacy, while third-person perspectives enable broader structural awareness (Keen, 2007). The perspective-shift instrument reifies this narratological concept into an actionable operation: changing the visual viewpoint of a scene automatically regenerates the story fragment from a first-, third-, or second-person perspective. Through polymorphism, this instrument applies consistently across modalities, altering either an image or its accompanying text triggers a corresponding adjustment in the other. Structural mapping ensures that the focalization shift carries the same meaning across text and image: a new camera angle in the image corresponds to a new narrative voice in the text, allowing writers to explore empathy, distance, and structural awareness in a synchronized manner (Figure 4 (d)).
4.1.3. Highlight Elements and Cluster.
Writers often struggle to integrate scattered highlights and annotations on the Canvas into coherent storylines, leaving ideas fragmented across cards (DG4). Vistoria addresses this by transforming the dispersed fragments into reusable narrative building blocks, aligning them structurally across characters, objects, and scenes. This process is centered in the cluster panel (Figure 5), which turns fragmented inputs into organized knowledge assets.
Figure 5. Writers highlight objects and text segments on cards(b-1); the Cluster panel aggregates these by character/object/scene and can auto-summarize settings/plot/description about a certain object to guide final writing(b-2); Clicking on a specific image reveals the corresponding highlights and comments from earlier phases left on canvas(b-3).
On the canvas, writers can highlight textual segments, edit stories directly, and add inline comments noting potential uses in later drafting. Story objects, such as characters, settings, or scenes, are represented as editable keywords that can be highlighted by themselves. The system automatically binds each highlighted object to its associated text, consolidating references across multiple cards.
The cluster panel then aggregates all highlighted objects into an organized overview of evolving narrative elements. This eliminates the need to manually scan scattered cards and provides writers with a dynamically updated, object-centered workspace. Selecting an object reveals its complete associated materials, such as linked images, highlighted text segments, and comments, which creates a multimodal, context-rich reference for downstream writing.
Beyond simple aggregation, the panel supports higher-level knowledge construction through its summary feature. When this feature is invoked, the system generates structured summaries of settings, descriptions, and plot elements derived from highlights and comments. These summaries distill fragmented annotations into narrative building blocks (Burroway et al., 2022), enabling writers to iteratively scaffold coherent storylines from previously disjointed ideas.
4.2. Implementation
System Architecture. We adopted a decoupled front-end/back-end architecture. The React front-end enables efficient rendering for complex interactive interfaces, while the Flask back-end flexibly handles model calls with minimal overhead. Axios manages asynchronous communication between layers. The canvas module uses React-Flow for dynamic node-edge relationships and integrates “path to SVG” for natural interactions like freehand drawing and lasso selection. A DOM-based screenshot function ensures visual consistency in exported images. The text editor integrates React-Quill for rich text editing with lightweight customizations. Global state (text content, canvas nodes, selections) is centrally managed via Zustand to ensure consistency across modules. Per-session persistence relies on sessionStorage, which preserves data during the study session and clears on tab close to protect privacy.
Back-end Multi-Agent flow. The back-end bridges front-end actions with language and vision models through multi-agent orchestration. Every canvas action (e.g., selection, lasso, collage, inline text) is rasterized into screenshots paired with local context (outline, prior cards, global theme). A text agent (o4-mini) examines screenshots using templated prompts from a prompt library to extract user intent, recover sketch layouts, and write text consistent with surrounding outlines. The agent returns structured JSON (story text, intent, layout hints, semantic tags) that downstream visual agents use for synthesis and alignment.
Image Generation and Editing. An image agent invokes Flux Kontext Pro with consolidated prompts that fuse screenshot-derived intent, story fragments, and style controls. Sketches serve as structural scaffolds respecting composition while avoiding duplication; reference images act as content anchors, encouraging variation over copying. When neither is present, Flux operates in free mode guided by text-agent intent and style.For editing, front-end instruments (e.g., perspective shifts, filters, collage) are translated into prompt deltas through GPT-4o’s image pathway, yielding coupled visual updates. In parallel, the text agent revises tone or viewpoint for synchronized multimodal transformation. All artifacts are persisted as cards (image, story, objects keywords), enabling highlight-and-cluster operations that help reassemble fragments into coherent narratives.
5. User Study
To evaluate the usability, effectivness and usefulness of Vistoria, we conduct a lab user study with 12 participants. The study aims to answer the following research questions:
  • RQ1: What interaction patterns and strategies emerge when participants use a image-text co-editing process?
  • RQ2: How does image-text co-editing facilitate and shape the cognitive process of story creation, in terms of ideation, iteration, and refinement?
5.1. Participants
We recruited 12 participants (6 males, 6 females, aged 21–32, M=25.5), all with prior creative writing experience. Most held Bachelor’s or Master’s degrees, with backgrounds spanning science, arts, design, or communication. Their creative practices included fiction writing, screenwriting, songwriting, advertising, research, and philosophy. Participants’ creative writing experience ranged from under one year (n=5) to over seven years (n=1), with others reporting 1–3 years (n=2) or 4–7 years (n=3).
All participants were familiar with large language models (e.g., ChatGPT, Gemini, Claude) and had used them for idea generation, editing, descriptive support, content expansion, worldbuilding, and style imitation. This diversity reflects our design goal of supporting not only expert writers but also a broader range of users. Each participant received \$40 USD compensation.
5.2. Procedure
5.2.1. Apparatus
Sessions were conducted on a laptop computer with keyboard and mouse for typing, dragging, and selecting. To support sketch input, we provided an external tablet (iPad) for freehand drawing on the canvas.
The baseline condition presented a side-by-side interface with a text editor and GPT-4o conversational panel, enabling both manual editing and LLM-assisted text/image generation. Participants completed two story-writing tasks (Appendix A.1), each extending a given story beginning into a 300–500 word draft. Tasks were counterbalanced across conditions (Baseline vs. Vistoria).
5.2.2. Study Procedure
The study followed a within-subjects design with counterbalanced condition order. After informed consent and a demographic survey, participants were introduced to Vistoria through a written guide and tutorial video, followed by a short hands-on exploration (15 minutes).
In each condition, participants first focused on worldbuilding and idea exploration (20 min) and then on refining and improving the story (20 min). We divided the writing task into two phases (exploration and refinement) to prevent participants from prematurely committing to a single storyline and to reduce fixation, thereby encouraging broader ideation before focused improvement (Suh et al., 2024).
After each condition, participants completed surveys including NASA-TLX (Hart and Staveland, 1988), Creativity Support Index (CSI)(Cherry and Latulipe, 2014), self-perceived ML experience(Choi et al., 2024; Wu et al., 2022), and (for Vistoria only) perceived feature usefulness. All instruments used 7-point Likert scales. Following both conditions, participants completed a 15–20 minute semi-structured interview.
All sessions were video-recorded via Zoom. We collected system logs, final story drafts, canvas artifacts, image–text pairs, and interview transcripts for analysis.
5.2.3. Data analysis
We employed a mixed-methods approach to systematically analyze three types of data.
First, interview transcripts were independently open-coded by two authors following an inductive, grounded theory–informed process (Thornberg, 2012); initial codes were iteratively refined into a shared codebook, with themes developed through constant comparison. Disagreements were resolved via negotiated agreement, and analytic memos were maintained for transparency.
Second, interaction data were analyzed through structured video coding by two authors and aligned with system logs on an event-by-event basis. We examined the sequences of function use and compared exploration patterns across the baseline and our condition. To characterize participants’ divergent–convergent behaviors during the creative process, for each task, we reconstructed exploration structures by defining Directions as top-level trajectories toward a goal, Branches as the diversity of possibilities generated within a direction, and Depth as the mean number of iterative steps within each branch to compare the exploration across the baseline and our condition.
Finally, for survey measures, we conducted paired-sample t-tests under the assumptions of normality and homogeneity of variance. When assumptions were violated, we used the Wilcoxon signed-rank test.
6. Study Results
Figure 6. Interaction records of all participants. The creative workflow begins with multimodal generation—primarily text, complemented by sketches or images to express intentions, followed by refinement and iteration using visual instruments, during which textual descriptions are continuously revised in parallel. Definition of specific behaviors: Generation includes multimodal creation of new cards using Creative Spark, Exact Craft, or Collage; image-based editing refers to operations such as Lasso, perspective shift, and filters; text-based editing covers modifications of generated story segments on the canvas as well as edits made in the text editor; other operations include updates through highlight, cluster, and upgrading global settings.
Vistoria Baseline Statistics
mean std mean std p Sig.
Transparent of ML model Transparency 4.250 1.865 3.417 2.234 0.2015
Controllability 4.750 1.357 4.833 1.586 0.8664
Sense of Collaboration 5.583 1.505 4.333 1.670 0.0206 *
Support for Thought Process 5.333 1.371 4.750 1.960 0.3172
Match to goal 5.500 0.905 4.833 1.337 0.1201
NASA-TLX Mental 5.16 1.528 3.167 1.337 0.0000 **
Physical 4.667 1.723 2.083 0.669 0.0002 **
Temporal 2.917 1.443 2.750 1.215 0.7723
Effort 4.083 1.443 3.500 1.834 0.3388
Performance 5.250 1.712 5.083 1.564 0.7986
Frustration 2.750 1.183 1.750 0.622 0.0204 *
Creativity Support Index Exploration 4.917 1.240 4.750 1.485 0.7126
Expressiveness 6.083 0.996 4.333 1.775 0.0232 *
Immersion 4.917 1.505 2.750 1.545 0.0006 **
Enjoyment 5.333 1.435 4.917 1.379 0.1753
Results Worth Effort 5.250 1.357 5.583 0.793 0.5166
Collaboration 5.500 0.674 4.583 1.505 0.0418 *
Table 2. Comparison of survey results: Vistoria vs. Baseline. Sig.: * $p<.05$; ** $p<.01$
6.1. Strategies for Integrating Multimodality in Creative Writing
To address RQ1, we identified four key interaction strategies employed when using Vistoria.
First, participants leveraged multimodal elements to externalize and interconnect different branches of their ideas. Second, they utilized image-text pairs as catalysts to generate richer and more dynamic narrative descriptions. Third, participants employed visual instruments as a mechanism to drive pivotal shifts in their storytelling. Finally, they treated images as both checkpoints and traces, enabling them to maintain narrative coherence and facilitate recall throughout the story development process.
6.1.1. Use Multimodality to Support Expressing and Connecting Ideas
Log data (Figure 6) and observations (Figure 13) reveal that participants followed a consistent creative workflow when using the system. The process typically began with multimodal content generation, where participants combined generation functions across multiple modalities to visualize their intended content. During this exploratory phase, participants iteratively collected and refined story elements through generated cards(images-text pairs), at the same time, frequently modifying text in the editor or regenerating images based on updated text.
As participants explored multiple visualization directions, they employed the Collage function to combine existing canvas content and discover new narrative possibilities, or used visual instruments (Lasso, Filter, Perspective Shifts) to simultaneously modify text and images. This process involved continuous comparison with LLM outputs to determine desired versus undesired elements (P1, P2, P4, P8, P10). Participants systematically highlighted useful images and text for final translation into the text editor, often using Cluster functions to summarize character or object settings for incorporation into their final writing.
Multimodal Expression Enhances Creative Intention. The multimodal integration enhanced participants’ ability to convey creative intentions. As shown in Figure 6 (b), text served as the primary medium but was consistently supplemented with sketches and images for context, yielding significantly higher expressiveness ratings than the baseline ($M_{\text{Vistoria}}=6.08$ vs. $M_{\text{Baseline}}=4.33$, p = .023) (Table 2). P6 captured this benefit: “Even though I’m not very good at drawing, when I type some texts and sketch a few rough images, the cards with story and images generated can really capture the scene I have in mind.” (Figure 11). Nearly all participants (P2, P6, P10, P11) stressed that combining sketches with text and images aligned outputs more closely with their creative intentions. P7 illustrated this with a concrete case: she sketched a rough flower and archway, then added text specifying a glowing flower and a Gothic gate. The system fused the spatial layout from the sketch with textual details to generate multiple fitting images. She highlighted the unique value of sketching: “Drawing is pretty cool. In GPT, I want to draw a mental image, but GPT cannot… the geometry is always different from what’s in my head.“ (Figure 7). This suggests that multimodal expression enabled participants to externalize their mental imagery and refine it into concrete, shareable representations, bridging the gap between vague internal visions and precise creative outputs.
Figure 7. Multimodal Expression. Example where sketch structure + textual details jointly yield multiple relevant images; participants valued sketches to visualize spatial layout beyond what text-only tools could provide.
Collage Function Enables Creative Recombination. The Collage function emerged as a central strategy for creative recombination, receiving positive usability ratings (M=5.33, SD=1.5) (Figure 8). All participants used collage to merge extracted objects or scenes from different images, building connections across disparate elements. For instance, when P6 generated a scene of Maya entering a castle, he envisioned a larger structure with taller stairs. He sketched a bigger castle and mountain while expressing his intention through text, resulting in a generated card that matched his vision and was directly incorporated into his story (Figure 9). P11 articulated the creative freedom this technique provided: “This technique doesn’t limit me; I can create abstract or non-abstract sketches, and I can incorporate whatever I want.” This multimodal recomposition enabled participants to quickly express envisioned scenes (P2) and provided “more freedom to envision and create the story” (P10). These practices highlight that collage is not merely a usability feature but a catalyst for multimodal recomposition, enabling participants to externalize, reconfigure, and expand their mental imagery into coherent narrative possibilities that text or images alone could not achieve.
Furthermore, multimodal input also increased participants’ immersion with Vistoria compared to the baseline ($M_{\text{Vistoria}}=4.92$ vs. $M_{\text{Baseline}}=2.75$, p = .0006) (Table 2). This heightened immersion stemmed from the playful, enjoyable nature of multimodal creation and the emotional reinforcement of freely expressing intentions across modalities. As P1 reflected: “Using the tool feels more like… doing something on a whiteboard or a big piece of paper, where you can do almost anything—it’s very free and interesting.” Together, multimodal inputs and outputs fostered a more engaging and emotionally resonant creative experience.
Figure 8. The perceived usefulness of each key functionality in Vistoria
6.1.2. Instrumental Interaction Tools Drive Narrative Development
Analysis of interaction logs (Figure 6 and 13) shows that participants often engaged with visual instruments in the mid-phase of creation, and these instruments directly shaped how stories were developed and revised.
Perspective Shift as a narrative frame-shifter.
The Perspective Shift instrument, which altered both the image and the story perspective, was rated among the most useful features (M=5.92, SD=1.56) (Figure 8). P8 described how shifting perspectives changed the story direction: “Adopting the water’s viewpoint anthropomorphized the water spirit and introduced a regretful undertone that established the story’s emotional framework. Changing the perspective of the voice of the story added personal sentiment and changed the development direction, inspiring new writing possibilities.” P5 also experimented with this feature, incorporating a first-person voice (“I didn’t expect this to be so heavy!”) adopted from the system-generated segments into her third-person story (Figure 10 (a)). Perspective Shift allowed authors to flexibly reconfigure narrative viewpoint and voice, surfacing new emotional framings and redirecting story trajectories without disrupting their ongoing writing flow.
Lasso as a granularity controller for local-to-global rewriting.
The Lasso instrument was also valued as useful (M=5.0, SD=1.28) (Figure 8) for enabling writers to zoom between different narrative scales. P8 emphasized, “You can write in different scales, especially when you use the Lasso tool, in which you can extract out that specific detail, so [the story] generated in the card is more heterogeneous on the specific point.”(Figure 10 (b)) This reflects how the lasso tool enables a narrative “zoom” functionality, allowing writers to oscillate between macro-level story development and micro-level detail refinement within a single interface. Similarly, P9 and P11 described how the lasso allowed them to deepen or refine specific text and emphasize the key points with more focused attention. In this way, the Lasso operates as a narrative instrument, turning macro-level edits into micro-level adjustments while preserving precision, enabling rollback, and sustaining fluent exploration.
Figure 9. Collage. Participants used Collage for creative recombination, merging extracted objects/scenes (often mixing sketches with images) to specify visualization and advance story ideas; usability was positively rated.
Filters as affective parameterization for tone alignment.
The Filter instrument shaped narrative emotion and tone by visually parameterizing affect. P11 noted, “The image provides the style, which influences my story’s tone and direction… After using the filter, I can’t determine tone from text alone—I need to choose between suspenseful or romantic, and the visuals help me decide.” Similarly, P9 observed, “Filter visualizes the emotions in my mind when creating stories… I created a sad mood with filters, and the girl suddenly turned into a sad face. It’s very easy to see what it’s doing.” (Figure 10 (c)). The immediate visual feedback fostered a more intuitive workflow, aligning emotional intent with visual representation. The immediate visual feedback aligned emotional intent with representation, streamlining tone-setting decisions.
Taken together, these tools acted as narrative instruments that transformed localized operations into meaningful shifts in viewpoint, scale, and tone. They enhanced authors’ control, amplified expressiveness, and sustained a fluid perception–action loop throughout the creative process.
Figure 10. (a) Perspective Shift was rated highly useful (M=5.92); changing the image viewpoint also reframed narrative voice and redirected story development, as participants reported; (b) Filters (M=4.25) synchronize mood and style across media—applying a visual filter also rewrites the associated text to match the intended emotional tone, ensuring stylistic coherence; (c) Lasso tool (M=5.0) enabled writers to focus at different narrative scales by extracting or isolating elements to steer local and global edits.
6.1.3. Structural Mapping between Images-Text Paris as a Catalyst for Vivid Story Writing
Participants emphasized that when text and images appeared together, images served as direct references during writing, easing detailed description and enriching narrative depictions. This integration worked through two cognitive mechanisms that supported vivid story construction.
Images-Text Paris as cognitive scaffolding for unfamiliar scenarios. Visual references helped participants imagine actions or settings beyond their lived experience. As P11 reflected, “I’m more comfortable describing conversations and overall scenarios than people’s actions—especially kids, who might have more interesting reactions. Visualizing with text descriptions helps me imagine unfamiliar actions, which I can then use in my story.” Here, the image acted as a structural anchor by aligning the visual cues of children’s postures and expressions with possible textual descriptions. In other words, structural mapping between what was seen (visual structure) and what could be narrated (linguistic structure) scaffolded richer and more plausible action descriptions.
Images-Text Paris enabling descriptive language translation. Participants also converted visual elements into vivid prose, effectively mapping structures across modalities. As P5 noted, “Looking at an image makes it easier to kind of, like, describe what something is supposed to look like or feel like… so when you describe it to [readers], and they can picture it for themselves in their head, that is what makes the story more interesting.” From observation, P5 closely described the final images in her writing and directly adopted text from the generated cards, integrating them into her final story. (Figure 12) This visual-to-textual translation was also evident in concrete writing outcomes. For example, P1 used visual cues from an image showing Claire touching a letter and adopted the descriptions in the text such as “Claire steadies the box on her lap” and “She runs her fingers over the letters, heartbeats echoing in her ears.” to form his final story. (Figure 11) Here, the structural alignment between gesture in the image and embodied description in text allowed P1 to carry emotional undertones across modalities, turning static imagery into dynamic, sensory-rich narrative moments.
Overall, structural mapping between visual and textual representations scaffolded descriptive richness, enhanced vividness, and supported schema formation for scenarios participants might otherwise have struggled to imagine.
Figure 11. Visual-to-text translation: writers turn visual cues into vivid prose that readers can picture; sample lines show how P1 used sensory-rich descriptions derived from an image-text pair.
6.1.4. Structural Mapping across Multiple Dimensions to Trace the Story
Participants used not only text but also images as visual checkpoints to trace story direction and development. Because images were more immediately perceptible than text, they helped participants align outputs with intended style or mood and, when viewed in sequence, detect progressions and gaps (P6, P7, P2). Within Vistoria, this process evolved into structural mapping: participants systematically compared visual–textual structures across dimensions, gaining multimodal visibility into narrative development.
This layered view created a relational space where main storylines and branching alternatives could be mapped against one another, allowing participants to perceive both continuity and divergence, thereby fostering coherence while supporting flexible redirection. As P12 noted, visuals alongside text set the mood and made it easier to recall earlier ideas, illustrating how structural mapping provided mnemonic scaffolding that supported continuity in storytelling. Compared to GPT’s linear text-only workflow, participants found that Vistoria’s visual traces better supported narrative management and reduced idea drift or loss (P12, P11, P2). P2 reflected on this benefit explicitly:
“Using the system, we started with a baseline picture and then came up with another picture… it was easy to trace the development of the story. If you get too far down that chain and don’t like it, you can just delete that node and go in a different direction. I liked visually being able to see the progress from generation to generation.”
Here, structural mapping across successive images enabled participants to align branches and selectively prune or redirect them. Such multi-dimensional mapping clarified narrative progression, improved recall and coherence, and supported flexible branching and gap detection, sustaining sensemaking throughout the evolving story.
Figure 12. Images act as cognitive scaffolds, helping writers describe unfamiliar actions or contexts more concretely. P5 constructed her final writing by referring to both images and adopted and edit related text.
6.2. Multimodal Approaches Creation to Inspire Creativity
To answer RQ2, we found that image-text co-editing facilitated creativity in fictional story writing through three key cognitive processes: (1) divergent exploration, where participants spatially arranged multimodal components and pursued multiple narrative directions; (2) serendipitous discovery, where unexpected LLM outputs sparked new creative possibilities; and (3) multimodal cues activating memory and cultural connections. Together, these processes enhanced participants’ sense of authorship by positioning the system as a collaborative partner that expanded creative possibilities rather than replacing human imagination.
Figure 13. Behavioral diagram contrasting Vistoria vs. baseline: participants cycle through multimodal generation, collage/recombination, and coupled image–text editing before collecting highlights for integration. Data from P8-P12 show that, compared with baseline, participants using Vistoria explored more directions, with greater divergence (Branches) within each direction. Participants also tended to pursue deeper exploration within specific directions when using Vistoria.
Vistoria Baseline
Mean # of directions 6.92 $\pm$ 2.81 1.42 $\pm$ 1.08
Mean # of branches 3.00 $\pm$ 1.35 1.92 $\pm$ 0.67
Mean depth 1.70 $\pm$ 1.18 2.00 $\pm$ 1.22
Table 3. Descriptive statistics (mean $\pm$ SD) for Vistoria vs. Baseline. When using Vistoria, participants exhibited broader exploration; at the same time, as shown in Figure 13, they also tended to pursue individual directions with greater depth. Specifically, Directions denote the number of distinct aspects or dimensions explored when co-creation with Vistoria or baseline. Branches represent the diversity of possibilities generated within a given direction. Depth indicates the mean number of iterative steps within each branch.
6.2.1. Bottom-up Creation to Support Divergent Exploration
As shown in Figure 13 and Table 3, participants demonstrated significantly greater breadth and depth of exploration with Vistoria compared to the GPT baseline. They pursued a wider range of narrative directions ($M_{\text{System}}=6.92,SD=2.81$ vs. $M_{\text{Baseline}}=1.42,SD=1.08$) while also producing more variations within single paths (branches: $M_{\text{System}}=3.00,SD=1.35$ vs. $M_{\text{Baseline}}=1.92,SD=0.67$). Although the average depth per path was slightly lower for Vistoria ($M_{\text{System}}=1.70,SD=1.18$ vs. $M_{\text{Baseline}}=2.00,SD=1.22$), this reflected participants’ substantially broader exploration overall; within individual directions, they often engaged in deeper iterations, as evident in Figure 13. Such enhanced exploration transformed their creative process from passively consuming LLM outputs to actively curating multimodal components.
The Vistoria canvas functioned as an exploratory space where participants pursued multiple storylines in parallel without linear constraints. This freeform environment encouraged spatial arrangement, clustering, and repositioning of ideas through manipulation and recombination of alternative story–image cards to explore divergent scenarios (P7, P4, P5, P8). Rather than committing immediately to single narratives, participants typically generated multiple alternatives in early phases, positioning the system as an expressive medium rather than merely a text generator. This exploratory approach helped participants avoid early fixation and sustained creative engagement. As P8 observed: “GPT workflow is more streamlined… top-down. Using the visual app feels more bottom-up. You are open to possibilities, so there’s not a finite result and more possibilities being explored.” Ultimately, this shift highlights how multimodal, spatially organized exploration fosters divergent thinking, sustains creative momentum, and supports a more generative form of narrative authorship.
6.2.2. Serendipity from LLM Randomness Enriches Narrative Development
Another recurring theme was that the inherent randomness of LLM-generated image-text output often introduced serendipitous elements into participants’ creative process. Rather than being perceived as noise, these unexpected details were frequently embraced as inspiration, enriching story development and occasionally opening entirely new narrative directions.
For example, when P1 generated an image-text pair of the mysterious box, the system unexpectedly inserted the logo “LEH.” on the box and in the text description. Intrigued, P1 subsequently asked the system to visualize who or what “LEH” could be, gradually evolving this accidental motif into a recurring symbol within their final story (Figure 11). Nearly all participants described how surprises in LLM-generated images or texts gave them fresh possibilities to explore. P12 captured this dynamic vividly: “The system gave me a riddle at an early stage, I didn’t understand it at first, but as I kept working, I slowly decoded it, and this process of unraveling became part of my story. It helped me construct something unique that I wouldn’t have created otherwise.” (Figure 14) Together, these examples show how the interplay of multimodal generation and LLM’s inherent unpredictability fueled creativity. Randomness supported divergence by surfacing novel ideas, while participants retained agency in selectively integrating these elements, leading to unique, personally authored narratives.
Figure 14. In the creation process of P12, unexpected LLM randomness served as serendipitous prompts: a generated riddle was decoded and transformed into a key-and-note motif guiding Claire toward the lighthouse, illustrating how stochastic details became generative cues that enriched development and redirected the plot.
6.2.3. Unlocking Creativity Through Memory and Cultural Resonance
Participants reported that multimodal cues routinely triggered personal memories and cultural references, enriching their stories. Visual props, e.g., a medieval mask, readily invoked film/literary associations; as P4 noted, “Seeing the mask immediately brought Batman/Phantom to mind… it made choosing a style much easier.” (Figure 15) Crucially, image–text pairs co-triggered lived experiences. In P7’s case, a fountain scene image together with the phrase “mossy stone” jointly evoked a vivid memory, which she then translated into a concrete narrative element, the envisioned fall that became her story’s ending(Figure 15). These multimodal traces turned associative recall into usable story material, weaving authenticity and imaginative depth into the narrative.
6.2.4. Strengthened Authorship and Agency
Enhanced exploration with Vistoria strengthened participants’ sense of authorship. While GPT outputs in the baseline condition were consistently described as “surface-level” (P2, P5, P7) and felt like “someone else’s work” (P3), almost all participants reported markedly stronger authorship with Vistoria, enabling deeper exploration and richer plots (P1, P2, P3, P5, P12, P10). P5 articulated this shift:
“When using Vistoria, every idea originated from my own imagination, and the final story was formed by combining these different self-generated ideas. This gave me a strong sense that the story was truly my own creation. I think of the system less as a generator and more as a companion sketchbook. While you explore different directions, you always remain in charge of the final piece.”
P9 reinforced this sentiment: “This tool is more ‘me’… I control characters and plots; ChatGPT may not become my story.” Quantitative results corroborated these perceptions: participants rated Vistoria as providing stronger collaborative experiences (CSI: $M_{\text{Vistoria}}=5.50$ vs. $M_{\text{Baseline}}=4.58$, p = .0418) (Table 2). Participants framed Vistoria as a supportive co-pilot rather than a replacement; as P10 noted: “I consider this system as a copilot that does not help you with everything, and this system that you have to create by yourself.” Together, these findings mark a transition from passively adopting others’ ideas to actively curating one’s own generative outputs, reinforcing creative agency and ownership.
Figure 15. Evolving cards leave visual traces that cue personal memories and cultural references (e.g., a mask → Batman/Phantom; a “mossy stone” → place memory), which writers externalize into concrete scenes. These multimodal traces act as mnemonic scaffolds, helping retrieve earlier thinking and weave lived experience into the story world—transforming associative recall into narrative material.
6.3. Tradeoffs of Multimodal Interaction for Creative Writing
While image–text co-editing enhanced authorship, coherence, and expressiveness, it also increased workload. NASA–TLX scores showed significantly higher mental demand ($M_{\text{Vistoria{}}}=5.16$ vs. $M_{\text{baseline}}=3.17$, p=.0000) and physical demand ($M_{\text{Vistoria{}}}=4.67$ vs. $M_{\text{baseline}}=2.08$, p=.0002), plus moderately higher frustration ($M_{\text{Vistoria{}}}=2.75$ vs. $M_{\text{baseline}}=1.75$, p=.0204) (Table 2). These increases reflect the effort of coordinating across modalities and actively curating outputs relative to the GPT baseline. Part of the mental load may have stemmed from first-time use—learning new image/text operations and switching between modalities (P1, P2, P5, P9, P10): “The biggest burden is switching between tools to sketch or type; it takes time to learn and adapt, even though the functions are useful.” Additional difficulty arose from unfamiliarity with the canvas interface (P3, P10).
6.3.1. Higher burden of authorship.
Almost all participants valued the ability to maintain control of the story, and the creation process gave them a stronger sense of ownership. Yet, this empowerment came with a cost: several participants noted that they had to actively develop details within their own text, especially at the early stages, which could feel “a little bit frustrating” (P9). Unlike GPT, which could quickly produce long passages or propose questions to guide brainstorming (P10), the system required users to supply and elaborate on their own ideas before meaningful generation occurred, which may lead to a higher mental workload.
This shift demanded more cognitive and physical effort: even though sketching and annotation helped externalize mental imagery, participants noted that it felt more demanding than simply receiving GPT’s ready-made text (P1, P2, P11). Thus, stronger authorship agency came at the cost of higher workload.
6.3.2. Validating ideas rather than generating them.
Some participants noted that the system’s strengths lay in validating or expanding concrete mental images rather than generating abstract directions(P1, P11). As P1 explained, “when I have a vague impression in my mind, I tend to generate some image-text pairs. But sometimes, once the visual appears, it fixes my imagination in a certain way in my mind, and I can no longer imagine other possibilities. In contrast, only plain text can inspire limitless imagination.” This reveals a tension: images act as concrete anchors that aid detailed development, yet their representational specificity can induce fixation by prematurely crystallizing fuzzy concepts and narrowing exploration.
Similarly, P10 noted that the baseline GPT condition was superior at breaking down initial story points and offering prompt-like guidance, whereas the multimodal Vistoria system primarily served to elaborate or diverge from existing visions. This made the multimodal approach particularly valuable for writers with partially formed concepts, but less helpful during the most open-ended phases of ideation when abstract exploration is more important than visual specificity.
6.3.3. Cognitive offloading and efficiency despite higher load.
Despite a higher mental workload, participants repeatedly emphasized that multimodality enabled more efficient allocation of attention. From participants’ view, the image-text pairs externalized fleeting ideas, preserved spatial and sensory detail, and reduced information loss when translating imagination into tangible artifacts (P4, P7). As P7 noted, “By highlighting and collaging, I externalized formed ideas into image–text pairs, clearing mental space to pre-plan the next line and concentrate on the next plot beat.”
Thus, while Vistoria required more effort and decision-making, it also freed participants from lower-level memory maintenance and allowed them to focus on higher-order creative synthesis. With greater familiarity, the efficiency gains from this offloading may outweigh the initial overhead.
7. Discussion
7.1. Cross-modal LLM-Mediated Instrumental Interaction
We extend the notion of Instrumental Interaction into an LLM-mediated, cross-modal paradigm where generated artifacts repeatedly re-enter the workflow as manipulable outputs. In classic formulations, instruments mediate between users and domain objects, maintaining a separation between “tool” and “content” (Beaudouin-Lafon, 2000). In generative contexts, this dichotomy breaks down: each card (image–text pair) is simultaneously an outcome and a renewed object of operation, enabling recursive, non-linear composition (Riche et al., 2025).
This reframing also reshapes reification. Vague creative intents expressed through text, sketches, or images materialize across modalities into manipulable, cross-linked artifacts that remain editable and reusable. Reification thus spans modalities rather than existing in a single channel, supporting direct manipulation and reflection on both structure and semantics of evolving narratives (Riche et al., 2025; Masson et al., 2024b).
A key mechanism enabling this circulation is structural mapping (Gentner, 1983; Norman, 2013; Masson et al., 2024b). Participants made sense of their evolving materials not only by inspecting the surface properties of each medium but also by exploiting the correspondences in shared properties, concepts, or identities established between them. Refining a visual scene, for example, carried implications for narrative structure, tone, or pacing in text, while textual edits translated into compositional adjustments in images. In this way, mappings transformed what is usually a cognitive act of alignment into an operational resource: they stabilized coherence while keeping alternatives visible and accessible. Instrumental interaction and structural mapping thus operate as mutually reinforcing processes. Instruments generate the artifacts that mappings align, while mappings scaffold the interpretability and reuse of instruments.
This dynamic carries several design implications for future systems. They should not only offer isolated multimodal tools but cultivate instrumental ecosystems in which mappings are first-class and inspectable, and where families of instruments carry coherent semantics across domains (Riche et al., 2025). Furthermore, instead of overwhelming users with immediate complexity, operations can be gradually disclosed, allowing mappings to surface progressively and adapt to context (Shneiderman and Maes, 1997; Allen et al., 1999). Provenance views that expose the cascade of edits across modalities could support both divergence and recombination, helping individuals and teams understand how narratives branch and converge (Masson et al., 2024b).
Ultimately, the LLM can be positioned less as a generator of content and more as a mediator of instruments that align structures, propagate semantics, and synthesize new operations from prior materials. This shift reimagines authorship as an orchestration of evolving, cross-modal instruments, opening a design space where exploration and coherence can be sustained simultaneously (Sawyer, 2017; Masson et al., 2024b).
7.2. Orchestration and Curatorial Authorship to Form Creativity in Cross-Modal Creation
Traditional creativity theories emphasize individual expression through language and pre-formed intent (Smith, 2020; Donnelly, 2011), positioning the creator as the primary originator of content. Our findings challenge this view by demonstrating that creativity in Vistoria manifested as orchestration. Participants conducted and harmonized streams of multimodal LLM outputs through selection, combination, and reframing. Creative goals emerged through interaction with generated materials rather than preceding them, with meaning composed across representational boundaries. This suggests creativity in AI-augmented contexts should be theorized as the strategic coordination of evolving possibilities, repositioning authorship from origination to curation and synthesis within a collaborative ecosystem.
While existing multimodal creativity research often focuses on seamless integration between modes (Yen et al., 2025; Rosenberg et al., 2024; Shi et al., 2025), our study reveals that productive tensions between text and image became central to the creative process. Participants actively negotiated the trade-offs between imagistic specificity and narrative openness, treating these tensions not as problems to solve but as resources for exploration. The ongoing adjudication process—accept, adapt, defer, or discard–became the primary engine of creative coherence. This extends theories of creative constraint by showing how representational tensions across modalities can drive rather than hinder creative development.
Furthermore, our findings contribute to debates about human agency in LLM-assisted creation by demonstrating that curatorial work enhanced rather than diminished participants’ sense of ownership. Authorship was experienced as steering a dialogue with the model, including deciding what counts, where elements belong, and how disparate fragments accrue into a unified voice. This sense of agency grew precisely from having meaningful choices to negotiate among viable alternatives (Resnick et al., 2005). This suggests a new model of distributed authorship where human control emerges through deliberate curation within a space of LLM-generated possibilities.
As for design implications, our findings suggest that future system design should prioritize support for orchestration over seamless, one-shot generation. Instead of producing singular outputs, tools could by default surface multiple cross-modal alternatives, accompanied by lightweight mechanisms that enable users to accept, adapt, defer, or discard options as part of an ongoing multimodal curatorial process (Resnick et al., 2005). Moreover, systems should not only present outputs but also make the inherent tensions between modalities visible and adjustable. This can be done, for example, through interactive controls such as ambiguity sliders that allow users to navigate between imagistic specificity and narrative openness.
Beyond simple interface controls, future systems could also provide deeper support for curatorial authorship. For example, they might keep track of where each fragment or idea came from across different iterations, or offer reusable “curation moves” that let users quickly repeat and adapt common ways of combining or editing materials (Shneiderman, 2007). In this way, design emphasis shifts away from achieving frictionless integration and toward helping users actively steer evolving creative possibilities, foregrounding authorship in AI-mediated contexts as a matter of curation, synthesis, and negotiated choice (Ding, 2024).
7.3. Limitation and Future Work
While our study provides valuable insights into multimodal story writing, several limitations constrain the generalizability and scope of our findings.
Task Scope and Short-Term Focus. The 300–-500 word story task, while manageable for controlled evaluation, does not reflect the demands of long-form fictional writing, where authors build sustained voices, complex arcs, and intricate structures. We did not have an opportunity to study Vistoria’s suitability for extended projects, iterative revisions, or complex narratives, leaving open questions on the consistency in long works, scalability with larger volumes, or risks of over-reliance on LLM over the time of use. Moreover, evaluation relied primarily on self-reports of creativity and user experience; we did not include objective measures of story quality, originality, or literary merit.
In the future, we plan on conducting field studies to deploy this system with writers of varying expertise levels in their authentic creative contexts, observing how they integrate the tool into real writing projects over extended periods. Such longitudinal research would assess the ecological validity of Vistoria and provide insights into how writers adapt Vistoria for planning and translating across longer creative cycles. We anticipate that writers might spontaneously capture inspirational moments from daily life, potentially increasing their reliance on clustering functionality as they generate more dispersed content fragments that require organization. These naturalistic studies would provide crucial insights into the tool’s role in sustained creative practice, revealing usage patterns, adaptation strategies, and long-term impacts on writers’ creative processes that controlled laboratory settings cannot capture.
Participant Sample. Our study involved only 12 participants, while this is typical for similar lab usability studies, a larger group could provide stronger statistical power, reveal more varied interaction patterns, and allow comparisons across subgroups. Furthermore, the participant pool has limited cultural and age diversity, which could have narrowed the range of narrative traditions, writing styles, and storytelling approaches represented. Future research should address these limitations by recruiting a larger and more diverse set of participants, including professional authors, older and younger writers, and individuals from varied cultural backgrounds, to more fully evaluate the system’s applicability and generalizability.
8. Conclusion
This paper introduced Vistoria, a multimodal text–image co-editing system designed to support the cyclical “planning–translation” process in fictional writing. Grounded in Wizard-of-Oz co-design study and theories of Instrumental Interaction and Structural Mapping, the system integrates four key capabilities: (1) representing sketches, texts, and references as manipulable card units; (2) enabling cross-modal alignment so that edits in one modality propagate to the other; (3) providing polymorphic operations that apply consistently across text and image; and (4) supporting organization and reuse through highlighting and clustering panels. A controlled study with 12 creative writers demonstrated that Vistoria enhanced writers’ expressiveness, immersion, and sense of collaboration, while broadening exploration of narrative branches and reinforcing authorship and control. These findings suggest that Vistoria transforms the writing process into a dialogic interaction with multimodal instruments, enabling authors to negotiate meaning and control across modalities.
Appendix A Appendix
A.1. Writing Topics Used in the User Study
The two writing topic prompts used in the user study were:
Topic 1—Claire steps outside her apartment and finds a small wooden box on her doorstep. The box is secured with an old brass clasp and feels unexpectedly heavy when she lifts it. Its surface is scratched, as if it has been handled many times, and faint traces of dried sea salt cling to the edges.
Topic 2—During her morning jog through the park, Maya discovers an ornate iron gate hidden behind overgrown ivy. Through the bars, she can see a path lined with luminescent flowers that pulse gently like soft heartbeats. The air carries faint whispers in a language that sounds hauntingly familiar, almost like someone calling her name.
A.2. Filters
The following are the types of filter supported by the filter instrument, showing how different types of filters are applied to image styles and mapped to text tone or emotion.
Filter Image Effect Text Effect
Warm Warm tones (gold, amber, red, orange, yellow), high exposure, strong contrast $\rightarrow$ evoke happiness, comfort, nostalgia Emphasizes positivity, vitality, intimacy
Calm Cool tones (blue, green, purple) with balanced or lower saturation $\rightarrow$ convey calmness, wisdom, introspection Reflects contemplative and stable moods
Dramatic Deep blacks, sharp whites, directional lighting $\rightarrow$ create intensity, mystery, urgency Heightens stakes and emotional tension
Dreamy Soft tones, lowered contrast, diffuse focus $\rightarrow$ suggest melancholy, intimacy, ethereality Supports subtle, nostalgic, introspective narration
Monochrome Removal of color, emphasis on light, shadow, texture $\rightarrow$ evoke nostalgia, timelessness, artistry Adopts reflective and universal tone
Table 4. Filter types with corresponding image and text effects.
A.3. LLM Prompts
Communication between different agents is carried out through JSON-formatted messages, ensuring efficient and reliable information transfer.
All prompts are structured in the following format:
  • Blue text indicates placeholders that will be replaced with task-specific information in different generation requests.
  • Pink text represents the mandatory output type of information generated by the agent.
The below is an example prompt for precise description generation of GPT-4o.
You are a visual story developer who analyzes screenshots to decode user visualization intent and creates detailed story segments that bring their creative vision to life. Process: 1) Examine the screenshot to understand what specific story content the user wants generated by identifying: printed text({text} ) which is the primary indicator of what the user wants you to generate, handwritten text expressing the user’s desired content direction and story focus where generation should address gaps and missing details, any images or illustrations with reference text {previous_text} for additional context, and hand-drawn sketches representing scenes from the user’s imagination. Synthesize these elements to understand the user’s envisioned story. The generated story should mainly focus on filling in content not covered in ({text} ) instead of still remain unknown. 2) If hand-drawn illustrations exist in the screenshot, return the information about the story scene conveyed in the illustration’s layout; if none exist, output ’none’. 3) Generate Focused Story Content: Using the existing written passages {full_text} only as background context to ensure logical consistency, create a NEW detailed story segment that elaborates on a specific scene or moment the user wants to visualize, focuses primarily on the intent expressed in the screenshot rather than expanding the existing text, contains concrete details, character emotions, environmental descriptions, dialogue, and interactions, maintains consistency with the global theme {global_theme}, and can reference previous text {previous_text} if relevant to the visualization goal. 4) The narrative should contain substantial plot or setting content, not just descriptive language. The generated story should introducing new, insightful elements based on the context and provide new direction of the story development that can reificate the story. The generated stories need to be imaginative with concrete content, not filled with uncertainties. Respond in JSON format: { "story": "A detailed paragraph of no larger than 100 words that creates NEW story content focused on the user’s screenshot intent, elaborating a specific scene with concrete details while maintaining logical consistency with the background context.", "intention": "The visualization intention read from the screenshot for story generation direction", "sketch_information":"Regarding line sketches, integrate them with story descriptions to capture the user’s envisioned layout and scene details communicated through hand-drawn imagery, directing the image generator to produce story scenes based on this layout guidance. Avoid generating stories that may trigger content moderation." }. Provide only the JSON response without markdown formatting or additional commentary. Let’s think step by step.

### v3

@@
-Vistoria: A Multimodal System to Support Fictional Story Writing through Instrumental Text–Image Co-Editing
+Vistoria: A Multimodal System to Support Fictional Story Writing through Instrumental Image-Text Co-Editing
 Abstract.
 Humans think visually—we remember in images, dream in pictures, and use visual metaphors to communicate.
-Yet, most creative writing tools remain text-centric, limiting how authors planing and translating ideas.
-We present Vistoria, a system for synchronized text–image co-editing in fictional story writing that treats visuals and text as co-equal narrative materials.
-A formative Wizard-of-Oz co-design study with 10 story writers revealed how sketches, images, and annotations serve as essential instruments for ideation and organization.
-Drawing on theories of Instrumental Interaction and Structural Mapping, Vistoria introduces multimodal operations-lasso, collage, filters, and perspective shifts that enable seamless narrative exploration across modalities.
-A controlled study with 12 participants shows that co-editing enhances expressiveness, immersion, and collaboration, enabling writers to explore divergent directions, embrace serendipitous randomness, and trace evolving storylines.
-While multimodality increased cognitive demand, participants reported stronger senses of authorship and agency.
+Yet, most creative writing tools remain text-centric, limiting how writers plan and translate ideas.
+We present Vistoria, a system for synchronized image-text co-editing in fictional story writing.
+A formative Wizard-of-Oz co-design study with 10 story writers revealed how sketches, images, and text serve as essential elements for ideation and organization.
+Drawing on theories of Instrumental Interaction, Vistoria introduces instrumental operations—Lasso, Collage, Perspective Shift, and Filter that enable seamless narrative exploration across modalities.
+A controlled study with 12 participants shows that co-editing enhances expressiveness, immersion, and collaboration, opening space for writers to follow divergent story directions and craft more vivid, detailed narratives.
+While multimodality increased cognitive demand, participants reported stronger senses of ownership and agency.
 These findings demonstrate how multimodal co-editing expands creative potential by balancing abstraction and concreteness in narrative development.
-Multimodality, Creativity Support, Storytelling, Creative Writing, Instrumental Interation, Structure Mapping, Direct Manipulation
-journalyear: 2026copyright: acmlicensedconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Creant el demà junts, Barcelonabooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Creant el demà junts, Barcelona
+Multimodality, Creativity Support, Storytelling, Creative Writing, Instrumental Interaction
+journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790400isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Interaction paradigms
 Figure 1.
 Vistoria supports a cyclic workflow in which multimodal artifacts and ideas co-evolve.
-(A) Instrument Interaction: a unified set of tools—Collage, Lasso, Filter, and Perspective Shift—enables coupled editing of image and text to produce image–text “cards.” (B) Artifacts → Idea: the resulting visual and textual artifacts stimulate new concepts and story directions.
-(C) Cognitive Support: a structural mapping view aligns image and text variations, making alternatives comparable and keeping narrative structure consistent.
-(D) Ideas → Artifacts: emerging ideas are materialized back into new cards, closing the loop and driving iterative ideation, exploration, and integration.
+(A) Instrumental operations enable image–text co-editing.
+(B) Resulting image–text artifacts prompt new story directions.
+(C) Image–text alignment supports idea formation by coordinating verbal and non-verbal processing.
+(D) Emerging ideas are externalized into new artifacts, closing the loop for iterative ideation and integration.
+The figure illustrates Vistoria’s overall cyclic workflow for multimodal story writing.
+It shows how writers interact with image–text instruments to edit visual and textual artifacts, how these artifacts stimulate new story ideas, and how ideas are materialized back into new image–text cards.
+The cycle highlights continuous co-evolution between artifacts and ideas, emphasizing synchronized visual and verbal processing during iterative narrative development.
 1. Introduction
-Human cognition is deeply rooted in visual processing (Rezaee, 2011; Huang et al., 2024; Kanellopoulou et al., 2019).
-We recall experiences as spatial scenes, build mental models through imagery, and rely on techniques such as memory palaces to organize abstract information.
-Everyday reasoning often involves manipulating internal visual representations—planning a route, rearranging objects in mind, or picturing alternative outcomes.
-Language use is similarly tied to imagery—comprehension frequently evokes mental pictures, and abstract ideas are often conveyed through spatial metaphors such as path, framework, or perspective (Meng et al., 2023).
-This aligns with Dual Coding Theory, which argues that cognition integrates both verbal and nonverbal channels, with distinct representational capacities, to support imagination, reasoning, and communication (Clark and Paivio, 1991).
-Given this inherent multimodality of thought, a critical question arises: why do our most sophisticated creative practices, such as narrative writing, remain confined to unimodal tools?
-Fictional story writers often describe richly visualized scenes and manipulate complex spatial relationships in their imagination, yet must translate these processes through the narrow channel of sequential text (Mukramah et al., 2023; Listyani, 2019; Rao et al., 2024; Wang et al., 2025c).
-This mismatch between the multimodal nature of creative cognition and the unimodal design of creative tools constrains expressive potential (Zhao et al., 2025; Lee et al., 2024).
-In fictional story writing, the most demanding phases are planning and translating (Gero et al., 2022; Flower and Hayes, 1981).
-Authors must organize imagined scenes, characters, and events into coherent structures, then render these abstractions into linear prose (Doyle, 1998; Bal and Veltkamp, 2013; Chung et al., 2025; Dang et al., 2023).
-Visuals can scaffold these processes by stimulating ideas, supporting organization, and enriching narrative detail (Mukramah et al., 2023; Listyani, 2019; Ali and Parikh, 2021).
-However, traditional visual aids have been limited in scope and scalability, offering only passive inspiration rather than dynamic, interactive support for the iterative writing process (Kreminski et al., 2024; Choi et al., 2024; Yan et al., 2023; Mukramah et al., 2023; Listyani, 2019).
-Large language models (LLMs) have unlocked unprecedented generative capabilities, enabling rapid creation of diverse narrative elements through combined text and visual imagery at previously unattainable scales (Fu et al., 2024).
-LLM-powered systems have explored visual integration through character construction tools (Qin et al., 2024), visual analytics for branching storylines (Mishra et al., 2025; Wang et al., 2025a), script visualization (Rao et al., 2024), diagram-prose synchronization (Masson et al., 2024b), and text-editing metaphors from image manipulation (Masson et al., 2024a).
-However, the use of visuals remains peripheral across these approaches.
-They are often treated as prompts, overlays, or loosely coupled representations with unidirectional interaction.
-Our formative Wizard-of-Oz (WOz) co-design study with 10 writers revealed that sketches, annotations, and image manipulations are essential instruments for externalizing mental images and organizing narrative ideas.
-What is missing from existing LLM-powered creativity tools for fictional story writing is a unified multimodal substrate that directly supports iterative oscillation between planning and translating, where text and visuals serve as co-equal, manipulable materials in a bidirectional co-editing loop.
-To bridge this gap, we developed Vistoria, a system that transforms fictional story writing from a unimodal text process into a multimodal co-editing experience where synchronized text-image manipulations through polymorphic operations—lasso, collage, perspective shifts, and filters.
-The design enables writers to fluidly navigate between abstract conceptualization and concrete visualization while creating a unified multimodal substrate for narrative development.
-These operations enable writers to reorganize narrative elements, explore alternative viewpoints, and refine story details through a bidirectional loop where text and visuals continuously inform one another.
-The design of Vistoria was grounded in two key theoretical frameworks.
-Instrumental Interaction guides consistent operations across modalities to reduce switching costs and sustain flow (Beaudouin-Lafon, 2000; Shi et al., 2025; Riche et al., 2025), while Structural Mapping (Gentner, 1983; Norman, 2013) explains how verbal and visual representations can be aligned so edits in one modality inform the other.
-We conducted a controlled study with 12 participants to evaluate Vistoria’s effectiveness in supporting multimodal narrative creation.
-The results indicate that participants successfully integrated Vistoria’s synchronized text-image co-editing into their creative workflows, employing four distinct multimodal strategies to enhance storytelling.
-Participants demonstrated enhanced creative expression by using images as catalysts for richer descriptions, leveraging visual editing operations to drive narrative pivots, and employing multimodal elements to externalize and interconnect story branches.
-The cross-modal dialogue facilitated creativity through divergent exploration of multiple narrative directions, serendipitous discoveries from AI-generated content, and culture-driven associations that activated personal memories and references.
-While multimodality increased cognitive load, study participants reported heightened authorship, agency, and immersion, characterizing the system as a collaborative partner that expanded their imagination.
-Ultimately, creativity emerged from the dynamic interplay between modalities, where text and imagery continuously shaped one another.
+Human thinking involves multimodal processing.
+Visual processes play a central role in cognition: we recall experiences as spatial scenes, form mental models through imagery, and use visual structure to organize and interpret information  (Rezaee and Shoar, 2011; Huang et al., 2024; Kanellopoulou et al., 2019).
+Language is similarly entwined with imagery.
+Particularly, text comprehension often evokes mental pictures, and abstract ideas are commonly articulated through spatial metaphors such as path, framework, or perspective (Zhao et al., 2025; Clark and Paivio, 1991; Gärdenfors, 1996).
+Dual Coding Theory frames this coupling between imagery and language, positing that humans draw on both verbal and nonverbal channels to represent and support specific reasoning and communicative processes (Clark and Paivio, 1991).
+Story writing is particularly multimodal in nature.
+During the planning phase, experienced writers often use both imagery and language to construct the story world.
+They visualize spatial layouts, character interactions, and scene dynamics, while using textual notes to label, sequence, and reason about narrative structure (Doyle, 1998; Bal and Veltkamp, 2013; Chung et al., 2025; Dang et al., 2023).
+In the translating phase, visual details serve as an anchor that shapes how writers organize narrative detail and emotional tone, while texts linearize these visualized ideas into descriptions, dialogue, and narrative perspective that readers can follow (Mukramah et al., 2023; Listyani, 2019).
+Crucially, nonverbal and verbal channels do not operate in isolation: writers use imagery to trigger new wording, and emerging text in turn elicits further mental images (Zhao et al., 2025; Lee et al., 2024).
+Therefore, there is a great opportunity for tools that support story writing to match this multimodal complexity by accommodating the continuous interplay between visual and textual thinking within the same workflow.
+Yet, current writing tools remain overwhelmingly text-centric, treating linear text as the primary or sole medium of expression (Lee et al., 2024; Gero et al., 2023).
+Although recent systems powered by Large Language Models (LLMs) incorporate visual elements through image generation or retrieval, these visuals remain peripheral.
+They function mainly as prompts (Qin et al., 2024), static references (Rao et al., 2024), or organizational diagrams (Mishra et al., 2025; Wang et al., 2025b), rather than as tightly integrated, manipulable representations along with text.
+This requires writers to translate visual ideas back into text, increasing cognitive load (Brown and Kim, 2021; Zhao et al., 2025; Yan et al., 2023; Singh et al., 2023).
+Such translation overhead not only hinders cross-modal collaboration but also restricts the creative synthesis of semantically distant concepts (Singh et al., 2023) that would emerge more naturally in a truly integrated multimodal system.
+To examine this gap, we conducted a formative Wizard-of-Oz (WoZ) co-design study with 10 experienced writers to inform the design of a unified image-text multimodal system that supports the planning and translation phases of fictional story writing.
+We found that text and images play distinct yet complementary roles in the writing process.
+Writers expressed a need to directly manipulate text and images for fine-grained editing and alignment, valuing the ability to move fluidly between them.
+Based on these formative results, we developed Vistoria, a system for experienced, LLM-literate writers that reimagines fictional storytelling as a multimodal co-editing experience, integrating text, images, and sketches beyond traditional text-centric workflows.
+The design of Vistoria draws on the principle of Instrumental Interaction, designing a set of instrumental operations (lasso, collage, perspective shift, and filter).
+These functions can be applied to either text or images, where a single action simultaneously affects both images and text, minimizing switching costs and preserving creative flow (Beaudouin-Lafon, 2000; Shi et al., 2025; Riche et al., 2025).
+Based on Dual Coding Theory, Vistoria enables the alignment of text and visual representations, ensuring that edits in one modality are appropriately reflected in the other.
+We conducted a controlled study with 12 participants to examine how multimodal image–text co-editing supports fictional story writing, with a focus on evaluating the system’s usability and understanding its creative support.
+Overall, the study showed that Vistoria enhances expressiveness, immersion, and exploration, enabling participants to have more divergent ideas and write detailed narratives.
+Participants used the instrumental operations to refine ideas at multiple scales and explore alternative directions.
+While this workflow increased mental and physical workload, it also supported writers’ senses of agency and ownership, as they maintained greater operational control over narrative development.
 In summary, this work contributes:
-  • A WOz co-design study with 10 writers examined the practices and needs of using multimodal elements (text, sketches, and visual imagery) to externalize ideas and develop narratives in the planning and translating phase of fictional story writing;
-  • Vistoria, a multimodal co-editing system that unifies text and visual images through polymorphic operations and a bidirectional editing loop to support planning and translating in fictional story writing;
-  • A controlled usability study with 12 participants demonstrates the effectiveness of Vistoria, showing that multimodal co-editing enhances expressiveness, coherence, and authorial agency in fictional story writing.
+  • A WoZ co-design study with 10 writers examined the practices and needs of using multimodal elements to externalize ideas and develop narratives in the planning and translating phase of fictional story writing;
+  • Vistoria, a multimodal co-editing system designed for experienced, LLM-literate writers that unifies text and visual images through instrumental operations and a synchronized editing loop to support fictional story writing;
+  • A controlled usability study with 12 participants demonstrates the potential of Vistoria, suggesting that multimodal co-editing can enhance expressiveness, idea generation, and narrative development in fictional story writing.
 2. Related Work
@@
 Fictional story writing is distinct from argumentative or expository genres in its emphasis on imagination, world-building, and character development (Doyle, 1998).
-Writers must invent narrative worlds and characters while ensuring coherence, which poses unique cognitive challenges: abstract, non-verbal mental images must be transformed into structured narrative elements and then into text (Bal and Veltkamp, 2013).
+Writers must invent narrative worlds and characters while ensuring coherence, which poses unique cognitive challenges: abstract, nonverbal mental images must be transformed into structured narrative elements and then into text (Bal and Veltkamp, 2013).
 The Cognitive Process Model of Writing (Gero et al., 2022; Flower and Hayes, 1981) frames writing as recursive processes of planning, translating, and reviewing.
-In fictional story writing, the transition from planning to translating is especially demanding, as writers move from imaginative constructs to linear verbal representation, imposing a high cognitive load.
+In fictional story writing, the phases from planning to translation are especially demanding, as writers move from imaginative constructs to linear verbal representation, imposing a high cognitive load due to the need for simultaneous translation and structural organization.
 However, visual representations can scaffold this process by externalizing abstract ideas.
-Research shows that picture prompts improve writing coherence (Mukramah et al., 2023), and visual images stimulate creativity in narrative writing (Listyani, 2019).
-In practice, sketches, maps, and diagrams help structure plot, setting, and character relationships during planning, and serve as cognitive anchors during translating to support coherence and consistency (Zhao et al., 2025).
-Dual Coding Theory (Clark and Paivio, 1991) explains these benefits: verbal and imagery systems function separately but also interact, creating richer memory traces when information is encoded in both modalities.
+Research shows that picture prompts improve writing coherence (Mukramah et al., 2023), and visual images stimulate creativity in narrative writing (Listyani, 2019; Han et al., 2024).
+In practice, sketches, maps, and diagrams externalize plot, setting, and character relationships that writers actively manipulate during planning and revision, while also serving as cognitive anchors during translation to maintain coherence and consistency.
+(Zhao et al., 2025).
+Dual Coding Theory (Clark and Paivio, 1991) explains these benefits: verbal and nonverbal systems function separately but also interact, creating richer memory traces when information is encoded in both modalities.
 In fictional writing, visual representations of narrative elements complement verbal planning, making abstract concepts more concrete and retrievable.
 When writers encounter difficulties in translation, visual anchors provide alternative access to imaginative content, reducing cognitive load and enabling more fluid expression (Barsalou, 2008).
-Building on this foundation, a variety of creativity support tools have incorporated visuals into different phases of writing.
-Planning-focused systems such as CCI, Sketchar, and CharacterMeet assist authors in character and world development, often through image-guided backgrounds or conversational refinement of characters (Park et al., 2024; Long et al., 2024; Qin et al., 2024).
+These visual structures are not merely supportive; they function as alternative representational spaces in which writers perform cognitive operations that parallel textual editing and support non-linear narrative leaps (Yan et al., 2023; Singh et al., 2023).
+However, existing creativity-support systems largely leverage visuals in limited ways, focusing on inspiration, reference, or structural overview rather than enabling writers to directly manipulate visual narrative elements and align with text editing.
+Planning-focused systems such as CCI, Sketchar, and CharacterMeet assist authors in character and world development, through image-guided backgrounds or conversational refinement of characters (Park et al., 2024; Ling et al., 2024; Qin et al., 2024).
 Translation-focused tools like ScriptViz and Script2Screen aim to align textual composition with visual referents, either by retrieving reference visuals from movie databases (Rao et al., 2024) or by synchronizing scriptwriting with audiovisual scene creation (Wang et al., 2025c).
-Complexity management systems, for example, WhatIF (Mishra et al., 2025), ClueCart (Wang et al., 2025a), and PlotMap (Wang et al., 2024), help writers maintain structural coherence by visualizing branched narratives, organizing narrative clues hierarchically, or integrating spatial layouts with textual plot structures.
-Despite these advances, visuals in prior systems remain peripheral: they serve primarily as references, scaffolds, or analytic overlays, rather than as co-equal, manipulable materials.
-This treatment constrains multimodal creative expression by limiting writers to either text-first workflows or loosely coupled visual aids (Zhao et al., 2025).
-To bridge this gap, our work introduces a system that treats visuals and text as fully integrated, co-editable artifacts, directly supporting the planning and translating phases of fictional story writing so that imagination can be externalized and refined seamlessly across both modalities.
+Complexity management systems, for example, WhatIF (Mishra et al., 2025), ClueCart (Wang et al., 2025b), and PlotMap (Wang et al., 2024), help writers maintain structural coherence by visualizing branched narratives, organizing narrative clues hierarchically, or integrating spatial layouts with textual plot structures.
+In these systems, writers may look at images to spark ideas, but operations such as cutting, re-ordering, or reframing narratives still have to be performed only in the verbal channel.
+Because the underlying story state is effectively defined only through text, visuals cannot serve as core writing operations such as restructuring events, adjusting focalization, or reorganizing character relationships (Flower and Hayes, 1981; Zhao et al., 2025).
+This separation requires writers to repeatedly convert visually grounded ideas back into verbal form for any narrative change to take effect, thereby increasing cognitive load and undermining many of the well-established benefits of external representations such as diagrams, sketches, and other forms of external cognition (Kirsh, 2010; Singh et al., 2023).
+As a result, images remain outside the recursive planning–translating loop.
+To address this gap, our work treats visual and text representations as synchronized and co-editing materials, allowing writers to manipulate narrative elements across image and text through the same set of operations and thereby more tightly aligning verbal and nonverbal with the cognitive processes of fictional story writing.
 2.2. LLM-powered Multimodality Tools for Creativity in Content Creation
-Recent multimodal creativity tools move beyond linear prompting by enabling direct manipulation of creative elements, helping creators express intentions that language alone cannot capture (Riche et al., 2025; Masson et al., 2024a).
-Powered by advanced LLMs, these systems address fundamental barriers through three complementary mechanisms.
-First, externalizing creative structures.
-Tools like AI-Instruments (Riche et al., 2025) and Brickify (Shi et al., 2025) turn abstract intentions into manipulable interface objects or reusable design tokens, making spatial, stylistic, and compositional relations tangible.
-Second, reducing cognitive load with multimodal input: DrawTalking (Rosenberg et al., 2024) and Code Shaping (Yen et al., 2025) distribute expression across sketch, voice, or annotations, lowering the reliance on precise verbal specification.
-Third, supporting iterative refinement.
-Inkspire (Lin et al., 2025) and AIdeation (Wang et al., 2025b) accelerate variation and exploration, enabling rapid cycles of sketch-to-output or recombination of references.
-However, most current systems remain fragmented: they either support isolated stages (e.g., ideation, editing) or couple modalities only loosely, often within a one-shot prompting paradigm that underutilizes LLMs’ potential for sustained dialogue and refinement (Dang et al., 2023).
-Seeking tighter coupling for fictional story writing, recent systems push integration of multimodal interaction in different ways (Chung et al., 2025; Chung and Kreminski, 2024).
-WorldSmith grounds worldbuilding in text–sketch edits and region-based filling to support users iteratively visualize and modify elements of their fictional world (Dang et al., 2023).
+Recent multimodal creativity tools move beyond linear prompting by enabling direct manipulation of creative elements, helping creators express intentions that language alone cannot capture (Riche et al., 2025; Masson et al., 2025a).
+Powered by LLMs, these systems address fundamental barriers through three complementary mechanisms.
+First, they externalize creative structures and support better intention expression.
+Tools like AI-Instruments (Riche et al., 2025) and Brickify (Shi et al., 2025) transform abstract intentions into manipulable interface objects or reusable visual tokens, rendering otherwise ineffable ideas as visible, persistent, and operable elements.
+Second, in recent multimodal systems, sketches always act as a nonverbal, spatially grounded modality that conveys structure, hierarchy, and relations far more efficiently than language.
+DrawTalking (Rosenberg et al., 2024) combines freehand sketching with spoken narration, enabling natural intention communication.
+Code Shaping (Yen et al., 2025) allows developers to make sketched annotations directly on the code editor to support fuzzy, incremental expression of intent.
+Third, supporting iterative refinement, Inkspire (Lin et al., 2025) and AIdeation (Wang et al., 2025a) accelerate variation and exploration, enabling rapid cycles of sketch-to-output or recombination of references.
+Similarly, the AI Drawing Partner (Davis and Rafner, 2025) combines sketch recognition, text-to-image generation, and direct feedback mechanisms to enable rapid cycles of creative exploration while automatically quantifying the interaction dynamics between human and AI collaborators.
+Seeking tighter coupling for fictional story writing, recent systems push the integration of multimodal interaction in different ways (Chung et al., 2025; Chung and Kreminski, 2024).
+These systems are designed in response to growing evidence that text-only, model-driven workflows cause LLM-assisted stories to converge toward similar narrative structures, limiting exploration, reducing originality, and diminishing writers’ expressive control (Chakrabarty et al., 2024a; Doshi and Hauser, 2024; Li et al., 2024).
+WorldSmith supports layered edits and hierarchical compositions through sketches, making it easier to grow a world piece by piece instead of through single and monolithic prompts (Dang et al., 2023).
 XCreation (Yan et al., 2023) supports cross-modal storybook creation by integrating an interpretable entity-relation graph, improving the usability of the underlying generative structures.
-Toyteller (Chung et al., 2025) maps symbolic motions to character actions.
-Visual Writing allows writers co-edit stories by directly manipulating visual representations of entities, actions, and events alongside text, enabling easier navigation, editing, and exploration of narrative variations (Masson et al., 2024b).
-These systems highlight the transformative potential of LLMs to couple narrative and visual modalities, yet they still lack a unified, synchronized substrate that supports fictional story writing as a continuous, multimodal workflow.
-Our work advances this direction by introducing such a substrate, where LLM-generated texts and visuals act as co-editable inspirational materials for iterative, coherent story development.
-2.3. Instrumental Interaction and Structural Mapping Theory
+Toyteller (Chung et al., 2025) maps symbolic motions to character actions, letting users express rich social and emotional interactions that are often hard to write down explicitly only using text.
+Visual Writing defines an approach where writers edit stories by manipulating visual representations to make the underlying narrative structure more comprehensible and easier to work with than linear text alone (Masson et al., 2025b).
+This line of multimodal research demonstrates that combining language with gestures, sketches, and direct manipulation can offload cognitive work from linear prompting and give creators more expressive, situated channels for specifying and revising intent.
+Building on this line of work, our system introduces a canvas-based interface that integrates images, sketches, and text to support the externalization of mental imagery and the articulation of narrative intent.
+2.3. Instrumental Interaction
 Instrumental Interaction is central to understanding how users control and refine digital systems.
@@
 In computing, this elevates implicit system descriptions into explicit first-class entities.
-In LLM-assisted workflows, this principle appears in modular prompt blocks for structured edits (Yen et al., 2023) and in Textoshop’s reification of abstract text editing commands (e.g., tone adjustment, boolean operations, layers) into direct manipulation tools for text (Masson et al., 2024a).
+In LLM-assisted workflows, this principle appears in modular prompt blocks for structured edits (Yen et al., 2024) and in Textoshop’s reification of abstract image editing commands (e.g., tone adjustment, boolean operations, layers) into direct manipulation tools for text (Masson et al., 2025a).
 A second principle is polymorphism, where the same instrument applies across contexts (Masson, 2023).
@@
 Together, these principles reduce cognitive burden by externalizing interaction histories, making them manipulable, transferable, and extensible.
-In our system, we extend this perspective to AI support for fictional story writing, where imagination must be continuously externalized and revised.
-We design interaction instruments that reify narrative elements into manipulable objects and, through polymorphism, let the same tools operate across LLM-generated text and images.
-This integration of reification and polymorphism enables writers to fluidly adjust multimodal outputs as persistent, tangible entities.
-Structural Mapping Theory (SMT) is central to understanding how users interpret and compare complex representations.
-Gentner (Gentner, 1983) proposed SMT to describe how people form abstractions through analogical alignment of relational structures.
-Rather than relying on surface similarities, SMT highlights the role of alignable differences, which reveal systematic variations and help learners construct new knowledge.
-Building on this foundation, SMT-informed design emphasizes preserving object structure rather than abstracting it away.
-At the collection level, showing complete objects allows users to perceive relationships across the dataset (Yan et al., 2022; Wu et al., 2020).
-At the part level, systems can highlight structural correspondences by surfacing identity matches (literal overlaps) (Glassman et al., 2015), property matches (shared attributes) (Jasim et al., 2022), and conceptual matches (semantic commonalities) (Glassman et al., 2018).
-These correspondences support the iterative process of aligning, re-aligning, and refining mental models.
-Prior visualization and learning systems demonstrate the benefits of making such correspondences explicit, between code solutions (Glassman et al., 2015), textual variants (Glassman et al., 2018), or large language model outputs (Arawjo et al., 2024; Gero et al., 2024), helping users notice critical differences, generalize across examples, and avoid disengagement when facing “walls of text.” SMT thus serves as both a cognitive theory and a design constant for interactive systems that deepen users’ reasoning about collections of related objects.
-In our system, SMT informs the design of synchronized text–image co-editing: textual fragments are explicitly mapped to visual elements based on their shared properties, matched concepts, or references to the same object, so edits in one modality lead to structurally aligned updates in the other, allowing users to compare, reorganize, and refine story elements across modalities.
+In our system, we extend this perspective to the design of multimodal tools for fictional story writing.
+We reify narrative development as a set of instrumental operations spanning text and imagery.
+Through polymorphism, the same operation can be applied to both LLM-generated text and images.
+This combination of reification and polymorphism enables writers to shape multimodal outputs fluidly, aligning with both verbal and nonverbal perceptions.
 3. Formative Study
 Previous research shows that multimodal tools enhance fictional story writing by making abstract concepts tangible, reducing cognitive load, and improving creativity and coherence (Chung et al., 2022; Chung and Kreminski, 2024; Dang et al., 2023; Zhao et al., 2025).
-However, current tools treat images and texts as supplementary rather than integral to the creative process, leaving unclear how writers actually integrate multiple content types.
-To address this gap, we conducted a Wizard-of-Oz co-design study (Dahlbäck et al., 1993) examining how creators use multimodal content (images, text, sketches) when planning and drafting fictional stories (Zhao et al., 2025; Gero et al., 2022; Flower and Hayes, 1981).
+However, current tools treat images as supplementary rather than integral to the creative process, leaving unclear how writers actually integrate multiple content types into a cohesive workflow.
+To address this gap, we conducted a Wizard-of-Oz (WoZ) co-design study (Dahlbäck et al., 1993) examining how creators use multimodal content (images, text, sketches) when planning and drafting fictional stories (Zhao et al., 2025; Gero et al., 2022; Flower and Hayes, 1981).
 Our investigation focused on three questions: (1) Multimodal information use: what types of multimodal content users employ and how they leverage these materials for idea generation; (2) Iteration and integration: how creators refine and combine multimodal artifacts in world-building and narrative development; and (3) Organization of inspirations: how creators organize, connect, and refine dispersed inspirations through multimodal manipulation.
 The WoZ setup simulated AI-assisted visual and textual support while sustaining the impression of an intelligent, interactive system.
+Our system is designed for writers with intermediate to expert writing expertise, rather than novices who are still learning basic narrative composition.
+This target group typically possesses established writing habits and a solid understanding of narrative structure.
+As contemporary writers increasingly incorporate LLMs into their creative workflows (idea generation, style adjustment, etc.) (Chakrabarty et al., 2024b; Lee et al., 2024), we target writers who have hands-on experience using LLMs to assist their writing, even though they may not be experts in multimodal interaction or prompting.
 3.1. Process
-We designed a Wizard-of-Oz (WoZ) co-design study, positioning participants as active co-designers and treating text, sketches, and images as shared design materials (Dahlbäck et al., 1993; Sanders and Stappers, 2008; Vaajakallio and Mattelmäki, 2014).
+We designed a WoZ co-design study, positioning participants as active co-designers and treating text, sketches, and images as shared design materials (Dahlbäck et al., 1993; Sanders and Stappers, 2008; Vaajakallio and Mattelmäki, 2014).
 3.1.1. Participants
-We recruited 10 participants, each with an two or more years of creative writing experience.
+For the formative study, we recruited 10 participants through student organizations by sharing our study announcement in group chats, each with at least two years of experience in creative writing.
 The group included three fictional story writers, three animation scriptwriters, two visual film creators, one new media creator, and one online fiction writer.
 Eight participants held a master’s degree or higher, and two held a bachelor’s degree.
-Experimental Setup Three days prior to the session, participants were instructed to prepare a brief fictional story outline consisting of several sentences that followed one of the narrative structures from The Seven Basic Plots (Booker, 2004), which served as the foundation for subsequent ideation and content development.
-The 90-minute main session took place in Figma(Figma, 2025), which provided a collaborative canvas for free sketching, arranging text–image materials, and treating multimodal inputs as manipulable artifacts.
-Session Process During the session, participants engaged in fictional story co-design through activities including generative prompts, collage, and storyboard-like arrangement while interacting with a simulated AI assistant operated by the researcher (wizard).
+All participants had experience using LLMs to assist in their writing, AI familiarity ranging from casual use (fewer than two days per week, n=5) to daily workflow integration (five or more days per week, n=5).
+Experimental Setup.
+Three days before the session, participants were instructed to prepare a brief fictional story outline consisting of several sentences that followed one of the narrative structures from The Seven Basic Plots (Booker, 2004), which served as the foundation for subsequent ideation and content development.
+Each session concluded with a 30-minute semi-structured interview probing how multimodal materials mediated co-creation, and what interaction patterns and workflows participants desired.
+3.1.2. Wizard-of-Oz System and Session Process
+For the WoZ interface, we utilized the canvas in either Figma (31) or Miro (Miro, 2025) as a collaborative space, which was divided into (1) a user-facing “Text Editor” where participants can put in the outline and edit the story, (2) the “Canvas” where generated images and text, and participants’ notes were, and (3) a hidden “Wizard Control Center” (as shown in Figure 2).
+Participants communicated via voice, text (sticky notes), or hand-drawn sketches while two researchers acted as the “Wizards (system backend)” in real-time to generate outputs, ensuring responsiveness, copying user inputs into separate windows.
+Researchers ran Claude for text generation, ChatGPT (GPT-4o) and Midjourney [^1] All models were accessed via their commercial web interfaces: https://claude.ai (Anthropic, 2024), https://chatgpt.com (38), and https://www.midjourney.com/ (61) respectively, in June 2025., for simulating the visual engine, then pasting the results back onto the “Canvas.” To ensure consistency, the Wizards followed: (1) input the user’s sketch/text as a literal prompt.
+Use the initial outline as contextual information for prompts; (2) do not offer creative suggestions unless explicitly asked; (3) to ensure diversity of output style, one researcher generated both texts and images via ChatGPT, and the other researcher generated images via Midjourney and texts via Claude.
+The two wizards ensured that participants were provided with results that were both timely and diverse.
+Figure 2.
+WoZ System and Study Process (the example of P4) (a) The user edits text in the Text editor; (b) The user writes, sketches, and speaks out their intents; (c)-(d) both wizards paste in the user’s inputs into AI software windows; (e) wizards pasted the generated results back onto the canvas.
+The figure depicts the Wizard-of-Oz study setup and interaction process.
+It shows a text editor where the participant edits story text, a canvas where the participant writes notes, sketches, and expresses intentions, and a hidden wizard control area where researchers input user content into AI tools.
+The generated images and text are then returned to the canvas, illustrating the end-to-end co-design workflow.
+Session Process.
+During the session, participants engaged in fictional story co-design through activities including generative prompts, collage, and storyboard-like arrangement while interacting with the wizards.
 Rather than working toward a fixed output, participants iteratively developed stories of about 300 words while envisioning how the tool itself should behave.
-They freely requested content through voice, text, or hand-drawn sketches, and called upon the assistant for image and text editing as new needs emerged.
-The wizard responded in situ to simulate intelligent system behaviors, encouraging naturalistic exploration.
-To support this process, the experimenter used Claude (Anthropic, 2025) for text generation (drawing on outlines and randomness to stimulate ideation) and Midjourney (Midjourney, 2025) and GPT-4o (OpenAI, 2024) for image generation from prompts, sketches, and editing instructions.
-Each session concluded with a 30-minute semi-structured interview probing how multimodal materials mediated co-creation, and what interaction patterns and workflows participants desired.
 3.2. Formative Study Findings
 3.2.1. Using Multimodal Input to Reify Vague Ideas
-As shown in the Table LABEL:tab:formative_grouped, participants utilized multimodal expressions, including text, sketches, and images, as co-design materials to articulate and negotiate intentions with the wizarded system.
+As shown in the Appendix A.1, participants utilized multimodal expressions, including text, sketches, and images, as co-design materials to articulate and negotiate intentions with the wizarded system.
 Sketches externalized vague intentions and spatial imagination.
-For example, P4 envisioned a scene where clown JoJo appeared on stage and created a sketch showing “JoJo emerging from a giant spinning wheel” with textual annotations describing the intended atmosphere, hoping AI could elaborate narrative details.
-Text functioned as the primary medium for conveying intent, allowing participants to express connections between desired content and existing stories while clearly communicating their intentions.
-Participants frequently used textual annotations to specify story parts or image types for AI generation and guide generation direction.
+For example, P4 envisioned a scene where clown Joko appeared on stage and created a sketch with textual annotations describing the intended atmosphere, hoping AI could elaborate narrative details.
+Sketches were also frequently used to express ideas that participants found difficult to convey through prompts alone (P6, P7, P8), as they were seen as carrying richer layout and spatial information.
+Text functioned as the primary medium for conveying intent, allowing participants to express connections between desired content and existing stories.
+Participants also frequently used textual annotations to specify story parts or image types for AI-generated content and to guide the direction of content generation.
+They also used textual annotations to record the inspirations they received and how those ideas might be used in writing.
 Images communicated style and mood expectations.
-P3 altered an AI-generated picture’s style by supplying a reference image, while P1 noted, “if possible, I want to use a ’supporting image’–a vague reference picture—as a basis, expecting the AI to generate more detailed images derived from it.”
-Table 1. Results of user strategies for manipulating multimodal elements in the WOz co-design study.
-Stage Observed behavior User need / Insight N Interaction Example
-Multi-modal input Text Generated prose should match the established world while allowing the injection of new elements 83 [Uncaptioned image] [Uncaptioned image]
-Multi-modal input Sketches Make spatial relations/composition concrete 8 [Uncaptioned image] [Uncaptioned image]
-Planning LLM-Generated images Precisely locate inspiration from images 90 [Uncaptioned image] [Uncaptioned image]
-Planning LLM-Generated text Filter usable bits from many generations and reuse them 90 [Uncaptioned image] [Uncaptioned image]
-Planning/Translating Text notes / annotations Externalize temporary ideas for re-generation or final writing 30 [Uncaptioned image] [Uncaptioned image]
-Planning/Translating Collage elements Recompose fragments from cross-image to form new scenes 12 [Uncaptioned image] [Uncaptioned image]
-Planning/Translating Link elements Structure relationships between character / scene / object to connect the plot 40 [Uncaptioned image] [Uncaptioned image]
-Translating Reconfigure elements in canvas Reorder scattered ideas by role / time / space into place for global understanding 4 [Uncaptioned image] [Uncaptioned image]
-Translating Text (integration) Consolidate AI-generated text and image-inspired content into the draft 10 [Uncaptioned image] [Uncaptioned image]
-3.2.2. Text–Image Interplay as Complementary Design Moves
-Participants most often sought AI elaboration on characters, objects, and scenes, expressing the need for assistance with character design refinement, setting depictions, or object visualization.
-Combining image outputs with textual descriptions helped participants enrich their limited knowledge.
-For example, P7 requested AI-generated designs of an ancient Chinese poison bottle as a narrative element, noting her vague understanding of the concept.
-These findings highlight the value of systems that accept heterogeneous multimodal input and help co-designers transform nascent ideas into concrete narrative materials.
-Participants perceived text and images as distinct yet complementary in their creative writing.
-We observed an oscillation between abstraction (text) and concreteness (images) as a recurring co-design pattern, which reveals design opportunities for multimodal authoring tools.
+P3 altered an AI-generated picture’s style by supplying a reference image, while P1 noted, “If possible, I want to use a ‘supporting image’—a vague reference picture—as a basis, expecting the AI to generate more detailed images derived from it.” In addition, sketches were often paired with images or text to further articulate the intentions participants held in mind (P4, P7).
+These findings highlighted the value of systems that accept different modality input and help co-designers transform nascent ideas into concrete narrative materials.
+3.2.2. Image-Text Interplay as Complementary Design Moves
+Participants perceived text and images as distinct yet complementary, and we observed them often switching between abstraction (text) and concreteness (images) as a recurring co-design pattern.
 Text as open imagination.
-Participants described text as a “blank canvas” for associative and structural thinking.
-P4 noted, “text allows me to imagine many things in my mind,” and P1 emphasized its “infinite imagination on a blank page.” P8 highlighted that text helped set up narrative structure before layering in visuals.
-This suggests systems should treat text as a flexible space for ideation and intent communication, where ambiguity can be preserved rather than prematurely resolved.
+Participants described text as a “blank canvas” for boundless imagination.
+P4 noted, “Text allows me to imagine many things in my mind,” and P1 emphasized text as “infinite imagination on a blank page.” P8 highlighted that text helped set up the narrative structure before layering in visuals.
+This suggests systems should treat text as a flexible space for ideation and intent communication, where ambiguity can be preserved rather than early resolved.
 Images as concreteness, inspiration, and feedback.
 Images grounded abstract ideas, while their randomness often sparked unexpected inspiration.
-P1 explained: “The randomness in AI-generated images goes beyond what I want or can express; it helps me imagine the next step of the story.” Similarly, P4 refined Lily’s behavior based on an unexpected visual detail, and P3 used images as feedback for progressive refinement.
-P10 likened it to a “look-and-write exercise” that scaffolds scene construction.
+P1 explained: “The randomness in AI-generated images goes beyond what I want or can express; it helps me imagine the next step of the story.” Similarly, P10 regarded referring to images as a “look-and-write exercise” that scaffolds scene construction.
 These accounts highlight the value of image outputs not just as illustrations but as provocations.
-Designers can leverage it through image-based iteration, selection, and reinterpretation.
 Complementary interplay.
 Participants emphasized that neither modality sufficed alone: images “set the vibe,” while text reframed meaning.
-P8 noted, “Images can serve as references for appearance when I don’t have many ideas, while text quickly triggers associations.” We observed participants iteratively moving between text for open-ended imagination and images for concrete grounding, forming a cycle of divergence and convergence.
-This interplay suggests design opportunities for systems that orchestrate smooth transitions between modalities, enabling users to fluidly oscillate between abstract exploration and concrete elaboration.
+P8 noted, “Images can serve as references for appearance when I don’t have many ideas, while text quickly triggers associations.” We observed participants iteratively moving between text for open-ended imagination and images for concrete grounding, forming a cycle of divergence and convergence (P4, P5, P8, P9, P10).
+Four participants (P1, P2, P5, P7) also expressed a desire for text and image changes to be synchronized, so they would not need to constantly cross-check and compare updates across modalities.
+In addition, the concurrent presentation of text and images further facilitates narrative expression.
+As P7 explained, “ When text and images appear together, I can describe the scene while referencing the image and directly adopt AI-generated text when it proves useful.
+This also allows inspiration to emerge.”
+This interplay suggested design opportunities for systems that incorporated smooth transitions between text and images while aligning these two modalities, enabling participants to fluidly move between abstract exploration and concrete elaboration
 3.2.3. Direct Manipulation of Multimodal Artifacts
 Participants expressed a strong interest in treating text and images as manipulable, recombinable design materials.
-Two recurring practices pointed to design needs for more fluid multimodal manipulation.
+Two recurring practices pointed to design needs for more direct multimodal manipulation.
 Collaging and Recombination.
@@
 P9 envisioned combining “the house from the first AI-generated images with the street from the second picture” to construct scenes, while P7 highlighted that “randomly combining characters and scenes” could inspire unexpected connections when accompanied by textual descriptions.
-As she explained, “if I can see an AI-generated image of my protagonist in one of the scenes, it helps me better imagine potential connections between elements that might otherwise seem unrelated.” Such practices illustrate the potential of collage and recombination as creative strategies.
+She also expressed that visualizing hard-to-imagine scenes makes it easier to write richer narrative descriptions.
+Such practices illustrated the potential of collage and recombination as creative strategies.
 Granular Editing and Annotation.
 Beyond recombination, participants desired fine-grained control over outputs.
-Sticky notes captured details for iteration and served as prompts for later development.
+They use sticky notes to capture details for iteration and to serve as prompts for later development.
 Participants also wanted more localized operations, such as regenerating specific regions (P1, P3), extracting and reusing circled image elements (P5), or annotating character personas for refinement (P2).
-They left narrative prompts for later translation, e.g., P2’s note “Ending could be related to why this postman job even exists.” These behaviors emphasize annotation as both a vehicle for iteration and a bridge to subsequent writing.
-Together, these findings suggest systems should enable flexible recombination, localized editing, and traceable annotations to help creators turn multimodal inspiration into actionable writing material.
+These behaviors emphasized editing and annotation as both vehicles for iteration and a bridge to subsequent writing.
+Together, these findings suggested systems should enable flexible recombination, localized editing, and traceable annotations to help creators iteratively refine narrative materials.
 3.2.4. From Fragmented Inspirations to Coherent Storylines
-While participants often highlighted text or circled inspiring image details, organizing these dispersed fragments into coherent narratives was a persistent challenge.
-As P2 noted, “everything quickly became too messy on the canvas”, and P4 likened fragments on the canvas to “many cards that required connections,” where narrative coherence depends on linking passages, characters, and settings from scattered parts.
+While participants often highlighted text or circled inspiring image details, organizing these dispersed fragments into coherent narratives was a persistent challenge in the planning phase.
+As P2 noted, “Everything quickly became too messy on the canvas,” and P4 likened fragments on the canvas to “many cards that required connections,” where narrative coherence depends on linking passages, characters, and settings from scattered parts.
 Furthermore, P6 wished for mind map–like tools to scaffold this process.
-Participants also requested clustering notes, surfacing latent relations (e.g., by character/object/setting), and consolidating materials into reusable “setting cards” for ensuring cross-chapter consistency and avoiding logic conflicts (P4, P7).
-These challenges point to opportunities for systems that transform fragmented inspirations into structured storylines by supporting clustering, relation mapping, and the creation of reusable narrative units that preserve coherence across iterations.
+Participants also requested clustering notes, surfacing latent relations (e.g., by character/object/setting), and consolidating materials into reusable “setting cards” to ensure cross-chapter consistency and avoid logic conflicts (P4, P7).
+These challenges pointed to opportunities for systems that transform fragmented inspirations into structured storylines by supporting clustering, relation mapping, and the creation of reusable narrative units that preserve coherence across iterations.
 3.3. Design Goals
 Drawing on insights from our WoZ co-design study, prior work on multimodal LLM tools, and theories of Structural Mapping and Instrumental Interaction (Section 2.3), we identify four design goals for a multimodal content creation interface that supports the planning and translating phase in fictional story writing (Gero et al., 2022; Flower and Hayes, 1981; Zhao et al., 2025).
-  • DG1: Reifying Multimodal Intentions into Manipulable Artifacts.
-Informed by findings in Section 3.2.1, our system should reify multimodal inputs (text, sketches, images) into first-class artifacts that can be annotated, rearranged, and recombined.
-Drawing on Instrumental Interaction, such reification should preserve fleeting intentions as reusable materials for later iteration; following SMT, these artifacts should function as alignable units that support comparison, linkage, and cross-modal reorganization.
+Figure 3. Vistoria’s interface: a left Text Editor, a central Cluster panel (can be collapsed when not used), and a right free-form Canvas.
+The figure presents Vistoria’s interface layout.
+It consists of a left-side Text Editor displaying the current story draft, a central Cluster panel that aggregates highlighted narrative elements, and a right-side free-form Canvas for sketching, placing images, writing text, and performing image–text co-editing operations.
+  • DG1: Supporting the Expression of Ideas through Multimodality.
+Grounded in the findings in Section 3.2.1, our system should provide multimodal mechanisms combining sketches, text, and images to capture early intention, imprecise expressions, and help transform them into concrete narrative materials for further iterative editing or re-organizing.
   • DG2: Aligning Text and Images for Iterative Creative Exploration.
-Informed by findings in Section 3.2.2, our system should enable fluid cross-modal iteration: textual edits can be re-visualized, image refinements can inform descriptions, and alternative versions should make differences explicit.
-Grounded in Dual Coding Theory and SMT, this alignment should support cycles of divergence and convergence by allowing users to oscillate between abstraction (text) and concreteness (images).
+Informed by findings in Section 3.2.2, our system should enable fluid cross-modal iteration: textual edits can be re-visualized, and image refinements can inform text descriptions.
+Grounded in Dual Coding Theory, text and image updates should also be synchronized to more effectively align verbal and nonverbal perception.
   • DG3: Enabling Polymorphic Cross-Modal Manipulation.
-Informed by findings in Section 3.2.3, our system should support direct manipulation interactions that apply uniformly across modalities, guided by Instrumental Interaction’s principle of polymorphism.
-This uniformity should reduce switching costs and empower writers to fluidly manipulate and edit textual and visual fragments while maintaining narrative coherence.
+Informed by findings in Section 3.2.3, our system should support direct manipulation interactions for both text and images.
+Guided by Instrumental Interaction’s principle of polymorphism, we should design the same instrument for cross-modal editing to reduce switching costs and enable writers to manipulate textual and visual fragments while maintaining narrative coherence.
   • DG4: Organizing and Reusing Fragments into Coherent Narratives.
-Informed by findings in Section 3.2.4, our system should cluster and organize fragments and fleeting ideations, surface hidden connections, and consolidate dispersed inspirations into coherent, evolving narratives.
-It should further support the reuse of narrative units (e.g., characters, settings, objects) to ensure consistency and foster pathways for story progression.
+Informed by the findings in Section 3.2.4, our system should support clustering and organizing fragments and fleeting ideations during the exploratory phase, surface latent connections, and consolidate dispersed inspirations into coherent, evolving narrative structures that support translation into final writing.
 4. Vistoria System
@@
 As shown in Figure 3, the interface comprises three primary components: a left Text Editor, a central collapsible Cluster panel, and a right Canvas interface.
-The text editor houses the current story draft, serving as contextual information for content generation.
-The right canvas supports freeform sketching, text input, and comprehensive editing tools.
-The central cluster panel aggregates highlights and annotations from canvas content, organizing them by characters, objects, and scenes for reference and summary.
-Together, these components create a workflow where narrative text, visual artifacts, and organizational structures remain continuously aligned, enabling writers to externalize, manipulate, and recombine ideas and materials across modalities.
-4.1. Key Features
+The Text Editor displays the current story draft, serving as contextual information for content generation.
+The right Canvas supports freeform sketching, text input, and image-text generation and editing tools.
+The central Cluster panel aggregates highlights and annotations from Canvas, displaying related plots, settings, and descriptions of each highlighted element for easy reference and overview.
+Figure 4. Vistoria analyzes selected text, sketches, images, and broader story context to generate image–text cards that externalize narrative development directions.
+The figure shows how Vistoria generates image–text cards from multimodal input.
+Selected regions containing text, sketches, and images on the Canvas are analyzed together with the broader story context.
+The system then produces image–text cards that externalize potential narrative development directions.
 Figure 5.
-Vistoria’s interface: a left text editor, a central Cluster panel(can be collapsed when not used), and a right free-form Canvas.
-Cards (image-text pairs with main object keywords) are generated from user intentions and serve as the basic units; two modes—Exact Craft and Creative Spark—support precise realization vs.
-exploratory variation.
-4.1.1. Reifying Intention through Multimodal Generation
-Writers often externalize early intentions through sketches, short notes, or reference images, but these expressions are typically fleeting and difficult to refine into coherent narrative materials.
-Supporting the preservation and refinement of such multimodal intentions is therefore essential for sustaining ideation (DG1).
-To address this, Vistoria reifies multimodal inputs into persistent Cards that bundle an image, a story fragment, and extracted narrative objects (characters, objects, scenes).
-Drawing on the principle of Reuse, Vistoria reifies sketches and text from ephemeral marks into manipulable and reusable artifacts that remain on the canvas and can be invoked again as elements for regeneration, enabling writers to carry forward and iteratively develop their creative intentions.
-At the same time, we treat cards as alignable units: text and their corresponding images are designed to carry the same underlying meaning, with images serving as visualizations of the narrative described in text.
-Vistoria further balances precision and exploration by offering two complementary generation modes (Figure 3).
-In Exact Craft mode, single cards closely adhere to the author’s expressed intention to concretize specific ideas.
-In Creative Spark mode, three cards were generate represent diverse options based on writer’s intention, the system deliberately introduces variation around characters, settings, or objects, providing alternative prompts that can inspire new directions.
-Figure 3.
-Vistoria balances precision and exploration through two complementary generation modes: in Exact Craft mode, a single card closely follows the author’s intention to concretize ideas; in Creative Spark mode, three cards are generated with deliberate variations in characters, settings, or objects, offering diverse options to inspire new directions.
-Figure 4.
-A set of visual instruments for image-text co-editing to enhance planning and translating of fictional story writing: (a) Lasso selects regions for coupled image–text edits.
-(b) Writers extract elements and compose across cards to discover new narrative directions, using collage alongside other instruments to recombine sketches, images, and text on the canvas.
+A set of instrumental operations for image-text co-editing to enhance planning and translating of fictional story writing: (a) Lasso selects regions for coupled image–text edits.
+(b) Collage enables writers to extract elements and compose across cards to discover new narrative directions.
 (c) Perspective Shift changes an image’s viewpoint and automatically regenerates the story’s point of view (first/third/second person).
 (d) Filters align visual style and textual tone (e.g., melancholic/dreamy) by jointly altering image effects and rewriting prose.
-4.1.2. Synchronized Image–Text Co-Editing through Visual Instruments
-Fictional story writing benefits from fluid movement between abstract textual reasoning and concrete visual imagination, yet existing tools separate these modalities and disrupt creative flow (DG2, DG3).
-To address this, we introduce a set of visual instruments (Figure 4) designed around three principles: (1) Reification, which draws on familiar operations from image editing (e.g., lassoing, collaging, filtering) and transforms them into persistent, manipulable operations that apply equally to text and image elements (Swerzenski, 2021; Masson et al., 2024a); (2) Polymorphism, which ensures the same operations apply uniformly across text and images; and (3) Structural Mapping, which synchronizes changes so that visual and textual representations carry the same narrative meaning.
-Together, these principles reduce modality switching costs while maintaining coherence between story segments and imagery.
+The figure illustrates four instrumental operations for synchronized image–text co-editing.
+The Lasso operation selects localized regions for focused edits, the Collage operation combines extracted elements across cards, the Filter operation aligns visual style with textual tone, and the Perspective-Shift operation changes visual viewpoint and narrative voice together.
+Each operation affects images and text simultaneously.
+4.1. Key Features
+4.1.1. Reifying Intention through Multimodal Generation
+The system enables writers to externalize early, vague ideas using multimodal inputs (DG1).
+To support this, Vistoria transforms multimodal input (sketches, text, and images) into cards that pair a representative image with its corresponding narrative segment.
+Writers select a region of interest, capturing all multimodal elements within that area.
+Vistoria then infers from these multimodal elements and broader story context to generate image–text cards that reflect the writer’s intention for narrative direction (Figure 4).
+These cards serve a dual purpose: they transform ephemeral multimodal input into persistent, regenerable artifacts on the Canvas, while also functioning as semantically aligned units where image and text convey the same underlying meaning.
+Vistoria further balances precision and exploration by offering two complementary generation modes.
+In Exact Craft mode, single cards closely adhere to the author’s expressed intention to concretize specific ideas.
+In Creative Spark mode, three cards are generated to represent diverse options based on the writer’s intention.
+The system deliberately introduces variation around characters, settings, or objects, providing alternative prompts that can inspire new directions.
+Figure 6.
+Writers highlight objects and text segments on cards (b-1); the Cluster panel aggregates these by character/object/scene and can auto-summarize settings/plot/description about a certain object to guide final writing (b-2); Clicking on a specific image reveals the corresponding highlights and comments from earlier phases left on Canvas (b-3).
+The figure shows how highlighted text segments and objects on cards are aggregated in the Cluster panel.
+Narrative elements are grouped by character, object, or scene, and the system generates structured summaries of plots, settings, or descriptions.
+Selecting an item reveals its associated images, text highlights, and comments to support coherent story construction.
+4.1.2. Synchronized Image–Text Co-Editing through Instrumental Operations
+Fictional story writing benefits from fluid movement between abstract textual reasoning and concrete visual imagination.
+The system should tightly align text and images so that edits in one modality fluidly inform the other (DG2, DG3).
+To address this, we introduce a set of Instrumental Operations (Figure 5) designed around three principles: (1) Reification (instrumental interaction), which draws on familiar image-editing operations to make abstract image–text co-editing actions more concrete and manipulable; (2) Polymorphism (instrumental interaction), designing a set of instrumental operations (Lasso, Collage, Filter, Perspective Shift) which ensure the same operations apply uniformly across text and images to lower switching cost; and (3) Dual Coding Theory, which indicates that verbal and nonverbal changes should be aligned to maintain coherence across cognitive channels.
 Lasso.
-The lasso instrument exemplifies reification by turning the abstract action of “focusing on part of a story” into a manipulable unit: selecting a region in either an image or a fragment of text generates a new card focusing on the selected part with enriched narrative and visual details.
+The Lasso instrument exemplifies reification by turning the abstract action of “focusing on part of a story” into a manipulable unit: selecting a region in either an image or a fragment of text triggers the generation of a new card focusing on the selected part with enriched narrative and visual details.
 Through polymorphism, the same selection logic applies across modalities—whether circling a visual detail or isolating a text segment—providing a consistent interaction pattern.
@@
 The extracted portion of text is used to regenerate the corresponding image.
-Structural mapping ensures that the lassoed image region and the expanded story fragment correspond to one another, aligning visual and textual perspectives within the same narrative unit (Figure 4 (a) ).
+The lassoed image region and the expanded story fragment correspond to one another, aligning visual and textual perspectives within the same narrative unit (Figure 5 (a)).
 Collage.
 The Collage instrument reifies the abstract act of “recombining inspirations” into a tangible manipulation: fragments of images, sketches, or text can be directly composed within a collage frame to form a new card.
-Through polymorphism, the same cut–paste–combine logic applies uniformly across modalities—an image region, a text excerpt, or a sketch element can all be treated as compositional materials for intention-based generation.
-Guided by structural mapping, the system interprets the spatial arrangement of these multimodal pieces as narrative intent, generating a card where textual descriptions and visual depictions are aligned.
+The same cut–paste–combine logic applies uniformly across modalities—an image region, a text excerpt, or a sketch element can all be treated as compositional materials for intention-based generation.
+The system interprets the spatial arrangement of these multimodal pieces as narrative intent, generating a card where textual descriptions and visual depictions are aligned.
 For instance, merging two character fragments not only produces a combined image but also generates a new story segment situating them together, ensuring that narrative and imagery evolve in sync (Figure 5 (b)).
@@
 Stylistic coherence is critical in fictional story writing, as consistent affective and aesthetic cues sustain narrative transportation (Green and Brock, 2000), activate readers’ interpretive schemas (Arbib, 1992), and enhance the emotional resonance of literariness (Miall and Kuiken, 1999).
-In Vistoria, the filter instrument reifies this abstraction into a concrete tool: applying a “melancholic” or “dreamy” filter adjusts both the style of the image and the tone of the accompanying prose (Appendix Table 4).
-Through polymorphism and structural mapping, the same filter operation works seamlessly across modalities, leveraging the correspondence between visual style in images and emotional tone in text to act on both simultaneously.
+In Vistoria, the Filter instrument reifies this abstraction into a concrete tool: applying a “melancholic” or “dreamy” filter adjusts the visual style and rewrites the accompanying prose to match the emotional tone (Appendix Table 3).
+Through polymorphism, the same filter operation works seamlessly across modalities, leveraging the correspondence between visual style in images and emotional tone in text to simultaneously act on both.
 By making intangible stylistic intentions manipulable and synchronized, filters expand expressive possibilities while maintaining narrative immersion (Figure 5 (c)).
 Perspective Shift.
-Fictional story writing often utilizes perspective shifts, as narratology highlights that changes in voice and focalization fundamentally reshape how events and characters are perceived (Genette, 1980).
+Fictional story writing often utilizes perspective shift, and narratology highlights that changes in voice and focalization fundamentally reshape how events and characters are perceived (Genette, 1980).
 Cognitive poetics further shows that such shifts alter readers’ empathy and immersion.
@@
 Through polymorphism, this instrument applies consistently across modalities, altering either an image or its accompanying text triggers a corresponding adjustment in the other.
-Structural mapping ensures that the focalization shift carries the same meaning across text and image: a new camera angle in the image corresponds to a new narrative voice in the text, allowing writers to explore empathy, distance, and structural awareness in a synchronized manner (Figure 4 (d)).
+The shift carries the same meaning across text and image: a new camera angle in the image corresponds to a new narrative voice in the text, allowing writers to explore empathy, distance, and awareness in a synchronized manner (Figure 5 (d)).
 4.1.3. Highlight Elements and Cluster.
@@
 This process is centered in the Cluster panel (Figure 6), which turns fragmented inputs into organized knowledge assets.
-Figure 5.
-Writers highlight objects and text segments on cards(b-1); the Cluster panel aggregates these by character/object/scene and can auto-summarize settings/plot/description about a certain object to guide final writing(b-2); Clicking on a specific image reveals the corresponding highlights and comments from earlier phases left on canvas(b-3).
 On the Canvas, writers can highlight textual segments, edit stories directly, and add inline comments noting potential uses in later drafting.
 Story objects, such as characters, settings, or scenes, are represented as editable keywords that can be highlighted by themselves.
-The system automatically binds each highlighted object to its associated text, consolidating references across multiple cards.
+The system automatically links each highlighted object to its associated text, consolidating references across multiple cards.
 The Cluster panel then aggregates all highlighted objects into an organized overview of evolving narrative elements.
-This eliminates the need to manually scan scattered cards and provides writers with a dynamically updated, object-centered workspace.
-Selecting an object reveals its complete associated materials, such as linked images, highlighted text segments, and comments, which creates a multimodal, context-rich reference for downstream writing.
+This eliminates the need for manual scanning of scattered cards and provides writers with a dynamically updated, object-centered workspace.
+Selecting an object reveals its complete set of associated materials, including linked images, highlighted text segments, and comments, which creates a multimodal, context-rich reference for downstream writing.
 Beyond simple aggregation, the panel supports higher-level knowledge construction through its summary feature.
 When this feature is invoked, the system generates structured summaries of settings, descriptions, and plot elements derived from highlights and comments.
-These summaries distill fragmented annotations into narrative building blocks (Burroway et al., 2022), enabling writers to iteratively scaffold coherent storylines from previously disjointed ideas.
+These summaries distill fragmented annotations into narrative building blocks (Burroway et al., 2019), enabling writers to iteratively scaffold coherent storylines from previously disjointed ideas.
 4.2. Implementation
-System Architecture.
 We adopted a decoupled front-end/back-end architecture.
-The React front-end enables efficient rendering for complex interactive interfaces, while the Flask back-end flexibly handles model calls with minimal overhead.
+The React [^2] https://react.dev/ front-end enables efficient rendering for complex interactive interfaces, while the Flask [^3] https://flask.palletsprojects.com/ back-end flexibly handles model calls.
 Axios manages asynchronous communication between layers.
-The canvas module uses React-Flow for dynamic node-edge relationships and integrates “path to SVG” for natural interactions like freehand drawing and lasso selection.
-A DOM-based screenshot function ensures visual consistency in exported images.
-The text editor integrates React-Quill for rich text editing with lightweight customizations.
-Global state (text content, canvas nodes, selections) is centrally managed via Zustand to ensure consistency across modules.
-Per-session persistence relies on sessionStorage, which preserves data during the study session and clears on tab close to protect privacy.
-Back-end Multi-Agent flow.
-The back-end bridges front-end actions with language and vision models through multi-agent orchestration.
-Every canvas action (e.g., selection, lasso, collage, inline text) is rasterized into screenshots paired with local context (outline, prior cards, global theme).
-A text agent (o4-mini) examines screenshots using templated prompts from a prompt library to extract user intent, recover sketch layouts, and write text consistent with surrounding outlines.
-The agent returns structured JSON (story text, intent, layout hints, semantic tags) that downstream visual agents use for synthesis and alignment.
-Image Generation and Editing.
-An image agent invokes Flux Kontext Pro with consolidated prompts that fuse screenshot-derived intent, story fragments, and style controls.
-Sketches serve as structural scaffolds respecting composition while avoiding duplication; reference images act as content anchors, encouraging variation over copying.
-When neither is present, Flux operates in free mode guided by text-agent intent and style.For editing, front-end instruments (e.g., perspective shifts, filters, collage) are translated into prompt deltas through GPT-4o’s image pathway, yielding coupled visual updates.
-In parallel, the text agent revises tone or viewpoint for synchronized multimodal transformation.
-All artifacts are persisted as cards (image, story, objects keywords), enabling highlight-and-cluster operations that help reassemble fragments into coherent narratives.
+4.2.1. Front-end.
+The front-end consists of three main modules: the Canvas, the Cluster, and the Text Editor.
+Zustand [^4] https://github.com/pmndrs/zustand centrally manages the global state (including canvas nodes, cluster selections, and text content) to ensure consistency across all modules.
+To protect privacy during user studies, per-session data is stored in sessionStorage and automatically clears when the tab closes, while users can manually export Canvas nodes and text content via the top toolbar.
+Canvas Module.
+The right-side Canvas module consists of four distinct layers: (1).
+Node Interaction Layer (Bottom): This layer uses React-Flow [^5] https://reactflow.dev/ to maintain a dynamic node-edge graph.
+Node types include card, collage, text, sketch, handwriting, and image.
+All nodes share basic properties, such as ID, type, and coordinates, for edge linking, but the internal data structures of those node types vary for rendering.
+For example, the “card” node includes features such as image modification tools, image lasso selection, image highlighting, object manipulation, and basic information display.
+(2).
+Pen-based Input Capture Layer: This layer supports natural interactions like freehand sketching, writing, and lasso selection.
+It uses the perfect-freehand library [^6] https://github.com/steveruizok/perfect-freehand to smooth captured points and convert them into Scalable Vector Graphics (SVGs).
+Each captured SVG stroke is added to the graph as a new node.
+(3).
+Generation Selection Layer: A Document Object Model (DOM) based screenshot function ensures visual and positional consistency.
+After clicking generate, the front-end sends both the screenshot and structured data of the nodes inside the selected area to the backend, adds a new node to the graph, and awaits the returned data.
+(4).
+Tool Layer (Top): This is the most visible layer.
+It contains operation tools and Canvas control tools.
+Cluster Module.
+The middle module displays selected information, including objects, images, text, and annotations.
+Users filter this content from the node-edge graph by selecting specific objects.
+A button allows users to expand or collapse this entire area.
+Text Editor Module.
+The left-side Text Editor uses the React-Quill [^7] https://github.com/zenoamaro/react-quill component.
+It provides lightweight rich-text editing capabilities designed to align with the narrative structure.
+4.2.2. Back-End Multi-Agent Flow.
+The backend is organized as a multi-agent pipeline, where each prompt-specialized LLM agent is organized sequentially.
+Rather than relying on a single model, the system decomposes the workflow into three cooperating functional agents:
+Narrative Construction Agent.
+This agent takes multimodal context—such as Canvas screenshots which incorporate all information (sketches, text inputs, and images) contained within the user-selected region to produce structured textual outputs—including user intentions and story segments.
+Its prompts enforce consistency with contextual information such as the existing story content in the Text Editor and the global stylistic constraints.
+In essence, the agent transforms the user’s multimodal inputs into coherent narrative storylines.
+During implementation, the GPT-o4-mini [^8] We used OpenAI’s o4-mini model via the OpenAI API (model ID o4-mini) in August 2025 (OpenAI, 2025).
+is used to process Canvas screenshots and contextual information to infer user intentions because of its rapid inference speed and strong reasoning ability.
+GPT-4o [^9] We used OpenAI’s GPT-4o model via the OpenAI API (model ID gpt-4o) in August 2025 (OpenAI, 2024).
+is used to refine these inferred intentions from o4-mini and contextual information into polished, coherent story segments.
+Prompts for precise description generation of o4-mini are shown in Appendix A.5).
+When applying instrumental operations for editing, the story segment from the previous card is also read in prompts and modified accordingly.
+Visual Synthesis Agent.
+Using the narrative produced by the Narrative Construction Agent as prompts, this agent supports image generation using either the GPT-4o API or the FLUX diffusion model [^10] We used the FLUX.1 diffusion model via the Black Forest Labs API in August 2025 (Black Forest Labs, 2024)..
+When reference images or screenshots are available—such as during instrumental-operation edits that modify images or when multimodal inputs include a base image or sketches—GPT-4o is used for image generation, leveraging its strong capabilities in image understanding and re-generating based on base images.
+In all other cases, the FLUX diffusion model is used due to its faster generation speed.
+Memory Agent.
+The backend maintains a persistent, globally accessible memory of all previously generated image-text pairs.
+This agent coordinates read/write operations to this store, enabling multi-turn reuse, cross-scene integration, and contextual continuity.
+When the user later requests to merge scenes, change details, or perform local edits, this agent retrieves relevant image-text pairs and passes them back to the earlier agents to re-initiate the pipeline.
 5. User Study
-To evaluate the usability, effectivness and usefulness of Vistoria, we conduct a lab user study with 12 participants. The study aims to answer the following research questions:
-  • RQ1: What interaction patterns and strategies emerge when participants use a image-text co-editing process?
-  • RQ2: How does image-text co-editing facilitate and shape the cognitive process of story creation, in terms of ideation, iteration, and refinement?
+To evaluate the usability of Vistoria and to understand how these multimodal interactions design support creativity, we conducted an exploratory lab user study with 12 participants.
+We structure our user study around two complementary components:
+(1) A usability evaluation of Vistoria aiming to answer the following research questions:
+RQ1: How useful are the multimodal co-editing functions, and in what ways do they influence participants’ workflows?
+RQ2: How does multimodal image–text co-editing affect participants’ workload compared to a text-only baseline?
+(2) A pilot study examining the creativity support provided by Vistoria, focusing on:
+RQ3: How does multimodal image–text co-editing influence participants’ ideation and the development of fictional stories?
+RQ4: How does multimodal image–text co-editing influence participants’ sense of agency and ownership?
+RQ1 evaluates the usefulness of these multimodality functions and the strategies participants adopted in relation to DG3 (polymorphic instrumental operations).
+RQ3 explores how DG1 (reifying intentions through multimodal input) and DG2 (image-text alignment) support creative ideation and narrative development.
+The design of the cluster panel corresponding to DG4 functions mainly as auxiliary support and is not central to our user evaluation of multimodal interaction.
+Note that we used the sense of ownership to refer to the writer’s “sense of possession” over the resulting narratives within the system, even the AI-generated artifacts (He et al., 2025; Lee et al., 2024; Draxler et al., 2024b; Weber et al., 2024).
+The declaration of the sense of agency, on the other hand, refers to the writer’s awareness of “initiating, executing, and controlling” key actions in the writing and artifact editing process (Moruzzi, 2022; Legaspi et al., 2019; Lee et al., 2024).
 5.1. Participants
-We recruited 12 participants (6 males, 6 females, aged 21–32, M=25.5), all with prior creative writing experience.
-Most held Bachelor’s or Master’s degrees, with backgrounds spanning science, arts, design, or communication.
-Their creative practices included fiction writing, screenwriting, songwriting, advertising, research, and philosophy.
-Participants’ creative writing experience ranged from under one year (n=5) to over seven years (n=1), with others reporting 1–3 years (n=2) or 4–7 years (n=3).
-All participants were familiar with large language models (e.g., ChatGPT, Gemini, Claude) and had used them for idea generation, editing, descriptive support, content expansion, worldbuilding, and style imitation.
-This diversity reflects our design goal of supporting not only expert writers but also a broader range of users.
-Each participant received \$40 USD compensation.
+As the system targets intermediate to expert participants who understand narrative structure and already use LLMs in their creative workflows, we recruited 12 participants (6 males, 6 females, aged 21–32, M=25.5), all with prior creative writing experience.
+Participants were also recruited through student organizations by sharing our study announcement in their group chats.
+All held Bachelor’s degrees, with backgrounds spanning science, arts, design, or communication.
+Their creative practices included fiction writing, screenwriting, songwriting, advertising, and media production.
+Participants’ creative writing experience ranged from under one year (n=5) to over seven years (n=1), with others reporting 1–3 years (n=3) or 4–7 years (n=3).
+All participants were familiar with LLMs (e.g., ChatGPT, Gemini, Claude) and had used them for idea generation, editing, descriptive support, content expansion, world-building, and style imitation.
+Among the 12 participants, 6 use LLMs daily, while the remaining 6 are evenly split across several times a week, occasionally, and rarely (2 in each).
+Each participant received a \$40 USD compensation after finishing the experiment.
 5.2. Procedure
@@
 To support sketch input, we provided an external tablet (iPad) for freehand drawing on the canvas.
-The baseline condition presented a side-by-side interface with a text editor and GPT-4o conversational panel, enabling both manual editing and LLM-assisted text/image generation.
+The baseline condition presented a side-by-side interface with a text editor and GPT-4o (38) conversational panel, enabling both manual editing and LLM-assisted text/image generation.
 Participants completed two story-writing tasks (Appendix A.3), each extending a given story beginning into a 300–500 word draft.
@@
 After informed consent and a demographic survey, participants were introduced to Vistoria through a written guide and tutorial video, followed by a short hands-on exploration (15 minutes).
-In each condition, participants first focused on worldbuilding and idea exploration (20 min) and then on refining and improving the story (20 min).
+In each condition, participants first focused on world-building and idea exploration (20 min) and then on refining and improving the story (20 min).
 We divided the writing task into two phases (exploration and refinement) to prevent participants from prematurely committing to a single storyline and to reduce fixation, thereby encouraging broader ideation before focused improvement (Suh et al., 2024).
-After each condition, participants completed surveys including NASA-TLX (Hart and Staveland, 1988), Creativity Support Index (CSI)(Cherry and Latulipe, 2014), self-perceived ML experience(Choi et al., 2024; Wu et al., 2022), and (for Vistoria only) perceived feature usefulness.
-All instruments used 7-point Likert scales.
-Following both conditions, participants completed a 15–20 minute semi-structured interview.
-All sessions were video-recorded via Zoom. We collected system logs, final story drafts, canvas artifacts, image–text pairs, and interview transcripts for analysis.
+After each condition, participants completed surveys including NASA Task Load Index (TLX) (Hart and Staveland, 1988) and Creativity Support Index (CSI) (Cherry and Latulipe, 2014).
+These surveys were chosen to be consistent with the standard measures employed in previous HCI system work on multimodal LLM-assisted ideation and storytelling (Choi et al., 2024; Suh et al., 2025; Darejeh et al., 2024).
+All surveys used 7-point Likert scales.
+After both conditions finished, participants also completed a 15–20 minute semi-structured interview.
+All sessions were video-recorded via Zoom.
+We collected system logs, final story drafts, canvas artifacts, image–text pairs, and interview transcripts for analysis.
 5.2.3. Data analysis
 We employed a mixed-methods approach to systematically analyze three types of data.
-First, interview transcripts were independently open-coded by two authors following an inductive, grounded theory–informed process (Thornberg, 2012); initial codes were iteratively refined into a shared codebook, with themes developed through constant comparison.
-Disagreements were resolved via negotiated agreement, and analytic memos were maintained for transparency.
-Second, interaction data were analyzed through structured video coding by two authors and aligned with system logs on an event-by-event basis.
-We examined the sequences of function use and compared exploration patterns across the baseline and our condition.
-To characterize participants’ divergent–convergent behaviors during the creative process, for each task, we reconstructed exploration structures by defining Directions as top-level trajectories toward a goal, Branches as the diversity of possibilities generated within a direction, and Depth as the mean number of iterative steps within each branch to compare the exploration across the baseline and our condition.
-Finally, for survey measures, we conducted paired-sample t-tests under the assumptions of normality and homogeneity of variance. When assumptions were violated, we used the Wilcoxon signed-rank test.
+For the qualitative interview data, we conducted an inductive, grounded theory-informed analysis (Thornberg, 2012).
+First, two authors independently performed open-coding on 33% of the data, generating an initial set of 30 distinct codes.
+The coders then met to compare code applications, resolve discrepancies through negotiated agreement, and refine the wording and boundaries of each code.
+Through several rounds of discussion, they reached full consensus on all coded segments and consolidated the initial codes into a shared codebook.
+Using the refined codebook, the two authors independently coded half of the remaining transcripts, meeting regularly to prevent coding drift and to determine whether newly emerging codes should be incorporated.
+Ongoing constant comparison within and across interviews was used to further refine relationships between codes.
+Finally, we clustered the codes into four higher-level themes that map onto our design goals.
+The final codebook is shown in Appendix A.6.
+Second, interaction data were analyzed through structured video coding by two authors to quantify tool usage frequency and modality switching events, and researchers aligned them with system logs on an event-by-event basis.
+Finally, for survey measures, we conducted paired-sample t-tests under the assumptions of normality and homogeneity of variance.
+When assumptions were violated, we used the Wilcoxon signed-rank test.
+Given the sample size limitations, we treat these quantitative results as descriptive signals intended to triangulate with and support the qualitative themes.
 6. Study Results
-Figure 6.
-Interaction records of all participants.
-The creative workflow begins with multimodal generation—primarily text, complemented by sketches or images to express intentions, followed by refinement and iteration using visual instruments, during which textual descriptions are continuously revised in parallel.
-Definition of specific behaviors: Generation includes multimodal creation of new cards using Creative Spark, Exact Craft, or Collage; image-based editing refers to operations such as Lasso, perspective shift, and filters; text-based editing covers modifications of generated story segments on the canvas as well as edits made in the text editor; other operations include updates through highlight, cluster, and upgrading global settings.
+Table 1. Comparison of survey results: Vistoria vs. Baseline. Sig.: * $p<.05$; ** $p<.01$
 Vistoria Baseline Statistics
 mean std mean std p Sig.
-Transparent of ML model Transparency 4.250 1.865 3.417 2.234 0.2015
-Controllability 4.750 1.357 4.833 1.586 0.8664
-Sense of Collaboration 5.583 1.505 4.333 1.670 0.0206 *
-Support for Thought Process 5.333 1.371 4.750 1.960 0.3172
-Match to goal 5.500 0.905 4.833 1.337 0.1201
 NASA-TLX Mental 5.16 1.528 3.167 1.337 0.0000 **
@@
 Collaboration 5.500 0.674 4.583 1.505 0.0418 *
-Table 2. Comparison of survey results: Vistoria vs. Baseline. Sig.: * $p<.05$; ** $p<.01$
-6.1. Strategies for Integrating Multimodality in Creative Writing
-To address RQ1, we identified four key interaction strategies employed when using Vistoria.
-First, participants leveraged multimodal elements to externalize and interconnect different branches of their ideas.
-Second, they utilized image-text pairs as catalysts to generate richer and more dynamic narrative descriptions.
-Third, participants employed visual instruments as a mechanism to drive pivotal shifts in their storytelling.
-Finally, they treated images as both checkpoints and traces, enabling them to maintain narrative coherence and facilitate recall throughout the story development process.
-6.1.1. Use Multimodality to Support Expressing and Connecting Ideas
-Log data (Figure 6) and observations (Figure 13) reveal that participants followed a consistent creative workflow when using the system.
-The process typically began with multimodal content generation, where participants combined generation functions across multiple modalities to visualize their intended content.
-During this exploratory phase, participants iteratively collected and refined story elements through generated cards(images-text pairs), at the same time, frequently modifying text in the editor or regenerating images based on updated text.
-As participants explored multiple visualization directions, they employed the Collage function to combine existing canvas content and discover new narrative possibilities, or used visual instruments (Lasso, Filter, Perspective Shifts) to simultaneously modify text and images.
-This process involved continuous comparison with LLM outputs to determine desired versus undesired elements (P1, P2, P4, P8, P10).
-Participants systematically highlighted useful images and text for final translation into the text editor, often using Cluster functions to summarize character or object settings for incorporation into their final writing.
-Multimodal Expression Enhances Creative Intention.
-The multimodal integration enhanced participants’ ability to convey creative intentions.
-As shown in Figure 6 (b), text served as the primary medium but was consistently supplemented with sketches and images for context, yielding significantly higher expressiveness ratings than the baseline ($M_{\text{Vistoria}}=6.08$ vs.
-$M_{\text{Baseline}}=4.33$, p = .023) (Table 2).
-P6 captured this benefit: “Even though I’m not very good at drawing, when I type some texts and sketch a few rough images, the cards with story and images generated can really capture the scene I have in mind.” (Figure 11).
-Nearly all participants (P2, P6, P10, P11) stressed that combining sketches with text and images aligned outputs more closely with their creative intentions.
+6.1. The Usability of Vistoria
+To address RQ1 and RQ2, we analyzed how participants engaged with the designed instrumental operations (Lasso, Collage, perspective shift, and Filter) and the usage patterns.
+In addition, we incorporated quantitative survey results with qualitative data to assess how Vistoria affected participants’ workload.
+6.1.1. The Usefulness of Instrumental Operations
+Lasso as a granularity controller for local-to-global rewriting.
+The Lasso instrument helped participants shift between narrative scales, supporting both macro-level story planning and micro-level detail refinement.
+As P8 noted, “You can write in different scales, especially when you use the Lasso tool, in which you can extract out that specific detail, so [the story] generated in the card is more heterogeneous on the specific point.” (Figure 7(c)).
+Similarly, P9 and P11 described using the Lasso to foreground key points and sustain more focused attention during revision.
+Together, these accounts suggest that the Lasso affords a narrative “zoom,” enabling users to isolate a local fragment for targeted rewriting while maintaining continuity with the broader narrative trajectory within the same interface.
+Figure 7.
+(a) Perspective Shift changes image viewpoint to reframe narrative voice and redirect story development; (b) Filter synchronizes mood and style across media—applying a visual Filter also rewrites the associated text to match the intended emotional tone; (c) Lasso enables participants to operate at different narrative scales by extracting or isolating elements to steer local and global edits.
+The figure provides examples of instrumental operations in use.
+It shows how Perspective Shift reframes narrative voice, how Filters synchronize emotional tone across images and text, and how the Lasso enables writers to operate at different narrative scales by isolating or extracting specific elements for local or global edits.
+Collage Function Enables Creative Recombination.
+Participants used Collage to merge extracted objects or scenes from different images, building connections across disparate elements.
+For instance, when P6 generated a scene of Maya entering a castle, he envisioned a larger structure with taller stairs.
+He sketched a bigger castle and mountain while expressing his intention through text, resulting in a generated card that matched his vision and was directly incorporated into his story.
+P11 articulated the creative freedom Collage technique provided: “This technique doesn’t limit me; I can create abstract or non-abstract sketches, and I can incorporate whatever I want.” (Figure 8).
+Collage thus serves not merely as a usability feature but as a catalyst for externalizing, reconfiguring, and expanding mental imagery into narrative possibilities unachievable through text or images alone.
+Perspective Shift as a narrative frame-shifter.
+The Perspective Shift alters both the image and the story perspective to provide new directions for story development.
+P8 described how shifting perspectives changed the story direction: “Adopting the water’s viewpoint anthropomorphized the water spirit and introduced a regretful undertone that established the story’s emotional framework… changed the development direction, inspiring new writing possibilities.” P5 also experimented with this feature, incorporating a first-person voice (“I didn’t expect this to be so heavy!”) adopted from the system-generated segments into her third-person story (Figure 7 (a)).
+Perspective Shift allowed participants to flexibly reconfigure narrative viewpoint and voice, surfacing new emotional framings and redirecting story trajectories without disrupting their ongoing writing flow.
+Filter as affective parameterization for tone alignment.
+The Filter instrument shaped narrative emotion and tone by visually parameterizing affect.
+P11 noted, “The image’s style influences my story’s tone and direction.
+Text alone doesn’t reveal tone.
+The Filter’s visual changes help me decide the desired feeling when I need to choose between suspenseful or romantic approaches.” Similarly, P9 observed, “Filter visualizes the emotions in my mind when creating stories… I created a sad mood with Filters, and the girl suddenly turned into a sad face.
+It’s very easy to see what it’s doing and easy for me to describe later.” (Figure 7 (b)).
+The immediate visual feedback aligned emotional intent with text, streamlining tone-setting decisions to evaluate the usability.
+Taken together, these instrumental operations transformed localized operations into meaningful viewpoints, scales, and tones.
+They appear to support participants’ intended operational precision and expressiveness, while potentially reinforcing the perception–action loop during the planning and translating of story writing.
+Figure 8.
+Collage.
+Participants used Collage for creative recombination, merging extracted objects/scenes (often mixing sketches with images) to specify visualization and advance story ideas; usability was positively rated.
+The figure demonstrates the Collage operation during story writing.
+Participants combine extracted objects and scenes from multiple cards, often mixing sketches with images, to specify visualization and explore new narrative ideas.
+The example highlights creative recombination as a mechanism for advancing story development.
+6.1.2. Tradeoffs of Multimodal Image-Text Co-editing
+Increased Workload.
+While image–text co-editing may support aspects of narrative coherence and expressiveness, our results suggest that it can also introduce additional workload.
+NASA–TLX scores showed significantly higher mental demand (NASA-TLX: $M_{\text{Vistoria{}}}=5.16$ vs.
+$M_{\text{baseline}}=3.17$, p=.0000; Table 1) and physical demand (NASA-TLX: $M_{\text{Vistoria{}}}=4.67$ vs.
+$M_{\text{baseline}}=2.08$, p=.0002; Table 1), plus moderately higher frustration (NASA-TLX: $M_{\text{Vistoria{}}}=2.75$ vs.
+$M_{\text{baseline}}=1.75$, p=.0204; Table 1).
+Qualitative analysis reveals that this increased workload reflected the higher cognitive and physical effort of coordinating across modalities and actively curating outputs compared with the GPT baseline .
+Part of the mental load may have stemmed from first-time use—learning new image/text operations and switching between modalities (P1, P2, P5, P9, P10).
+Additional difficulty also arose from unfamiliarity with the canvas interface compared with the traditional GPT interface (P3, P10).
+Below we present the tension of Image-Text co-editing based on qualitative feedback.
+The Cognitive Effort of Enhanced Sense of Agency.
+Nine participants (all except P2, P7, P12) felt they have higher sense of agency, and valued maintaining story control and felt they were directing the narrative through their sketches, inputs, and use of system tools (P4, P6, P10).
+However, P1, P3, P9, and P10 noted that they had to actively develop details within their own text, especially at the early stages, which could feel “a little bit frustrating” (P9).
+Unlike GPT, which could quickly produce long passages or propose questions to guide brainstorming (P10), the system requires participants to supply and elaborate on their own ideas before meaningful generation occurs, potentially leading to a higher mental workload.
+This shift demanded more cognitive and physical effort: even though sketching and annotation helped externalize mental imagery, participants noted that it felt more demanding than simply inputting and receiving GPT’s ready-made text (P1, P2, P11).
+Thus, a stronger sense of agency may come at the cost of a higher workload.
+Validating Ideas Rather than Generating Them.
+Some participants noted that the system’s strengths lay in validating or expanding existing concrete mental images rather than generating new and abstract directions (P1, P11).
+As P1 explained, “When I have a vague impression in my mind, I tend to generate some image-text pairs.
+But sometimes, once the visual appears, it fixes my imagination in a certain way in my mind, and I can no longer imagine other possibilities.
+In contrast, only plain text can inspire limitless imagination.” This reveals a tension: images act as concrete anchors that aid detailed development, yet their representational specificity can induce fixation by prematurely crystallizing fuzzy concepts and narrowing exploration.
+Similarly, P10 noted that the baseline GPT condition was superior at breaking down initial story points and directly providing additional suggestions and directions, whereas the Vistoria system primarily served to elaborate or diverge from existing visions the user already has.
+This suggests this multimodal approach may be valuable for participants with partially formed concepts, though potentially less helpful during the open-ended phases of ideation when abstract exploration is more important than visual specificity.
+Externalization Frees Cognitive Space.
+Although participants reported experiencing higher mental workload, several accounts suggest that multimodality may have supported a more efficient allocation of attention (P4, P7).
+As P7 described, “By highlighting and collaging, I externalized formed ideas into image–text pairs, clearing mental space to pre-plan the next line and concentrate on the next plot beat.” It indicates that the creation process has the potential to externalize fleeting ideas, preserve sensory and spatial detail, and reduce information loss when translating imagination into concrete artifacts.
+6.2. Creativity Support of Vistoria
+To address RQ3 and RQ4, we examine how Vistoria supports intention expression through multimodal input, facilitates the ideation process through divergent exploration, as well as how image–text alignment design contributes to narrative development.
+We also describe how this workflow potentially enhances participants’ sense of agency and ownership.
+Figure 9.
+Multimodal Expression.
+Example where sketch structure and textual details jointly yield multiple relevant images; participants valued sketches for visualizing spatial layout beyond what text-only tools could provide.
+The figure shows an example of multimodal expression where rough sketches and textual descriptions jointly guide image generation.
+The resulting images reflect the spatial layout and visual structure conveyed by sketches that would be difficult to express using text alone, supporting more precise externalization of mental imagery.
+6.2.1. Enhancing Intent Expression Through Multimodal Input.
+As shown in Appendix Figure 13 (b), text served as the primary medium for generation but was consistently supplemented with sketches and images to provide more spatial information.
+Eight participants (P1, P2, P3, P6, P7, P9, P10, P11) stressed that combining sketches with text and images aligned outputs more closely with their creative intentions, yielding significantly higher expressiveness ratings than the baseline ($M_{\text{Vistoria}}=6.08$ vs.
+$M_{\text{Baseline}}=4.33$, p = .023; Table 1).
+P6 captured this benefit: “Even though I’m not very good at drawing, when I type some texts and sketch a few rough images, the cards with story and images generated can really capture the scene I have in mind.”.
 P7 illustrated this with a concrete case: she sketched a rough flower and archway, then added text specifying a glowing flower and a Gothic gate.
@@
 She highlighted the unique value of sketching: “Drawing is pretty cool.
-In GPT, I want to draw a mental image, but GPT cannot… the geometry is always different from what’s in my head.“ (Figure 7).
-This suggests that multimodal expression enabled participants to externalize their mental imagery and refine it into concrete, shareable representations, bridging the gap between vague internal visions and precise creative outputs.
+In GPT, I want to draw a mental image, but GPT cannot… the geometry of GPT-generated image is always different from what’s in my head.” (Figure 9).
+This suggests that multimodal expression enabled participants to externalize their mental imagery and refine it into concrete, shareable representations, bridging the gap between vague internal visions and precise outputs.
+6.2.2. Bottom-up Creation to Support Divergent Exploration
+As shown in Appendix Figure 12 and Appendix Table 2, participants demonstrated greater breadth and depth of exploration using Vistoria and produced visibly more divergent narrative structures compared to the linear outputs of the GPT baseline.
+The Vistoria canvas functioned as an exploratory space where participants pursued multiple storylines in parallel without linear constraints.
+Rather than committing immediately to single narratives, participants typically generated multiple alternatives in early phases.
+As P8 observed: “GPT workflow is more streamlined… top-down.
+Using the system feels more bottom-up.
+You are open to possibilities, and then you choose one way to go deep, so there’s not a finite result and more possibilities being explored.”
+The exploratory nature of Vistoria was also described as playful and enjoyable (P1, P7).
+As P1 reflected: “Using the tool feels more like doing Collage or drawing on a whiteboard or a large sheet of paper, where you can do almost anything—it’s very free and interesting.” This experiential quality aligns with the higher immersion reported for Vistoria compared to the baseline (CSI: $M_{\text{Vistoria}}=4.92$ vs.
+$M_{\text{Baseline}}=2.75$, p = .0006; Table 1).
 Figure 10.
-Multimodal Expression.
-Example where sketch structure + textual details jointly yield multiple relevant images; participants valued sketches to visualize spatial layout beyond what text-only tools could provide.
-Collage Function Enables Creative Recombination.
-The Collage function emerged as a central strategy for creative recombination, receiving positive usability ratings (M=5.33, SD=1.5) (Figure 8).
-All participants used collage to merge extracted objects or scenes from different images, building connections across disparate elements.
-For instance, when P6 generated a scene of Maya entering a castle, he envisioned a larger structure with taller stairs.
-He sketched a bigger castle and mountain while expressing his intention through text, resulting in a generated card that matched his vision and was directly incorporated into his story (Figure 9).
-P11 articulated the creative freedom this technique provided: “This technique doesn’t limit me; I can create abstract or non-abstract sketches, and I can incorporate whatever I want.” This multimodal recomposition enabled participants to quickly express envisioned scenes (P2) and provided “more freedom to envision and create the story” (P10).
-These practices highlight that collage is not merely a usability feature but a catalyst for multimodal recomposition, enabling participants to externalize, reconfigure, and expand their mental imagery into coherent narrative possibilities that text or images alone could not achieve.
-Furthermore, multimodal input also increased participants’ immersion with Vistoria compared to the baseline ($M_{\text{Vistoria}}=4.92$ vs.
-$M_{\text{Baseline}}=2.75$, p = .0006) (Table 2).
-This heightened immersion stemmed from the playful, enjoyable nature of multimodal creation and the emotional reinforcement of freely expressing intentions across modalities.
-As P1 reflected: “Using the tool feels more like… doing something on a whiteboard or a big piece of paper, where you can do almost anything—it’s very free and interesting.” Together, multimodal inputs and outputs fostered a more engaging and emotionally resonant creative experience.
-Figure 8. The perceived usefulness of each key functionality in Vistoria
-6.1.2. Instrumental Interaction Tools Drive Narrative Development
-Analysis of interaction logs (Figure 6 and 13) shows that participants often engaged with visual instruments in the mid-phase of creation, and these instruments directly shaped how stories were developed and revised.
-Perspective Shift as a narrative frame-shifter.
-The Perspective Shift instrument, which altered both the image and the story perspective, was rated among the most useful features (M=5.92, SD=1.56) (Figure 8).
-P8 described how shifting perspectives changed the story direction: “Adopting the water’s viewpoint anthropomorphized the water spirit and introduced a regretful undertone that established the story’s emotional framework.
-Changing the perspective of the voice of the story added personal sentiment and changed the development direction, inspiring new writing possibilities.” P5 also experimented with this feature, incorporating a first-person voice (“I didn’t expect this to be so heavy!”) adopted from the system-generated segments into her third-person story (Figure 10 (a)).
-Perspective Shift allowed authors to flexibly reconfigure narrative viewpoint and voice, surfacing new emotional framings and redirecting story trajectories without disrupting their ongoing writing flow.
-Lasso as a granularity controller for local-to-global rewriting.
-The Lasso instrument was also valued as useful (M=5.0, SD=1.28) (Figure 8) for enabling writers to zoom between different narrative scales.
-P8 emphasized, “You can write in different scales, especially when you use the Lasso tool, in which you can extract out that specific detail, so [the story] generated in the card is more heterogeneous on the specific point.”(Figure 10 (b)) This reflects how the lasso tool enables a narrative “zoom” functionality, allowing writers to oscillate between macro-level story development and micro-level detail refinement within a single interface.
-Similarly, P9 and P11 described how the lasso allowed them to deepen or refine specific text and emphasize the key points with more focused attention.
-In this way, the Lasso operates as a narrative instrument, turning macro-level edits into micro-level adjustments while preserving precision, enabling rollback, and sustaining fluent exploration.
-Figure 9.
-Collage.
-Participants used Collage for creative recombination, merging extracted objects/scenes (often mixing sketches with images) to specify visualization and advance story ideas; usability was positively rated.
-Filters as affective parameterization for tone alignment.
-The Filter instrument shaped narrative emotion and tone by visually parameterizing affect.
-P11 noted, “The image provides the style, which influences my story’s tone and direction… After using the filter, I can’t determine tone from text alone—I need to choose between suspenseful or romantic, and the visuals help me decide.” Similarly, P9 observed, “Filter visualizes the emotions in my mind when creating stories… I created a sad mood with filters, and the girl suddenly turned into a sad face.
-It’s very easy to see what it’s doing.” (Figure 10 (c)).
-The immediate visual feedback fostered a more intuitive workflow, aligning emotional intent with visual representation.
-The immediate visual feedback aligned emotional intent with representation, streamlining tone-setting decisions.
-Taken together, these tools acted as narrative instruments that transformed localized operations into meaningful shifts in viewpoint, scale, and tone.
-They enhanced authors’ control, amplified expressiveness, and sustained a fluid perception–action loop throughout the creative process.
-Figure 10.
-(a) Perspective Shift was rated highly useful (M=5.92); changing the image viewpoint also reframed narrative voice and redirected story development, as participants reported; (b) Filters (M=4.25) synchronize mood and style across media—applying a visual filter also rewrites the associated text to match the intended emotional tone, ensuring stylistic coherence; (c) Lasso tool (M=5.0) enabled writers to focus at different narrative scales by extracting or isolating elements to steer local and global edits.
-6.1.3. Structural Mapping between Images-Text Paris as a Catalyst for Vivid Story Writing
-Participants emphasized that when text and images appeared together, images served as direct references during writing, easing detailed description and enriching narrative depictions.
-This integration worked through two cognitive mechanisms that supported vivid story construction.
-Images-Text Paris as cognitive scaffolding for unfamiliar scenarios.
-Visual references helped participants imagine actions or settings beyond their lived experience.
-As P11 reflected, “I’m more comfortable describing conversations and overall scenarios than people’s actions—especially kids, who might have more interesting reactions.
-Visualizing with text descriptions helps me imagine unfamiliar actions, which I can then use in my story.” Here, the image acted as a structural anchor by aligning the visual cues of children’s postures and expressions with possible textual descriptions.
-In other words, structural mapping between what was seen (visual structure) and what could be narrated (linguistic structure) scaffolded richer and more plausible action descriptions.
-Images-Text Paris enabling descriptive language translation.
-Participants also converted visual elements into vivid prose, effectively mapping structures across modalities.
-As P5 noted, “Looking at an image makes it easier to kind of, like, describe what something is supposed to look like or feel like… so when you describe it to [readers], and they can picture it for themselves in their head, that is what makes the story more interesting.” From observation, P5 closely described the final images in her writing and directly adopted text from the generated cards, integrating them into her final story.
-(Figure 12) This visual-to-textual translation was also evident in concrete writing outcomes.
-For example, P1 used visual cues from an image showing Claire touching a letter and adopted the descriptions in the text such as “Claire steadies the box on her lap” and “She runs her fingers over the letters, heartbeats echoing in her ears.” to form his final story.
-(Figure 11) Here, the structural alignment between gesture in the image and embodied description in text allowed P1 to carry emotional undertones across modalities, turning static imagery into dynamic, sensory-rich narrative moments.
-Overall, structural mapping between visual and textual representations scaffolded descriptive richness, enhanced vividness, and supported schema formation for scenarios participants might otherwise have struggled to imagine.
-Figure 11. Visual-to-text translation: writers turn visual cues into vivid prose that readers can picture; sample lines show how P1 used sensory-rich descriptions derived from an image-text pair.
-6.1.4. Structural Mapping across Multiple Dimensions to Trace the Story
-Participants used not only text but also images as visual checkpoints to trace story direction and development.
-Because images were more immediately perceptible than text, they helped participants align outputs with intended style or mood and, when viewed in sequence, detect progressions and gaps (P6, P7, P2).
-Within Vistoria, this process evolved into structural mapping: participants systematically compared visual–textual structures across dimensions, gaining multimodal visibility into narrative development.
-This layered view created a relational space where main storylines and branching alternatives could be mapped against one another, allowing participants to perceive both continuity and divergence, thereby fostering coherence while supporting flexible redirection.
-As P12 noted, visuals alongside text set the mood and made it easier to recall earlier ideas, illustrating how structural mapping provided mnemonic scaffolding that supported continuity in storytelling.
-Compared to GPT’s linear text-only workflow, participants found that Vistoria’s visual traces better supported narrative management and reduced idea drift or loss (P12, P11, P2).
-P2 reflected on this benefit explicitly:
-“Using the system, we started with a baseline picture and then came up with another picture… it was easy to trace the development of the story.
-If you get too far down that chain and don’t like it, you can just delete that node and go in a different direction.
-I liked visually being able to see the progress from generation to generation.”
-Here, structural mapping across successive images enabled participants to align branches and selectively prune or redirect them.
-Such multi-dimensional mapping clarified narrative progression, improved recall and coherence, and supported flexible branching and gap detection, sustaining sensemaking throughout the evolving story.
-Figure 12.
-Images act as cognitive scaffolds, helping writers describe unfamiliar actions or contexts more concretely.
-P5 constructed her final writing by referring to both images and adopted and edit related text.
-6.2. Multimodal Approaches Creation to Inspire Creativity
-To answer RQ2, we found that image-text co-editing facilitated creativity in fictional story writing through three key cognitive processes: (1) divergent exploration, where participants spatially arranged multimodal components and pursued multiple narrative directions; (2) serendipitous discovery, where unexpected LLM outputs sparked new creative possibilities; and (3) multimodal cues activating memory and cultural connections.
-Together, these processes enhanced participants’ sense of authorship by positioning the system as a collaborative partner that expanded creative possibilities rather than replacing human imagination.
+Images act as cognitive scaffolds, helping participants describe unfamiliar actions or contexts more concretely.
+P5 constructed her final writing by referring to both images and adopting and editing related text.
+The figure shows an example canvas during early exploration, where multiple image–text cards, sketches, and annotations are distributed spatially.
+Different narrative directions are pursued in parallel, illustrating non-linear exploration before committing to a single storyline.
+6.2.3. Leveraging Image-Text Alignment for Rich Descriptions and Story Progression
+When text and images appear together, the two modalities reinforce one another, enabling unfamiliar or imagined elements to be both visualized and verbalized.
+This cross-modal grounding supports the production of more detailed and enriched descriptions.
+Visual references, in particular, helped participants imagine actions or settings beyond their lived experience.
+As P5 noted, “Looking at an image makes it easier to kind of, like, describe what something is supposed to look like or feel like… so when you describe it to [readers], and they can picture it for themselves in their head, that is what makes the story more interesting.” From observation, P5 closely described the scene in her imagination with the cliff, the necklace, and the cat in the image in her writing and directly adopted some text generated, integrating them into her final story (Figure 10).
+Using the generated text and image also helped produce more concrete, detailed narratives.
+For example, P1 used visual cues from an image showing Claire touching a letter and adopted the descriptions in the text, such as “Claire steadies the box on her lap” and created the narrative of “She runs her fingers over the letters, heartbeats echoing in her ears.” to form his final story.
+Here, the text description with visualization allow participant to capture more dynamic, sensory-rich narrative moments.
+6.2.4. Preserving Sense of Agency and Ownership.
+Unlike the baseline, where participants often felt like they were editors of GPT-generated text, Vistoria supported exploratory editing while preserving a sense of agency and ownership.
+Participants felt that Vistoria “having more sense of mastery over the content,” in contrast to the baseline (P4, P6, P9).
+P7 felt that when using Vistoria, she was actively cutting, combining, filtering, and directing the story’s trajectory, whereas with GPT, she was mostly receiving and editing what the model produced.
+P9 explicitly emphasized a heightened sense of agency, noting: “This tool is more ‘me’… I control characters and plots.”
+Dissatisfied with the baseline condition, in which GPT produced most of the content and left participants primarily in the role of adopting it (P4, P6, P7), P10 also characterized Vistoria as a supportive co-pilot rather than a substitute for their own work.
+Quantitative results are consistent with these perceptions: participants rated Vistoria as providing a stronger collaborative experience than the baseline (CSI: $M_{\text{Vistoria}}=5.50$ vs.
+$M_{\text{Baseline}}=4.58$, p = .0418; Table 1).
+This preserved sense of agency also led to the preserved sense of ownership.
+GPT outputs in the baseline condition were repeatedly described as “surface–level” (P2, P5, P7) and as “someone else’s work” (P3), whereas most participants reported a stronger sense of ownership with Vistoria (P1, P2, P3, P4, P5, P6, P8, P9, P10, P11).
+As P5 explained: “When using Vistoria, every idea originated from my own imagination, and the final story was formed by manipulating and combining these different self-generated ideas.
+This gave me a strong sense that the story was truly my own creation.”
+Together, these findings suggest a shift from passively adopting model suggestions to actively creating and curating one’s own generative outputs.
+7. Discussion
+7.1. Multimodal Instrumental Interaction
+Instrumental Interaction (Beaudouin-Lafon, 2000) conceptualizes instruments as mediators that translate writers’ actions into operations on domain objects.
+We operationalize this principle by reifying a set of multimodal instrumental operations (Collage, Lasso, Filter, and Perspective Shift) that simultaneously act upon both text and image narrative materials.
+Rather than treating images and text as separate interface elements, these instruments serve as unified interactional units that enable writers to zoom between narrative scales, reorganize multimodal story fragments, and explore divergent directions within a shared representational space to edit image and text simultaneously.
+From the perspective of Designing Interaction, not Interfaces (Beaudouin-Lafon, 2004), our design moves away from adding more interface widgets and focuses instead on shaping the quality of writers’ ongoing activity.
+Beaudouin-Lafon argues that transformative interfaces must shift attention from surface-level user-interface (UI) components toward the underlying interactional structures that support creative work (Beaudouin-Lafon, 2004).
+Following this perspective and Dual Coding Theory’s suggestions for verbal and nonverbal perception alignment (Clark and Paivio, 1991), Vistoria’s design prioritizes fluid transitions between modalities, persistent manipulable artifacts, and an iterative loop in which narrative ideas and multimodal materials co-evolve.
+This interaction-centered framing explains why writers perceived the system as increasing expressiveness.
+Although participants generally found the instruments effective, they also reported a substantial learning curve.
+For novice system writers, the system imposed considerable mental and physical workload.
+These insights suggest future design opportunities to lower workload and adapt to writers’ evolving needs.
+In the early stages of system use, writers could express their intentions through natural language, allowing the system to suggest the most appropriate functions on their behalf (Bourgault et al., 2025).
+As writers become more familiar with the system, they can choose functions by themselves and even define customized functions that better fit their evolving needs.
+This approach aligns with the emphasis on supporting diverse, situated practices rather than enforcing a fixed interface vocabulary (Beaudouin-Lafon, 2004).
+Together, these directions point toward a design space where multimodal creativity systems integrate explicit instruments with adaptive, activity-centered interaction models to better support real-world writing workflows.
+7.2. Designing Dynamic Mixed-Initiative Multimodal Workflows
+In the traditional turn-based GPT workflow, writers often occupy a relatively passive or evaluative role: they receive model output and act primarily as examiners who check, accept, or correct the result (Kwan et al., 2024).
+In Vistoria, writers actively manipulate elements on the surface, decide which multimodal materials to combine, and select which operations to apply.
+From the perspective of the participants, Vistoria is not experienced as a detached “answer engine,” but as a co-pilot collaborator.
+Writers perceive themselves as those who decide how to create, what to keep, and which tools to invoke, preserving the sense of agency and ownership.
+However, this also brings cognitive effort.
+Precisely because the workflow is instrument and manipulation-driven, it demands that writers have a clearer sense of what they need, or at least which direction they wish to explore.
+When writers do not yet know what they want, the system requires them to specify intentions and choose operations that tend to lead to higher cognitive load.
+This reflects an inherent tension: high functional agency (agency is structurally distributed in the system) requires clear intentionality, but early-stage ideation is often characterized by ambiguity and exploration (Rafner et al., 2025).
+In contrast, a text-only GPT chat enables the rapid generation of a large amount of content, allowing for subsequent refinement through iterative prompting and selection to achieve a specific focus.
+This suggests that systems should adapt their agency distribution across creative phases as agency and ownership are not static end-states but fluctuate across the co-creation trajectory (Rafner et al., 2025).
+In early divergent exploration, lower functional agency may paradoxically support higher felt agency (e.g., through control, decision-making, or ownership) by reducing decision fatigue; in later convergent refinement, higher functional agency (granular control) aligns with writers’ desire to shape details (Rafner et al., 2025; Fu et al., 2024).
+On the other hand, ownership emerges from users’ active control over outputs and their ability to modify content (Draxler et al., 2024a).
+Critically, this control includes not only refinement flexibility but also rejection.
+Our results show that participants sometimes think image output “fixed imagination in a certain way,” which may threaten to constrain their creative vision.
+Previous work indicates that when AI outputs deviate from expectations, users experience diminished control and ownership uncertainty (Draxler et al., 2024a), and encouraging system design to prompt them to deploy adaptive strategies to restore agency.
+In that case, future designs could operationalize this through a function that, when invoked, asks “What specifically do you want to avoid?” and generates alternatives that explicitly diverge from the rejected output.
+This would transform counter-inspiration (Rafner et al., 2025) from an ad-hoc user strategy into operations that reframe dissatisfaction from a null action into a mechanism for articulating and preserving authorial boundaries and ownership.
+7.3. Limitation and Future Work
+While our study provides valuable insights into multimodal story writing, several limitations constrain the generalizability and scope of our findings.
+Task Scope and Short-Term Focus.
+The 300-to-500-word story task, while manageable for controlled evaluation, does not reflect the demands of long-form fictional writing, where authors build sustained voices, complex arcs, and intricate structures (Chen et al., 2025).
+We also did not have an opportunity to study Vistoria’s suitability for extended projects, iterative revisions, or complex narratives, leaving open questions on the consistency in long works, scalability with larger volumes, or risks of over-reliance on LLM over time of use.
+Moreover, evaluation relied primarily on self-reports of creativity and user experience; we did not include objective measures of story quality, originality, or literary merit.
+In the future, we plan to conduct field studies that deploy this system with writers of varying expertise levels in their authentic creative contexts, observing how they integrate the tool into real writing projects over extended periods.
+Such longitudinal research would assess the ecological validity of Vistoria and provide insights into how writers adapt Vistoria for planning and translating across longer creative cycles (Long et al., 2025).
+We anticipate that writers might spontaneously capture inspirational moments from daily life, potentially increasing their reliance on clustering functionality as they generate more dispersed content fragments that require organization.
+These naturalistic studies would provide crucial insights into the tool’s role in sustained creative practice, revealing usage patterns, adaptation strategies, and long-term impacts on writers’ creative processes that controlled laboratory settings cannot capture.
+Participant Sample.
+Our study involved only 12 participants, while this is typical for similar lab usability studies, a larger group could provide stronger statistical power, reveal more varied interaction patterns, and allow comparisons across subgroups.
+Furthermore, the group of participants has limited cultural and age diversity, which could have narrowed the range of narrative traditions, writing styles, and storytelling approaches represented.
+Future research should address these limitations by recruiting a larger and more diverse set of participants, including writers of various ages and individuals from diverse cultural backgrounds, to more fully evaluate the applicability and generalizability of the system.
+Construct Validity and Measurement Limitations.
+Although we discuss constructs such as creativity, sense of ownership, and agency, these observations arise primarily from qualitative reports.
+Our study does not include construct-grounded measurements or comparative baselines for these phenomena.
+Accordingly, the interpretations should be viewed as exploratory insights rather than empirically validated effects.
+Future work will incorporate construct-aligned measures and validated scales such as the Mixed-Initiative CSI (Lawton et al., 2023), or using frameworks situated within human–AI co-creativity like COFI (Rezwana and Maher, 2023) and CCDF (Davis et al., 2025) to quantify co-creativity interaction dynamics or better derive Human-AI co-creativity insights (Rezwana and Maher, 2025).
+Furthermore, while we assessed participants’ AI usage frequency (e.g., daily ChatGPT usage), we did not measure their collaborative AI literacy or metacognitive capabilities when working with AI systems.
+Future research should incorporate validated scales  (Sidra and Mason, 2025) to investigate how collaborative AI literacy influences the effectiveness of multimodal co-editing tools.
+Multimodality Scope.
+Our work focuses on multimodal support through text and images, but does not include other modalities.
+Prior research in creative writing suggests that audio can also serve as a useful medium (Singh et al., 2023), especially through nonverbal sounds and ambient effects that help shape mood and atmosphere.
+In future work, we plan to explore the addition of audio cues to the writing process.
+Such sound elements may support writers in building a stronger vibe, enhancing scene-setting, and offering an additional channel for creative inspiration (Rezwana and Maher, 2023).
+8. Conclusion
+This paper presents Vistoria, a multimodal image–text co-editing system that supports fictional story writing by tightly integrating image and text representations.
+Grounded in the WoZ co-design study, Vistoria introduces a unified set of instrumental operations (Lasso, Collage, Perspective Shift, and Filter) that reify writers’ intentions and enable synchronized manipulation across modalities.
+Through a controlled user study, we demonstrate that multimodal co-editing enhances expressiveness, immersion, and exploratory ideation.
+Although this multimodal workflow increases cognitive demand, participants reported preserved senses of agency and ownership, treating the system as a creative partner rather than a generative tool.
+We hope Vistoria highlights the opportunities for designing future writing systems that embrace multimodality as a core mechanism for ideation and narrative development.
+Acknowledgments
+This work is supported by City University of Hong Kong Teaching Development Grant (6000901), City University of Hong Kong Booster Fund (7030021), and Hong Kong Research Grants Council Theme-based Research Scheme (T45-205/21-N).
+We sincerely thank all anonymous reviewers for their insightful comments and constructive feedback that helped improve this paper.
+Appendix A Appendix
+A.1. User Behavior in the WoZ co-design study
+Figure 11. Results of user strategies for manipulating multimodal elements in the WoZ co-design study.
+Results of user strategies for manipulating multimodal elements in the WoZ co-design study.
+A.2. Behavioral interaction data gathered from participants in the user study
 Figure 12.
@@
 Participants also tended to pursue deeper exploration within specific directions when using Vistoria.
-Vistoria Baseline
-Mean # of directions 6.92 $\pm$ 2.81 1.42 $\pm$ 1.08
-Mean # of branches 3.00 $\pm$ 1.35 1.92 $\pm$ 0.67
-Mean depth 1.70 $\pm$ 1.18 2.00 $\pm$ 1.22
+The figure compares story development outcomes between the baseline text-only condition and Vistoria.
+It visualizes differences in narrative branching and structural diversity, showing how Vistoria supports parallel exploration and alternative story trajectories.
+We examined the sequences of function use and compared exploration patterns across the baseline and our condition.
+To characterize participants’ divergent–convergent behaviors during the creative process, for each task, we reconstructed exploration structures by defining Directions as top-level trajectories toward a goal, Branches as the diversity of possibilities generated within a direction, and Depth as the mean number of iterative steps within each branch to compare the exploration across the baseline and our condition.
 Table 2.
@@
 Depth indicates the mean number of iterative steps within each branch.
-6.2.1. Bottom-up Creation to Support Divergent Exploration
-As shown in Figure 13 and Table 3, participants demonstrated significantly greater breadth and depth of exploration with Vistoria compared to the GPT baseline.
-They pursued a wider range of narrative directions ($M_{\text{System}}=6.92,SD=2.81$ vs.
-$M_{\text{Baseline}}=1.42,SD=1.08$) while also producing more variations within single paths (branches: $M_{\text{System}}=3.00,SD=1.35$ vs.
-$M_{\text{Baseline}}=1.92,SD=0.67$).
-Although the average depth per path was slightly lower for Vistoria ($M_{\text{System}}=1.70,SD=1.18$ vs.
-$M_{\text{Baseline}}=2.00,SD=1.22$), this reflected participants’ substantially broader exploration overall; within individual directions, they often engaged in deeper iterations, as evident in Figure 13.
-Such enhanced exploration transformed their creative process from passively consuming LLM outputs to actively curating multimodal components.
-The Vistoria canvas functioned as an exploratory space where participants pursued multiple storylines in parallel without linear constraints.
-This freeform environment encouraged spatial arrangement, clustering, and repositioning of ideas through manipulation and recombination of alternative story–image cards to explore divergent scenarios (P7, P4, P5, P8).
-Rather than committing immediately to single narratives, participants typically generated multiple alternatives in early phases, positioning the system as an expressive medium rather than merely a text generator.
-This exploratory approach helped participants avoid early fixation and sustained creative engagement.
-As P8 observed: “GPT workflow is more streamlined… top-down.
-Using the visual app feels more bottom-up.
-You are open to possibilities, so there’s not a finite result and more possibilities being explored.” Ultimately, this shift highlights how multimodal, spatially organized exploration fosters divergent thinking, sustains creative momentum, and supports a more generative form of narrative authorship.
-6.2.2. Serendipity from LLM Randomness Enriches Narrative Development
-Another recurring theme was that the inherent randomness of LLM-generated image-text output often introduced serendipitous elements into participants’ creative process.
-Rather than being perceived as noise, these unexpected details were frequently embraced as inspiration, enriching story development and occasionally opening entirely new narrative directions.
-For example, when P1 generated an image-text pair of the mysterious box, the system unexpectedly inserted the logo “LEH.” on the box and in the text description.
-Intrigued, P1 subsequently asked the system to visualize who or what “LEH” could be, gradually evolving this accidental motif into a recurring symbol within their final story (Figure 11).
-Nearly all participants described how surprises in LLM-generated images or texts gave them fresh possibilities to explore.
-P12 captured this dynamic vividly: “The system gave me a riddle at an early stage, I didn’t understand it at first, but as I kept working, I slowly decoded it, and this process of unraveling became part of my story.
-It helped me construct something unique that I wouldn’t have created otherwise.” (Figure 14) Together, these examples show how the interplay of multimodal generation and LLM’s inherent unpredictability fueled creativity.
-Randomness supported divergence by surfacing novel ideas, while participants retained agency in selectively integrating these elements, leading to unique, personally authored narratives.
+Vistoria Baseline
+Mean # of directions 6.92 $\pm$ 2.81 1.42 $\pm$ 1.08
+Mean # of branches 3.00 $\pm$ 1.35 1.92 $\pm$ 0.67
+Mean depth 1.70 $\pm$ 1.18 2.00 $\pm$ 1.22
 Figure 13.
-In the creation process of P12, unexpected LLM randomness served as serendipitous prompts: a generated riddle was decoded and transformed into a key-and-note motif guiding Claire toward the lighthouse, illustrating how stochastic details became generative cues that enriched development and redirected the plot.
-6.2.3. Unlocking Creativity Through Memory and Cultural Resonance
-Participants reported that multimodal cues routinely triggered personal memories and cultural references, enriching their stories.
-Visual props, e.g., a medieval mask, readily invoked film/literary associations; as P4 noted, “Seeing the mask immediately brought Batman/Phantom to mind… it made choosing a style much easier.” (Figure 15) Crucially, image–text pairs co-triggered lived experiences.
-In P7’s case, a fountain scene image together with the phrase “mossy stone” jointly evoked a vivid memory, which she then translated into a concrete narrative element, the envisioned fall that became her story’s ending(Figure 15).
-These multimodal traces turned associative recall into usable story material, weaving authenticity and imaginative depth into the narrative.
-6.2.4. Strengthened Authorship and Agency
-Enhanced exploration with Vistoria strengthened participants’ sense of authorship.
-While GPT outputs in the baseline condition were consistently described as “surface-level” (P2, P5, P7) and felt like “someone else’s work” (P3), almost all participants reported markedly stronger authorship with Vistoria, enabling deeper exploration and richer plots (P1, P2, P3, P5, P12, P10).
-P5 articulated this shift:
-“When using Vistoria, every idea originated from my own imagination, and the final story was formed by combining these different self-generated ideas.
-This gave me a strong sense that the story was truly my own creation.
-I think of the system less as a generator and more as a companion sketchbook.
-While you explore different directions, you always remain in charge of the final piece.”
-P9 reinforced this sentiment: “This tool is more ‘me’… I control characters and plots; ChatGPT may not become my story.” Quantitative results corroborated these perceptions: participants rated Vistoria as providing stronger collaborative experiences (CSI: $M_{\text{Vistoria}}=5.50$ vs.
-$M_{\text{Baseline}}=4.58$, p = .0418) (Table 2).
-Participants framed Vistoria as a supportive co-pilot rather than a replacement; as P10 noted: “I consider this system as a copilot that does not help you with everything, and this system that you have to create by yourself.” Together, these findings mark a transition from passively adopting others’ ideas to actively curating one’s own generative outputs, reinforcing creative agency and ownership.
-Figure 15.
-Evolving cards leave visual traces that cue personal memories and cultural references (e.g., a mask → Batman/Phantom; a “mossy stone” → place memory), which writers externalize into concrete scenes.
-These multimodal traces act as mnemonic scaffolds, helping retrieve earlier thinking and weave lived experience into the story world—transforming associative recall into narrative material.
-6.3. Tradeoffs of Multimodal Interaction for Creative Writing
-While image–text co-editing enhanced authorship, coherence, and expressiveness, it also increased workload.
-NASA–TLX scores showed significantly higher mental demand ($M_{\text{Vistoria{}}}=5.16$ vs.
-$M_{\text{baseline}}=3.17$, p=.0000) and physical demand ($M_{\text{Vistoria{}}}=4.67$ vs.
-$M_{\text{baseline}}=2.08$, p=.0002), plus moderately higher frustration ($M_{\text{Vistoria{}}}=2.75$ vs.
-$M_{\text{baseline}}=1.75$, p=.0204) (Table 2).
-These increases reflect the effort of coordinating across modalities and actively curating outputs relative to the GPT baseline.
-Part of the mental load may have stemmed from first-time use—learning new image/text operations and switching between modalities (P1, P2, P5, P9, P10): “The biggest burden is switching between tools to sketch or type; it takes time to learn and adapt, even though the functions are useful.” Additional difficulty arose from unfamiliarity with the canvas interface (P3, P10).
-6.3.1. Higher burden of authorship.
-Almost all participants valued the ability to maintain control of the story, and the creation process gave them a stronger sense of ownership.
-Yet, this empowerment came with a cost: several participants noted that they had to actively develop details within their own text, especially at the early stages, which could feel “a little bit frustrating” (P9).
-Unlike GPT, which could quickly produce long passages or propose questions to guide brainstorming (P10), the system required users to supply and elaborate on their own ideas before meaningful generation occurred, which may lead to a higher mental workload.
-This shift demanded more cognitive and physical effort: even though sketching and annotation helped externalize mental imagery, participants noted that it felt more demanding than simply receiving GPT’s ready-made text (P1, P2, P11).
-Thus, stronger authorship agency came at the cost of higher workload.
-6.3.2. Validating ideas rather than generating them.
-Some participants noted that the system’s strengths lay in validating or expanding concrete mental images rather than generating abstract directions(P1, P11).
-As P1 explained, “when I have a vague impression in my mind, I tend to generate some image-text pairs.
-But sometimes, once the visual appears, it fixes my imagination in a certain way in my mind, and I can no longer imagine other possibilities.
-In contrast, only plain text can inspire limitless imagination.” This reveals a tension: images act as concrete anchors that aid detailed development, yet their representational specificity can induce fixation by prematurely crystallizing fuzzy concepts and narrowing exploration.
-Similarly, P10 noted that the baseline GPT condition was superior at breaking down initial story points and offering prompt-like guidance, whereas the multimodal Vistoria system primarily served to elaborate or diverge from existing visions.
-This made the multimodal approach particularly valuable for writers with partially formed concepts, but less helpful during the most open-ended phases of ideation when abstract exploration is more important than visual specificity.
-6.3.3. Cognitive offloading and efficiency despite higher load.
-Despite a higher mental workload, participants repeatedly emphasized that multimodality enabled more efficient allocation of attention.
-From participants’ view, the image-text pairs externalized fleeting ideas, preserved spatial and sensory detail, and reduced information loss when translating imagination into tangible artifacts (P4, P7).
-As P7 noted, “By highlighting and collaging, I externalized formed ideas into image–text pairs, clearing mental space to pre-plan the next line and concentrate on the next plot beat.”
-Thus, while Vistoria required more effort and decision-making, it also freed participants from lower-level memory maintenance and allowed them to focus on higher-order creative synthesis.
-With greater familiarity, the efficiency gains from this offloading may outweigh the initial overhead.
-7. Discussion
-7.1. Cross-modal LLM-Mediated Instrumental Interaction
-We extend the notion of Instrumental Interaction into an LLM-mediated, cross-modal paradigm where generated artifacts repeatedly re-enter the workflow as manipulable outputs.
-In classic formulations, instruments mediate between users and domain objects, maintaining a separation between “tool” and “content” (Beaudouin-Lafon, 2000).
-In generative contexts, this dichotomy breaks down: each card (image–text pair) is simultaneously an outcome and a renewed object of operation, enabling recursive, non-linear composition (Riche et al., 2025).
-This reframing also reshapes reification.
-Vague creative intents expressed through text, sketches, or images materialize across modalities into manipulable, cross-linked artifacts that remain editable and reusable.
-Reification thus spans modalities rather than existing in a single channel, supporting direct manipulation and reflection on both structure and semantics of evolving narratives (Riche et al., 2025; Masson et al., 2024b).
-A key mechanism enabling this circulation is structural mapping (Gentner, 1983; Norman, 2013; Masson et al., 2024b).
-Participants made sense of their evolving materials not only by inspecting the surface properties of each medium but also by exploiting the correspondences in shared properties, concepts, or identities established between them.
-Refining a visual scene, for example, carried implications for narrative structure, tone, or pacing in text, while textual edits translated into compositional adjustments in images.
-In this way, mappings transformed what is usually a cognitive act of alignment into an operational resource: they stabilized coherence while keeping alternatives visible and accessible.
-Instrumental interaction and structural mapping thus operate as mutually reinforcing processes.
-Instruments generate the artifacts that mappings align, while mappings scaffold the interpretability and reuse of instruments.
-This dynamic carries several design implications for future systems.
-They should not only offer isolated multimodal tools but cultivate instrumental ecosystems in which mappings are first-class and inspectable, and where families of instruments carry coherent semantics across domains (Riche et al., 2025).
-Furthermore, instead of overwhelming users with immediate complexity, operations can be gradually disclosed, allowing mappings to surface progressively and adapt to context (Shneiderman and Maes, 1997; Allen et al., 1999).
-Provenance views that expose the cascade of edits across modalities could support both divergence and recombination, helping individuals and teams understand how narratives branch and converge (Masson et al., 2024b).
-Ultimately, the LLM can be positioned less as a generator of content and more as a mediator of instruments that align structures, propagate semantics, and synthesize new operations from prior materials.
-This shift reimagines authorship as an orchestration of evolving, cross-modal instruments, opening a design space where exploration and coherence can be sustained simultaneously (Sawyer, 2017; Masson et al., 2024b).
-7.2. Orchestration and Curatorial Authorship to Form Creativity in Cross-Modal Creation
-Traditional creativity theories emphasize individual expression through language and pre-formed intent (Smith, 2020; Donnelly, 2011), positioning the creator as the primary originator of content.
-Our findings challenge this view by demonstrating that creativity in Vistoria manifested as orchestration.
-Participants conducted and harmonized streams of multimodal LLM outputs through selection, combination, and reframing.
-Creative goals emerged through interaction with generated materials rather than preceding them, with meaning composed across representational boundaries.
-This suggests creativity in AI-augmented contexts should be theorized as the strategic coordination of evolving possibilities, repositioning authorship from origination to curation and synthesis within a collaborative ecosystem.
-While existing multimodal creativity research often focuses on seamless integration between modes (Yen et al., 2025; Rosenberg et al., 2024; Shi et al., 2025), our study reveals that productive tensions between text and image became central to the creative process.
-Participants actively negotiated the trade-offs between imagistic specificity and narrative openness, treating these tensions not as problems to solve but as resources for exploration.
-The ongoing adjudication process—accept, adapt, defer, or discard–became the primary engine of creative coherence.
-This extends theories of creative constraint by showing how representational tensions across modalities can drive rather than hinder creative development.
-Furthermore, our findings contribute to debates about human agency in LLM-assisted creation by demonstrating that curatorial work enhanced rather than diminished participants’ sense of ownership.
-Authorship was experienced as steering a dialogue with the model, including deciding what counts, where elements belong, and how disparate fragments accrue into a unified voice.
-This sense of agency grew precisely from having meaningful choices to negotiate among viable alternatives (Resnick et al., 2005).
-This suggests a new model of distributed authorship where human control emerges through deliberate curation within a space of LLM-generated possibilities.
-As for design implications, our findings suggest that future system design should prioritize support for orchestration over seamless, one-shot generation.
-Instead of producing singular outputs, tools could by default surface multiple cross-modal alternatives, accompanied by lightweight mechanisms that enable users to accept, adapt, defer, or discard options as part of an ongoing multimodal curatorial process (Resnick et al., 2005).
-Moreover, systems should not only present outputs but also make the inherent tensions between modalities visible and adjustable.
-This can be done, for example, through interactive controls such as ambiguity sliders that allow users to navigate between imagistic specificity and narrative openness.
-Beyond simple interface controls, future systems could also provide deeper support for curatorial authorship.
-For example, they might keep track of where each fragment or idea came from across different iterations, or offer reusable “curation moves” that let users quickly repeat and adapt common ways of combining or editing materials (Shneiderman, 2007).
-In this way, design emphasis shifts away from achieving frictionless integration and toward helping users actively steer evolving creative possibilities, foregrounding authorship in AI-mediated contexts as a matter of curation, synthesis, and negotiated choice (Ding, 2024).
-7.3. Limitation and Future Work
-While our study provides valuable insights into multimodal story writing, several limitations constrain the generalizability and scope of our findings.
-Task Scope and Short-Term Focus.
-The 300–-500 word story task, while manageable for controlled evaluation, does not reflect the demands of long-form fictional writing, where authors build sustained voices, complex arcs, and intricate structures.
-We did not have an opportunity to study Vistoria’s suitability for extended projects, iterative revisions, or complex narratives, leaving open questions on the consistency in long works, scalability with larger volumes, or risks of over-reliance on LLM over the time of use.
-Moreover, evaluation relied primarily on self-reports of creativity and user experience; we did not include objective measures of story quality, originality, or literary merit.
-In the future, we plan on conducting field studies to deploy this system with writers of varying expertise levels in their authentic creative contexts, observing how they integrate the tool into real writing projects over extended periods.
-Such longitudinal research would assess the ecological validity of Vistoria and provide insights into how writers adapt Vistoria for planning and translating across longer creative cycles.
-We anticipate that writers might spontaneously capture inspirational moments from daily life, potentially increasing their reliance on clustering functionality as they generate more dispersed content fragments that require organization.
-These naturalistic studies would provide crucial insights into the tool’s role in sustained creative practice, revealing usage patterns, adaptation strategies, and long-term impacts on writers’ creative processes that controlled laboratory settings cannot capture.
-Participant Sample.
-Our study involved only 12 participants, while this is typical for similar lab usability studies, a larger group could provide stronger statistical power, reveal more varied interaction patterns, and allow comparisons across subgroups.
-Furthermore, the participant pool has limited cultural and age diversity, which could have narrowed the range of narrative traditions, writing styles, and storytelling approaches represented.
-Future research should address these limitations by recruiting a larger and more diverse set of participants, including professional authors, older and younger writers, and individuals from varied cultural backgrounds, to more fully evaluate the system’s applicability and generalizability.
-8. Conclusion
-This paper introduced Vistoria, a multimodal text–image co-editing system designed to support the cyclical “planning–translation” process in fictional writing.
-Grounded in Wizard-of-Oz co-design study and theories of Instrumental Interaction and Structural Mapping, the system integrates four key capabilities: (1) representing sketches, texts, and references as manipulable card units; (2) enabling cross-modal alignment so that edits in one modality propagate to the other; (3) providing polymorphic operations that apply consistently across text and image; and (4) supporting organization and reuse through highlighting and clustering panels.
-A controlled study with 12 creative writers demonstrated that Vistoria enhanced writers’ expressiveness, immersion, and sense of collaboration, while broadening exploration of narrative branches and reinforcing authorship and control.
-These findings suggest that Vistoria transforms the writing process into a dialogic interaction with multimodal instruments, enabling authors to negotiate meaning and control across modalities.
-Appendix A Appendix
-A.1. Writing Topics Used in the User Study
+Interaction records of all participants.
+The creative workflow begins with multimodal generation—primarily text, complemented by sketches or images to express intentions, followed by refinement and iteration using visual instruments, during which textual descriptions are continuously revised in parallel.
+Definition of specific behaviors: Generation includes multimodal creation of new cards using Creative Spark, Exact Craft, or Collage; image-based editing refers to operations such as Lasso, perspective shift, and filters; text-based editing covers modifications of generated story segments on the canvas as well as edits made in the text editor; other operations include updates through highlight, cluster, and upgrading global settings.
+The figure illustrates multimodal intent expression using text, sketches, and images together.
+User sketches provide spatial structure, while textual annotations specify narrative details.
+The system integrates these inputs to generate aligned image–text cards reflecting the intended scene.
+A.3. Writing Topics Used in the User Study
 The two writing topic prompts used in the user study were:
@@
 The air carries faint whispers in a language that sounds hauntingly familiar, almost like someone calling her name.
-A.2. Filters
+A.4. Filters
 The following are the types of filter supported by the filter instrument, showing how different types of filters are applied to image styles and mapped to text tone or emotion.
+Table 3. Filter types with corresponding image and text effects.
 Filter Image Effect Text Effect
-Warm Warm tones (gold, amber, red, orange, yellow), high exposure, strong contrast $\rightarrow$ evoke happiness, comfort, nostalgia Emphasizes positivity, vitality, intimacy
-Calm Cool tones (blue, green, purple) with balanced or lower saturation $\rightarrow$ convey calmness, wisdom, introspection Reflects contemplative and stable moods
+Warm Warm tones (gold, amber, red, orange, yellow), high exposure, strong contrast $\rightarrow$ evoke happiness, comfort, nostalgia Emphasize positivity, vitality, intimacy
+Calm Cool tones (blue, green, purple) with balanced or lower saturation $\rightarrow$ convey calmness, wisdom, introspection Reflect contemplative and stable moods
 Dramatic Deep blacks, sharp whites, directional lighting $\rightarrow$ create intensity, mystery, urgency Heightens stakes and emotional tension
-Dreamy Soft tones, lowered contrast, diffuse focus $\rightarrow$ suggest melancholy, intimacy, ethereality Supports subtle, nostalgic, introspective narration
-Monochrome Removal of color, emphasis on light, shadow, texture $\rightarrow$ evoke nostalgia, timelessness, artistry Adopts reflective and universal tone
-Table 4. Filter types with corresponding image and text effects.
-A.3. LLM Prompts
-Communication between different agents is carried out through JSON-formatted messages, ensuring efficient and reliable information transfer.
-All prompts are structured in the following format:
-  • Blue text indicates placeholders that will be replaced with task-specific information in different generation requests.
-  • Pink text represents the mandatory output type of information generated by the agent.
-The below is an example prompt for precise description generation of GPT-4o.
-You are a visual story developer who analyzes screenshots to decode user visualization intent and creates detailed story segments that bring their creative vision to life.
-Process: 1) Examine the screenshot to understand what specific story content the user wants generated by identifying: printed text({text} ) which is the primary indicator of what the user wants you to generate, handwritten text expressing the user’s desired content direction and story focus where generation should address gaps and missing details, any images or illustrations with reference text {previous_text} for additional context, and hand-drawn sketches representing scenes from the user’s imagination.
-Synthesize these elements to understand the user’s envisioned story.
-The generated story should mainly focus on filling in content not covered in ({text} ) instead of still remain unknown.
-2) If hand-drawn illustrations exist in the screenshot, return the information about the story scene conveyed in the illustration’s layout; if none exist, output ’none’.
-3) Generate Focused Story Content: Using the existing written passages {full_text} only as background context to ensure logical consistency, create a NEW detailed story segment that elaborates on a specific scene or moment the user wants to visualize, focuses primarily on the intent expressed in the screenshot rather than expanding the existing text, contains concrete details, character emotions, environmental descriptions, dialogue, and interactions, maintains consistency with the global theme {global_theme}, and can reference previous text {previous_text} if relevant to the visualization goal.
-4) The narrative should contain substantial plot or setting content, not just descriptive language.
-The generated story should introducing new, insightful elements based on the context and provide new direction of the story development that can reificate the story.
-The generated stories need to be imaginative with concrete content, not filled with uncertainties.
-Respond in JSON format: { "story": "A detailed paragraph of no larger than 100 words that creates NEW story content focused on the user’s screenshot intent, elaborating a specific scene with concrete details while maintaining logical consistency with the background context.", "intention": "The visualization intention read from the screenshot for story generation direction", "sketch_information":"Regarding line sketches, integrate them with story descriptions to capture the user’s envisioned layout and scene details communicated through hand-drawn imagery, directing the image generator to produce story scenes based on this layout guidance.
-Avoid generating stories that may trigger content moderation." }.
-Provide only the JSON response without markdown formatting or additional commentary.
-Let’s think step by step.
+Dreamy Soft tones, lowered contrast, diffuse focus $\rightarrow$ suggest melancholy, intimacy, ethereality Support subtle, nostalgic, introspective narration
+Monochrome Removal of color, emphasis on light, shadow, texture $\rightarrow$ evoke nostalgia, timelessness, artistry Adopt reflective and universal tone
+A.5. LLM Prompts
+A.6. Codebook
+Theme 1: Instrumental Interaction
+   $\vdash$ instrumental operations
+      $\vdash$ granularity control (using Lasso for detail extraction)
+      $\vdash$ multimodal recombination (using Collage)
+      $\vdash$ affective alignment (using Filters for tone)
+      $\vdash$ perspective shift (viewpoint transformation)
+      $\vdash$ open new narrative direction
+Theme 2: Cognitive Process
+   $\vdash$ externalization & traceability
+      $\vdash$ visual history / story evolution
+      $\vdash$ spatial organization (grouping)
+      $\vdash$ visual checkpoints
+   $\vdash$ cognitive offloading
+      $\vdash$ offloading working memory
+   $\vdash$ higher mental demand
+      $\vdash$ not familiar with operations
+      $\vdash$ substantial learning effort
+Theme 3: Creative Support through Multimodality
+   $\vdash$ bottom-up workflow
+      $\vdash$ open-ended exploration
+      $\vdash$ branching storylines
+      $\vdash$ comparing modalities
+      $\vdash$ divergent exploration
+   $\vdash$ inspiration
+      $\vdash$ serendipitous discovery (randomness as value)
+      $\vdash$ perspective transformation
+      $\vdash$ memory triggers
+   $\vdash$ visual–text interaction
+      $\vdash$ more vivid detailed description
+      $\vdash$ sense of immersion
+Theme 4: Ownership and Agency
+   $\vdash$ control & ownership
+      $\vdash$ resisting AI takeover
+      $\vdash$ active curation
+      $\vdash$ personal style alignment
+   $\vdash$ metaphors of use
+      $\vdash$ companion / sketchbook metaphor
+      $\vdash$ co-pilot
+      $\vdash$ free-exploration
+Figure 14. The final coding tree. Main themes are marked in bold; sub-codes represent specific strategies and behaviors observed in the study.
+The codebook tree of qualitative analysis